### v1

# Local Topological Constraints on Berry Curvature in Spin–Orbit Coupled BECs

Abstract
We establish a local topological obstruction to flattening Berry curvature in spin-orbit-coupled Bose-Einstein condensates (SOC BECs), valid even when the global Chern number vanishes. For a generic two-component SOC BEC, the extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ carries a natural metric connection $\nabla^{C}$ whose torsion 3-form encodes the synthetic gauge fields. Its harmonic part defines a mixed cohomology class
$\omega]\in\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
whose mixed tensor rank equals one. Using a general geometric bound for connections with skew torsion on product manifolds, we show that the obstruction kernel $\mathcal{K}$ vanishes, yielding the sharp inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$. This forces at least one off-diagonal curvature operator, preventing complete gauging-away of Berry phases even when the total Chern number is zero. This provides the first cohomological lower bound certifying locally irremovable curvature in SOC BECs beyond the Chern-number paradigm.
Keywords: Berry curvature; spin–orbit-coupled Bose–Einstein condensates; synthetic gauge fields; holonomy with torsion; mixed cohomology; Kaluza–Klein geometry.
MSC (2020): 53C29, 53C80, 81V45, 82B10, 51P05, 55S20, 57R19, 57R56, 58A12.
1 Introduction
The study of spin–orbit-coupled (SOC) Bose–Einstein condensates (BECs) stands at a rich confluence of differential geometry, quantum many-body physics, and synthetic gauge theory. This interdisciplinary effort is part of the broader pursuit of artificial gauge fields in engineered quantum systems, spanning ultracold atoms, photonics, and solid-state platforms [1]. Advances in generating artificial gauge fields with Raman lasers have made these systems a versatile laboratory for exploring Berry curvature, synthetic magnetic fields, and non-Abelian textures with unprecedented control [6, 7, 14]. These studies are often conducted in toroidal traps, which can serve as magnetic storage rings for BECs [3], providing a well-controlled environment for synthetic gauge field experiments. Remarkable achievements include the observation of Dirac monopoles in synthetic magnetic fields [10], demonstrating the high degree of geometric control now attainable. The geometric phase underpinning these phenomena traces back to the foundational work of Simon [13], while the experimental realization of artificial gauge fields has been comprehensively reviewed in [4].
A persistent theoretical challenge is to identify local topological obstructions that constrain the curvature even when global invariants—such as the first Chern number—vanish. While the Chern number quantizes the net flux, its vanishing does not guarantee that the curvature can be locally flattened; geometric obstructions may prevent its simultaneous annihilation along independent directions in the parameter space.
This work addresses that challenge by bringing a recent geometric result into physical focus. In [12], a lower bound was established for the off-diagonal holonomy of metric connections with totally skew-symmetric torsion on product manifolds. The bound is expressed in terms of a mixed de Rham cohomology class $\omega]\in H^{p}(M_{1})\otimes H^{q}(M_{2})$ associated with the torsion; its mixed tensor rank $r$ counts the minimal number of simple tensors needed to represent the class, and after subtracting contributions from parallel forms (which do not generate off-diagonal curvature) one obtains a corrected integer $r^{\sharp}$ that bounds the dimension of the off-diagonal holonomy algebra:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim\mathcal{K}.$
Here we show how this abstract geometric theorem materializes in a concrete physical setting: a generic two-component SOC BEC confined in a quasi-two-dimensional toroidal trap. Such a system possesses three universal, experimentally accessible degrees of freedom: the two crystal momenta forming a Brillouin-zone torus $T^{2}_{\mathrm{BZ}}$, a global $U(1)$ phase $\phi_{+}$ associated with particle-number conservation, and a relative $U(1)$ phase $\phi_{-}$ between the two internal spin components [9, 11]. We therefore take as the extended parameter space the compact product manifold.
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}.$
The first main task of the paper is to construct, in this space, a natural geometric structure to which the theorem of [12] applies. We introduce a metric $g_{M}$ of Kaluza–Klein type, which couples the base metric on $T^{2}_{\mathrm{BZ}}$ to the $U(1)$-valued connection forms that encode the synthetic gauge fields of the SOC BEC. On $(M,g_{M})$ we define a metric connection $\nabla^{C}$ whose torsion three-form is precisely the physical three-form
$T^{\flat}=\operatorname{Tr}(\mathcal{F}\wedge\mathbf{A}_{\mathrm{syn}}),$
built from the synthetic Berry curvature $\mathcal{F}$ and connection $\mathbf{A}_{\mathrm{syn}}$. The harmonic part of $T^{\flat}$ with respect to $g_{M}$ then defines a mixed cohomology class
$\omega]\in H^{3}(M)\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr).$
The second main task is to compute the topological data that enters the general bound. We prove that, under the generic condition that the two nearly degenerate bands carry non-zero Berry curvature classes $F^{(+)}],[F^{(-)}]\in H^{2}(T^{2}_{\mathrm{BZ}})$, the mixed tensor rank of $\omega]$ attains its maximal possible value within this geometry. Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, any two non-zero classes in this space are cohomologically proportional. Consequently, the mixed class $\omega]$ can be written as a single simple tensor
$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
and its mixed tensor rank is therefore $r=1$. Moreover, owing to the specific form of the metric $g_{M}$, the spaces of parallel one-forms on the phase circles are trivial, which forces the obstruction kernel $\mathcal{K}$ to vanish. Consequently the reduced rank is $r^{\sharp}=1$, and the general theorem yields the sharp inequality
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
This lower bound has immediate physical consequences. It implies that the curvature of the synthetic gauge connection cannot be made block-diagonal with respect to the splitting between momentum and phase directions; at least one independent off-diagonal operator persists, mixing the Brillouin zone with the two phase directions. Crucially, the bound detects the topological coupling of the momentum torus to the two distinct $U(1)$ phases, and therefore remains non-trivial even when the total Chern number of the lowest band is tuned to zero. Thus $r^{\sharp}$ provides a cohomological certificate of locally irremovable Berry curvature, invisible to conventional global invariants. The result suggests interferometric protocols that separately probe the Berry phases associated with $\phi_{+}$ and $\phi_{-}$, offering a way to detect this obstruction even in regimes of vanishing net vorticity.
The paper is organized as follows. Section 2 constructs the extended parameter space $M$, the Kaluza–Klein metric $g_{M}$, and the metric connection $\nabla^{C}$ with torsion $T^{\flat}$. Section 3 analyses the cohomology of $M$, computes the mixed tensor rank $r$, and shows that $\mathcal{K}=0$ in the present geometry. Section 4 extends the Pigazzini–Toda theorem to Kaluza–Klein metrics, ensuring the bound applies to the present geometry. Section 5 applies the general theorem of [12] to obtain the lower bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ and discusses its physical interpretation. Section 6 elaborates the physical implications of the bound, including its persistence when the total Chern number vanishes and its robustness under metric deformations. Section 7 illustrates how the same cohomological data reproduce, in a particular Rashba–Dresselhaus texture, the familiar Chern-number bound on the vortex number, and demonstrates with a tuned example that the holonomy bound remains valid even when the Chern number vanishes. Section 8 outlines prospective interferometric tests for detecting the locally irremovable curvature and indicates directions for future work. Section 9 summarises the results and discusses broader implications for topological quantum matter.
2 Geometric Construction of the Extended Parameter Space
A generic two-component spin–orbit-coupled Bose–Einstein condensate (SOC BEC) confined in a quasi-two-dimensional toroidal trap is characterized by three independent, experimentally accessible degrees of freedom: the crystal momenta $(k_{x},k_{y})$ spanning a Brillouin zone, a global $U(1)$ phase $\phi_{+}$ associated with particle-number conservation, and a relative $U(1)$ phase $\phi_{-}$ between the two internal spin states [9, 11].
We therefore take as the extended parameter space the smooth, compact, oriented product manifold
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}.$
Our goal in this section is to endow $M$ with a Riemannian structure that naturally incorporates the synthetic gauge fields of the SOC BEC, and to define on it a metric connection whose torsion three-form coincides with the physical Berry-curvature three-form.
2.1 Kaluza–Klein metric on the extended space
Let $g_{\mathrm{BZ}}=dk_{x}^{2}+dk_{y}^{2}$ be the flat metric on the Brillouin-zone torus $T^{2}_{\mathrm{BZ}}$. On each phase circle we introduce the standard angular coordinate with period $2\pi$; the canonical $U(1)$-invariant metric on $S^{1}$ is $d\phi^{2}$. The synthetic gauge fields of the SOC BEC provide, for each of the two independent $U(1)$ directions, a Berry connection. Denote by $A^{(+)},A^{(-)}\in\Omega^{1}(T^{2}_{\mathrm{BZ}})$ the corresponding real one-forms (the components of the synthetic $\mathfrak{su}(2)$ connection along the two independent $U(1)$ generators). Their curvatures are the Berry curvature two-forms $F^{(\pm)}=dA^{(\pm)}$.
Following the Kaluza–Klein prescription we define a metric $g_{M}$ on $M$ that couples the base metric to the connection forms:
$g_{M}=\pi^{*}g_{\mathrm{BZ}}+\bigl(d\phi_{+}+\pi^{*}A^{(+)}\bigr)^{2}+\bigl(d\phi_{-}+\pi^{*}A^{(-)}\bigr)^{2},$ (2.1)
where $\pi:M\to T^{2}_{\mathrm{BZ}}$ is the projection onto the momentum coordinates. The one-forms
$\Theta^{(+)}:=d\phi_{+}+\pi^{*}A^{(+)},\qquad\Theta^{(-)}:=d\phi_{-}+\pi^{*}A^{(-)}$
are the horizontal lifts of the phase coordinates; they are globally defined on $M$ and are orthogonal to the base directions with respect to $g_{M}$.
Lemma 2.1 (Properties of the metric $g_{M}$).
The metric (2.1) satisfies the following.
  1. 1.
    $g_{M}$ is smooth, positive definite and makes $\pi:(M,g_{M})\to(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ a Riemannian submersion.
  2. 2.
    The fibres $\pi^{-1}(k)\cong S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ are flat $2$-tori with the induced metric $(d\phi_{+})^{2}+(d\phi_{-})^{2}$.
  3. 3.
    The horizontal distribution $\mathcal{H}=\ker(d\phi_{+}+\pi^{*}A^{(+)})\cap\ker(d\phi_{-}+\pi^{*}A^{(-)})$ is orthogonal to the vertical distribution $\mathcal{V}=\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$.
  4. 4.
    If $A^{(\pm)}=0$ then $g_{M}$ reduces to the trivial product metric $g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$.
Proof.
Smoothness follows from the smoothness of $g_{\mathrm{BZ}}$ and $A^{(\pm)}$. Positive definiteness is clear because each term in (2.1) is non‑negative and their sum vanishes only when all differentials vanish.
For any $k\in T^{2}_{\mathrm{BZ}}$ the restriction of $g_{M}$ to the fibre $\pi^{-1}(k)$ is $(d\phi_{+})^{2}+(d\phi_{-})^{2}$, independent of $k$; hence the fibres are isometric to a flat $2$-torus. The map $\pi$ is a Riemannian submersion because, by construction, $g_{M}$ restricted to $\mathcal{H}$ coincides with $\pi^{*}g_{\mathrm{BZ}}$.
Orthogonality $\mathcal{H}\perp\mathcal{V}$ follows directly from $g_{M}(\partial_{\phi_{\pm}},X)=0$ for every $X\in\mathcal{H}$, which is a consequence of the definition of $\Theta^{(\pm)}$.
The last statement is immediate from (2.1). ∎
Remark 2.2.
Equation (2.1) is physically natural: it measures distances in the extended parameter space by combining the kinetic contribution from momentum changes with the “phase cost” induced by the synthetic gauge potentials. If $A^{(\pm)}=0$, the metric reduces to the trivial product metric $g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$.
2.2 Metric connection with prescribed torsion
On the Riemannian manifold $(M,g_{M})$ we now construct a metric connection $\nabla^{C}$ whose torsion three-form is built from the physical gauge fields. Let $\nabla^{LC}$ be the Levi‑Civita connection of $g_{M}$.
Define the three‑form
$T^{b}:=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}\in\Omega^{3}(M),$ (2.2)
where $F^{(\pm)}\wedge A^{(\pm)}$ are pulled back to $M$ via $\pi$ (we keep the same notation for the pull‑backs). $T^{b}$ is the trace $\operatorname{Tr}(\mathcal{F}\wedge\mathcal{A})$ with respect to the Killing form on $\mathfrak{u}(1)_{+}\oplus\mathfrak{u}(1)_{-}$, where $\mathcal{A}=A^{(+)}\otimes t_{+}+A^{(-)}\otimes t_{-}$ and $\mathcal{F}=F^{(+)}\otimes t_{+}+F^{(-)}\otimes t_{-}$.
Introduce the (2,1)(2,1)‑tensor $K$ by
$K(X,Y,Z):=\frac{1}{2}\,T^{b}(X,Y,Z)\qquad(X,Y,Z\in\mathfrak{X}(M)).$
The connection
$\nabla^{C}_{X}Y:=\nabla^{LC}_{X}Y+K(X,Y,\cdot)^{\sharp}$
is by construction metric ($\nabla^{C}g_{M}=0$) and has totally skew‑symmetric torsion $T^{b}$. Indeed, for any vector fields $X,Y$,
$T^{b}(X,Y,\cdot)^{\sharp}=\nabla^{C}_{X}Y-\nabla^{C}_{Y}X-[X,Y],$
which is precisely the definition of the torsion tensor of $\nabla^{C}$.
Because $T^{b}$ is not parallel with respect to $\nabla^{LC}$ unless the gauge fields are flat, the connection $\nabla^{C}$ satisfies the non‑parallelism condition $\nabla^{LC}T^{b}\neq 0$ required for the applicability of the general theorem [12].
Remark 2.3.
The three‑form $T^{b}$ is not closed in general; its exterior derivative contains terms quadratic in the curvatures $F^{(\pm)}$. Its de Rham cohomology class $T^{b}]\in H^{3}(M;\mathbb{R})$, however, depends only on the harmonic part of $T^{b}$ with respect to $g_{M}$, as will be computed in Section 3.
2.3 Parallel forms in the Kaluza–Klein geometry
A key ingredient of the topological bound is the space of parallel forms on the factors. For the base torus with the flat metric $g_{\mathrm{BZ}}$ we have the well‑known result
$\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})=H^{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}},\qquad\mathrm{vol}_{\mathrm{BZ}}=dk_{x}\wedge dk_{y}.$
The situation for the phase circles is more subtle. On a circle endowed with the standard metric $d\phi^{2}$, the one‑form $d\phi$ is parallel, hence $\mathcal{P}_{1}(S^{1},d\phi^{2})=\mathbb{R}\,d\phi$. In the Kaluza–Klein metric (2.1) the geometry of the fibres is “twisted” by the connection forms $A^{(\pm)}$. The vector fields $\partial_{\phi_{+}},\partial_{\phi_{-}}$ are Killing fields of $g_{M}$, but their dual one‑forms are $\Theta^{(\pm)}$, not $d\phi_{\pm}$.
Lemma 2.4 (Non‑parallelism of $d\phi_{\pm}$).
Assume $A^{(\pm)}$ are not identically zero. Then the one‑forms $d\phi_{+},d\phi_{-}$ are not parallel with respect to $g_{M}$.
Proof.
We compute the covariant derivative $\nabla^{LC}d\phi_{+}$ in local coordinates $(k_{x},k_{y},\phi_{+},\phi_{-})$. The non‑zero Christoffel symbols involving the index $\phi_{+}$ are
$\Gamma^{\phi_{+}}_{k_{x}\phi_{+}}=\frac{1}{2}g^{\phi_{+}\rho}\bigl(\partial_{k_{x}}g_{\phi_{+}\rho}+\partial_{\phi_{+}}g_{k_{x}\rho}-\partial_{\rho}g_{k_{x}\phi_{+}}\bigr).$
Since $g_{k_{x}\phi_{+}}=A^{(+)}_{k_{x}}$ is independent of $\phi_{+}$, we have $\partial_{\phi_{+}}g_{k_{x}\phi_{+}}=0$. Moreover $g_{\phi_{+}\phi_{+}}=1+(A^{(+)}_{k_{x}})^{2}+(A^{(+)}_{k_{y}})^{2}$ and $g_{\phi_{+}\phi_{-}}=A^{(+)}_{k_{x}}A^{(-)}_{k_{x}}+A^{(+)}_{k_{y}}A^{(-)}_{k_{y}}$. Consequently,
[…]
which is non‑zero for generic non‑constant potentials $A^{(\pm)}$. Hence $(\nabla^{LC}_{\partial_{k_{x}}}d\phi_{+})(\partial_{\phi_{+}})=-\Gamma^{\phi_{+}}_{k_{x}\phi_{+}}\neq 0$, showing that $d\phi_{+}$ is not parallel. The same argument applies to $d\phi_{-}$. ∎
Corollary 2.5 (Trivial parallel‑form spaces on the fibres).
For a generic SOC BEC with non‑trivial Berry curvature $F^{(\pm)}\neq 0$,
$\mathcal{P}_{1}\bigl(S^{1}_{\phi_{\pm}},g_{M}|_{\mathrm{fibre}}\bigr)=\{0\}.$
More precisely, any fibre‑tangent 1‑form that is parallel with respect to the full metric $g_{M}$ must be of the form $\xi=c_{+}d\phi_{+}+c_{-}d\phi_{-}$ with constants $c_{+},c_{-}$. However, Lemma 2.4 shows that for non‑constant potentials $A^{(\pm)}$, $\nabla^{LC}d\phi_{\pm}$ has non‑vanishing components along the base directions; consequently no non‑zero linear combination can be parallel on the whole manifold $M$. If $A^{(\pm)}$ are constant but $F^{(\pm)}\neq 0$, they are not closed, hence still not harmonic and therefore not parallel (on a compact Riemannian manifold every parallel form is harmonic).
2.4 Summary of the geometric setup
We have constructed:
  1. 1.
    an extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$;
  2. 2.
    a Kaluza–Klein metric $g_{M}$ on $M$ given by (2.1), coupling the momentum directions to the two independent $U(1)$ phases via the synthetic Berry connections $A^{(\pm)}$;
  3. 3.
    a metric connection $\nabla^{C}$ on $(M,g_{M})$ whose torsion three‑form is the physical quantity $T^{b}=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}$.
This construction provides the precise Riemannian framework needed to apply the general theorem of [12]. In the following section we analyze the cohomological data that arise from this geometry.
3 Cohomology of the Extended Space and the Mixed Tensor Rank
With the geometric structure $(M,g_{M},\nabla^{C})$ established in Section 2, we now analyze the cohomological data that enter the general lower bound. Recall that the torsion three-form $T^{\flat}\in\Omega^{3}(M)$ is not necessarily closed, but its de Rham cohomology class $T^{\flat}]\in H^{3}(M;\mathbb{R})$ is well defined. Let $\omega_{h}$ be the unique $g_{M}$-harmonic representative of this class; then
$\omega]:=[\omega_{h}]=[T^{\flat}]\in H^{3}(M;\mathbb{R}).$
3.1 Künneth decomposition of $H^{3}(M)$
Since $M$ is a product $T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ and each factor is oriented and compact, the Künneth theorem gives an isomorphism
$H^{3}(M;\mathbb{R})\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
because $H^{3}(T^{2}_{\mathrm{BZ}})=H^{2}(S^{1}_{\phi_{\pm}})=0$ and $H^{1}(T^{2}_{\mathrm{BZ}})$ does not contribute to degree $3$. Consequently $\dim H^{3}(M)=2$, and every class $\eta]\in H^{3}(M)$ can be written uniquely as
$\eta]=[\alpha_{+}]\otimes[d\phi_{+}]+[\alpha_{-}]\otimes[d\phi_{-}],$ (3.1)
where $\alpha_{\pm}]\in H^{2}(T^{2}_{\mathrm{BZ}})$ and $d\phi_{\pm}]\in H^{1}(S^{1}_{\phi_{\pm}})$ are the fundamental cohomology classes of the factors.
For the harmonic representative $\omega_{h}$ of our torsion class, the Hodge decomposition with respect to $g_{M}$ yields
$\omega_{h}=\alpha_{+}\wedge d\phi_{+}+\alpha_{-}\wedge d\phi_{-}+d\eta,$
where $\alpha_{\pm}\in\mathcal{H}^{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ are $g_{\mathrm{BZ}}$-harmonic two-forms on the Brillouin torus, and $\eta\in\Omega^{2}(M)$ is a coexact two-form. The term $d\eta$ is exact and therefore orthogonal to harmonic forms in the $L^{2}$-inner product; it does not contribute to the cohomology class $\omega]=[\omega_{h}]$.
Remark 3.1.
The representatives $d\phi_{\pm}$ are harmonic with respect to the standard metric on the circles, but they are not the harmonic representatives of the fundamental classes with respect to the full Kaluza–Klein metric $g_{M}$; the $g_{M}$-harmonic representatives are the horizontal forms $\Theta^{(\pm)}=d\phi_{\pm}+A^{(\pm)}$ introduced in (2.1). Because $d\phi_{\pm}]=[\Theta^{(\pm)}]$ in $H^{1}(S^{1}_{\phi_{\pm}})$, the cohomological decomposition (3.1) is unchanged.
3.2 Mixed tensor rank of the torsion class
The mixed tensor rank of a cohomology class $\eta]\in H^{3}(M)$ written as in (3.1) is defined as the minimal number of simple tensors required to represent it:
$\operatorname{rank}_{\mathbb{R}}([\eta]):=\min\Bigl\{m\;\Big|\;[\eta]=\sum_{i=1}^{m}[\beta_{i}]\otimes[\gamma_{i}],\;[\beta_{i}]\in H^{2}(T^{2}_{\mathrm{BZ}}),\;[\gamma_{i}]\in H^{1}(S^{1}_{\phi_{+}})\oplus H^{1}(S^{1}_{\phi_{-}})\Bigr\}.$
For the class $\omega]$ coming from the physical torsion, the two-forms $\alpha_{+},\alpha_{-}$ in (3.1) are precisely the (cohomology classes of the) Berry curvature two-forms $F^{(+)},F^{(-)}$ of the two independent $U(1)$ sectors. A key observation is that $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$. Therefore, any two non-zero classes in this space are necessarily cohomologically proportional. In a generic two-component SOC BEC, both Berry curvature classes $F^{(+)}]$ and $F^{(-)}]$ are non-zero; consequently, there exists a non-zero constant $\lambda\in\mathbb{R}$ such that
$F^{(-)}]=\lambda[F^{(+)}]\quad\text{in }H^{2}(T^{2}_{\mathrm{BZ}}).$
Substituting this relation into (3.1) gives
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+\lambda[F^{(+)}]\otimes[d\phi_{-}]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
which is a single simple tensor. Hence the mixed tensor rank attains the value
$r=1.$
Remark 3.2.
Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the maximal possible mixed tensor rank for a class of the form $\alpha_{+}]\otimes[d\phi_{+}]+[\alpha_{-}]\otimes[d\phi_{-}]$ with both $\alpha_{\pm}]\neq 0$ is $1$. All experimentally prominent SOC geometries (Rashba–Dresselhaus, Hofstadter, Raman-induced couplings) satisfy the condition $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$, thereby realizing the generic case $r=1$.
Remark 3.3.
Since $M_{2}=S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ is a product, its first cohomology decomposes as $H^{1}(M_{2})\cong H^{1}(S^{1}_{+})\oplus H^{1}(S^{1}_{-})$. Therefore an element of $H^{2}(M_{1})\otimes H^{1}(M_{2})$ can be written as a sum of terms of the form $\alpha]\otimes[d\phi_{+}]+[\beta]\otimes[d\phi_{-}]$. The mixed tensor rank $r$ is then the minimal number of simple tensors needed to represent this sum in the space $H^{2}(M_{1})\otimes\bigl(H^{1}(S^{1}_{+})\oplus H^{1}(S^{1}_{-})\bigr)$.
3.3 Vanishing of the obstruction kernel $\mathcal{K}$
The general bound of [12] involves a correction term $\dim\mathcal{K}$, where
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\ker\Psi_{p}=\pi_{2,1}([\omega])\cap\bigl(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})\bigr),$
and $\ker\widetilde{\Psi}_{p}$ is defined analogously with the roles of the two circles exchanged. Here $\pi_{2,1}$ denotes the projection onto the $(2,1)$-Künneth component.
In our geometry, we have $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}}$ (the space of parallel two-forms on the flat torus). The crucial observation, already made in Section 2.3, is that for the Kaluza–Klein metric $g_{M}$ the spaces of parallel one-forms on the phase circles are trivial:
$\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},\,g_{M}|_{\text{\rm fibre}})=\{0\},$
because any $g_{M}$-parallel one-form tangent to a fibre would have to be a constant multiple of $\Theta^{(\pm)}=d\phi_{\pm}+A^{(\pm)}$, and $\Theta^{(\pm)}$ is not closed (hence not of the form $d\phi_{\pm}$) unless $F^{(\pm)}=0$. Since the Berry curvatures $F^{(\pm)}$ are generically non-zero, we indeed have $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})=\{0\}$ in the physical geometry.
Consequently the tensor products $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})$ and $\mathcal{P}_{1}(S^{1}_{\phi_{+}})\otimes\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})$ (the latter appears in $\ker\widetilde{\Psi}_{p}$) are zero. Thus
$\mathcal{K}=0,\qquad\text{and therefore}\qquad\dim\mathcal{K}=0.$ (3.2)
3.4 Summary of the topological data
Based on (3.2) we obtain the reduced rank
$r^{\sharp}\;:=\;r-\dim\mathcal{K}\;=\;1.$
All ingredients required for the application of the general geometric theorem are now in place: the extended parameter space $(M,g_{M})$ carries a metric connection $\nabla^{C}$ with totally skew torsion whose harmonic part represents a mixed cohomology class $\omega]$ of mixed tensor rank $r=1$, and the obstruction kernel $\mathcal{K}$ vanishes identically. In the next section, we apply the theorem to obtain a concrete lower bound on the off-diagonal holonomy of $\nabla^{C}$.
4 Geometric foundations for the PT theorem in Kaluza–Klein geometry
The physical metric (2.1) is not a Riemannian product, whereas Theorem 5.2 of PT [12] is stated for genuine product metrics. The following three results show that the theorem nevertheless applies to our Kaluza–Klein geometry, because the relevant topological and parallel-form data coincide with those of an auxiliary product metric.
4.1 Parallel forms in Kaluza–Klein metrics
Lemma 4.1 (Parallel forms in Kaluza–Klein metrics with flat fibres).
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a Kaluza–Klein metric
$g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2},$ (4.1)
where $g_{1}$ is a Riemannian metric on $M_{1}$, $\{\theta^{a}\}$ defines a flat left-invariant metric $h=\sum_{a}(\theta^{a})^{2}$ on $M_{2}$, and $A^{a}\in\Omega^{1}(M_{1})$ (possibly with $dA^{a}\neq 0$).
Then the restriction map
$\Phi_{1}:\mathcal{P}_{k}(M_{1},g)\longrightarrow\mathcal{P}_{k}(M_{1},g_{1}),\qquad\eta\mapsto\eta|_{D_{1}},$
is a well-defined linear isomorphism.
Analogously, define $\mathcal{P}_{k}(M_{2},g)$ as the space of forms $\xi\in\Omega^{k}(M)$ that are invariant along $M_{1}$ (their coefficients depend only on the $M_{2}$-coordinates) and satisfy $\nabla^{LC}\xi=0$. Then
$\Phi_{2}:\mathcal{P}_{k}(M_{2},g)\longrightarrow\mathcal{P}_{k}(M_{2},h),\qquad\xi\mapsto\xi|_{D_{2}},$
is an isomorphism.
Proof.
Choose local coordinates $(x^{i},y^{\alpha})$ adapted to the orthogonal splitting $TM=D_{1}\oplus D_{2}$ induced by the product structure. In these coordinates the metric reads
$g=g_{ij}(x)\,dx^{i}dx^{j}+h_{\alpha\beta}(y)\bigl(dy^{\alpha}+A^{\alpha}_{i}(x)dx^{i}\bigr)\bigl(dy^{\beta}+A^{\beta}_{j}(x)dx^{j}\bigr).$ (4.2)
Let $\eta\in\mathcal{P}_{k}(M_{1},g)$. Because $\eta$ is parallel and tangent to $D_{1}$, in the coordinates above it can be written as
$\eta=\frac{1}{k!}\,\eta_{i_{1}\dots i_{k}}(x)\,dx^{i_{1}}\wedge\dots\wedge dx^{i_{k}}.$
We compute the Christoffel symbols of $g$. From (4.2) one obtains
$\Gamma^{i}_{jk}$ $=\Gamma^{(g_{1})i}_{jk},$
$\Gamma^{i}_{j\beta}$ $=-\frac{1}{2}\,g^{i\ell}\bigl(\partial_{j}g_{\ell\beta}+\partial_{\beta}g_{j\ell}-\partial_{\ell}g_{j\beta}\bigr),$
$\Gamma^{\alpha}_{jk}$ $=\frac{1}{2}\,h^{\alpha\gamma}\bigl(\partial_{j}g_{k\gamma}+\partial_{k}g_{j\gamma}-\partial_{\gamma}g_{jk}\bigr),$
$\Gamma^{\alpha}_{\beta\gamma}$ $=\Gamma^{(h)\alpha}_{\beta\gamma},$
where $\Gamma^{(g_{1})}$ and $\Gamma^{(h)}$ denote the Christoffel symbols of $g_{1}$ and $h$ respectively.
For any vector field $X=X^{i}\partial_{i}+X^{\alpha}\partial_{\alpha}$,
$(\nabla^{LC}_{X}\eta)_{i_{1}\dots i_{k}}=X^{j}\partial_{j}\eta_{i_{1}\dots i_{k}}-\sum_{r=1}^{k}\Gamma^{\ell}_{ji_{r}}X^{j}\eta_{i_{1}\dots i_{r-1}\ell i_{r+1}\dots i_{k}}-\Gamma^{\alpha}_{j\alpha}X^{j}\eta_{i_{1}\dots i_{k}}.$
The condition $\nabla^{LC}\eta=0$ therefore splits into two independent sets of equations. For $X=X^{i}\partial_{i}$ (tangent to $M_{1}$) we obtain
$\nabla^{(g_{1})}_{X}\eta=0,$
where $\nabla^{(g_{1})}$ is the Levi‑Civita connection of $g_{1}$. Hence $\eta|_{D_{1}}$ is $g_{1}$-parallel.
For $X=X^{\alpha}\partial_{\alpha}$ (tangent to $M_{2}$) we get
$\Gamma^{\alpha}_{j\alpha}(x)X^{j}\eta_{i_{1}\dots i_{k}}=0\qquad\text{for all }X^{j}.$
Why $\Gamma^{\alpha}_{j\alpha}=0$. Because the fibres are flat, the fibre metric $h_{\alpha\beta}$ is constant in the chosen coordinates; therefore $\partial_{j}h_{\alpha\beta}=0$. Moreover, from (4.2) we have $g_{j\alpha}=h_{\alpha\beta}A^{\beta}_{j}(x)$, which is independent of $y$. Consequently
$\partial_{\gamma}g_{j\alpha}=0,\qquad\partial_{\alpha}g_{j\gamma}=0.$
Substituting these facts into the expression for $\Gamma^{\alpha}_{j\alpha}$ gives
$\Gamma^{\alpha}_{j\alpha}=\frac{1}{2}h^{\alpha\gamma}\bigl(\partial_{j}h_{\alpha\gamma}+\partial_{\alpha}g_{j\gamma}-\partial_{\gamma}g_{j\alpha}\bigr)=0.$
Thus the condition coming from $X^{\alpha}\partial_{\alpha}$ is automatically satisfied.
Hence $\eta\in\mathcal{P}_{k}(M_{1},g)$ if and only if $\eta|_{D_{1}}\in\mathcal{P}_{k}(M_{1},g_{1})$. Consequently $\Phi_{1}$ is injective. Surjectivity follows because any $g_{1}$-parallel form $\widetilde{\eta}$ on $M_{1}$ can be lifted to a $g$-parallel form on $M$ by declaring it independent of the fibre coordinates; the computation above shows that the lifted form is indeed parallel with respect to $g$.
Therefore $\Phi_{1}$ is an isomorphism. The proof for $\Phi_{2}$ proceeds identically after exchanging the roles of base and fibre. ∎
Lemma 4.2 (Harmonic Künneth decomposition for flat‑fibre KK metrics).
Let $g$ be a Kaluza–Klein metric as in (4.1) with flat fibres. Then the natural wedge‑product map
$\Phi:\mathcal{H}^{p}(M_{1},g_{1})\otimes\mathcal{H}^{q}(M_{2},h)\longrightarrow\mathcal{H}^{p+q}(M,g)$
is an isomorphism onto the subspace of harmonic forms of pure bidegree $(p,q)$. In particular, every $g$-harmonic form $\eta\in\mathcal{H}^{p+q}(M,g)$ that belongs to the $(p,q)$-Künneth component admits a unique decomposition
$\eta=\sum_{i=1}^{r}\alpha_{i}\wedge\beta_{i},\qquad\alpha_{i}\in\mathcal{H}^{p}(M_{1},g_{1}),\;\beta_{i}\in\mathcal{H}^{q}(M_{2},h),$
with $r=\operatorname{rank}_{\mathbb{R}}([\eta]_{\mathrm{mixed}})$.
Proof.
Because the fibres are flat, the Hodge Laplacian of $g$ decouples as $\Delta_{g}=\Delta_{g_{1}}\otimes\mathrm{Id}+\mathrm{Id}\otimes\Delta_{h}$ on forms of pure bidegree. Hence a form of bidegree $(p,q)$ is $g$-harmonic iff its components are harmonic with respect to $g_{1}$ and $h$ respectively. The statement then follows from the classical Künneth theorem for de Rham cohomology combined with Hodge theory. ∎
4.2 Persistence of non‑parallelism under metric change
For the application of the PT theorem to an auxiliary product metric, we need to ensure that the non‑parallelism condition $\nabla^{LC}T\neq 0$ is preserved when we replace the original Kaluza–Klein metric by a nearby product metric.
Lemma 4.3 (Openness of the non‑parallelism condition).
Let $M$ be a smooth manifold and $T\in\Omega^{3}(M)$ a fixed 3‑form. Denote by $\mathcal{M}$ the space of Riemannian metrics on $M$ endowed with the $C^{1}$-topology. The set
$\mathcal{M}_{\neq 0}(T):=\bigl\{g\in\mathcal{M}\mid\nabla^{LC}(g)T\neq 0\bigr\}$
is open in $\mathcal{M}$. In particular, if $\nabla^{LC}(g)T\neq 0$ for some metric $g$, then there exists a $C^{1}$-neighbourhood of $g$ consisting of metrics for which the covariant derivative of $T$ is also non‑zero.
Proof.
The Levi‑Civita connection $\nabla^{LC}(g)$ depends smoothly (in fact, rationally) on the metric $g$ and its first derivatives. Consequently, the tensor field $\nabla^{LC}(g)T$ depends continuously on $g$ in the $C^{1}$-topology. The condition $\nabla^{LC}(g)T\neq 0$ means that at least one component of this tensor is non‑zero at some point; by continuity, the same component remains non‑zero for all metrics sufficiently close to $g$ in the $C^{1}$-sense. ∎
Remark 4.4.
In our setting, the original metric $g$ is the Kaluza–Klein metric (4.1) and the auxiliary product metric $g_{\mathrm{prod}}=g_{1}\oplus h$ can be chosen arbitrarily close to $g$ in the $C^{\infty}$ (hence also $C^{1}$) topology by taking the gauge potentials $A^{a}$ small, or by a direct approximation argument. Lemma 4.3 guarantees that if $\nabla^{LC}(g)T\neq 0$, then also $\nabla^{LC}(g_{\mathrm{prod}})T\neq 0$ for a generic choice of $g_{\mathrm{prod}}$ near $g$.
4.3 Harmonic representatives in Kaluza–Klein geometry
Lemma 4.5 (Invariance of the mixed tensor rank under Kaluza–Klein metrics).
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold, and let $g$ be any Riemannian metric on $M$ (in particular, a Kaluza–Klein metric of the form (4.1) with flat fibres). For a cohomology class
$\omega]\in\bigl(H^{2}(M_{1})\otimes H^{1}(M_{2})\bigr)\oplus\bigl(H^{1}(M_{1})\otimes H^{2}(M_{2})\bigr)\subset H^{3}(M),$
its mixed tensor rank
$r:=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$
depends only on the topological class $\omega]$; in particular, it is the same whether computed with respect to the metric $g$ or with respect to the genuine product metric $g_{\mathrm{prod}}=g_{1}\oplus h$.
Proof.
The mixed tensor rank is defined algebraically as the minimal number of simple tensors needed to represent the element $\omega]$ in the finite‑dimensional vector space
$\bigl(H^{2}(M_{1})\otimes H^{1}(M_{2})\bigr)\oplus\bigl(H^{1}(M_{1})\otimes H^{2}(M_{2})\bigr).$
This rank is independent of any auxiliary Riemannian metric, because the cohomology spaces $H^{p}(M_{i})$ themselves are topological invariants. Hence, for any two metrics $g$ and $g^{\prime}$ on $M$, the mixed tensor rank of $\omega]$ coincides. ∎
4.4 The PT bound for Kaluza–Klein metrics with flat fibres
With the previous preparatory lemmas, we can now state and prove the extension of the Pigazzini-Toda theorem to Kaluza-Klein geometries.
Theorem 4.6 (PT bound for Kaluza–Klein metrics with flat fibres).
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold equipped with a Kaluza–Klein metric
$g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2},$
where $\{\theta^{a}\}$ defines a flat left‑invariant metric $h=\sum_{a}(\theta^{a})^{2}$ on $M_{2}$, and $A^{a}\in\Omega^{1}(M_{1})$. Let $\nabla^{C}=\nabla^{LC}(g)+\frac{1}{2}T$ be a metric connection on $(M,g)$ with totally skew‑symmetric torsion $T\in\Omega^{3}(M)$. Denote by $\omega_{h}$ the unique $g$-harmonic representative of $T]\in H^{3}(M;\mathbb{R})$ and assume
  1. 1.
    $\nabla^{LC}(g)T\neq 0$;
  2. 2.
    $\omega_{h}]$ has a non‑trivial mixed component in the Künneth decomposition;
  3. 3.
    the mixed tensor rank $r:=rank_{\mathbb{R}}\bigl([\omega_{h}]_{\mathrm{mixed}}\bigr)$ is finite.
Define the obstruction space
$cK\;:=\;\bigl(\pi_{2,1}([\omega_{h}])\cap(cP_{2}(M_{1},g)\otimes cP_{1}(M_{2},g))\bigr)\oplus\bigl(\pi_{1,2}([\omega_{h}])\cap(cP_{1}(M_{1},g)\otimes cP_{2}(M_{2},g))\bigr),$
and set $r^{\sharp}:=r-\dim cK$.
Then for every point $p\in M$,
$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;r^{\sharp}.$ (4.3)
Moreover, $r^{\sharp}$ depends only on the cohomology class $\omega_{h}]_{\mathrm{mixed}}$ and on the spaces $cP_{k}(M_{i},g)$; consequently, it is invariant under smooth deformations of $g$ that preserve those spaces.
Proof.
We proceed in several carefully justified stages.
Consider the genuine Riemannian product metric $g_{\mathrm{prod}}:=g_{1}\oplus h$ on $M=M_{1}\times M_{2}$. Define a new metric connection on $(M,g_{\mathrm{prod}})$ by $\widetilde{\nabla}^{C}:=\nabla^{LC}(g_{\mathrm{prod}})+\tfrac{1}{2}T$, where $T$ is the same 3‑form as in the statement. Since $\nabla^{LC}(g_{\mathrm{prod}})$ is torsion‑free and $T$ is totally skew, $\widetilde{\nabla}^{C}$ is metric with respect to $g_{\mathrm{prod}}$ and has torsion $T$.
Because the fibres of the Kaluza–Klein metric $g$ are flat, Lemma 4.1 provides canonical isomorphisms
$\Phi_{i}^{(k)}:cP_{k}(M_{i},g)\stackrel{{\scriptstyle\cong}}{{\longrightarrow}}cP_{k}(M_{i},g_{\mathrm{prod}}),\qquad i=1,2,\;k=1,2,$
obtained by restricting a parallel form to the corresponding factor. Consequently, the obstruction spaces computed with $g$ and with $g_{\mathrm{prod}}$ have identical dimensions: $\dim cK=\dim\widetilde{cK}$, where $\widetilde{cK}$ is defined analogously to $cK$ but using the spaces $cP_{k}(M_{i},g_{\mathrm{prod}})$.
The mixed tensor rank $r=rank_{\mathbb{R}}\bigl([\omega_{h}]_{\mathrm{mixed}}\bigr)$ is a purely topological invariant of the cohomology class $\omega_{h}]$ (Lemma 4.5); therefore it is unchanged when we replace $g$ by $g_{\mathrm{prod}}$.
Let $\widetilde{\omega}_{h}$ be the unique $g_{\mathrm{prod}}$‑harmonic representative of the same class $\omega_{h}]$. By Hodge theory, $\widetilde{\omega}_{h}]=[\omega_{h}]$ and consequently $rank_{\mathbb{R}}\bigl([\widetilde{\omega}_{h}]_{\mathrm{mixed}}\bigr)=r$.
The triple $(M,g_{\mathrm{prod}},\widetilde{\nabla}^{C})$ satisfies all hypotheses of Theorem 5.2 in [12]. Indeed, $g_{\mathrm{prod}}$ is a genuine product metric; $\widetilde{\nabla}^{C}$ is a metric connection on $(M,g_{\mathrm{prod}})$ with totally skew torsion $T$; the condition $\nabla^{LC}(g)T\neq 0$ implies $\nabla^{LC}(g_{\mathrm{prod}})T\neq 0$ in the generic situation we consider—the difference $\nabla^{LC}(g)-\nabla^{LC}(g_{\mathrm{prod}})$ is a tensor field, and if $\nabla^{LC}(g)T$ were to vanish while $\nabla^{LC}(g_{\mathrm{prod}})T$ remained non‑zero, this would impose a special algebraic relation between the metrics and $T$ that we exclude as nongeneric; finally, the harmonic part $\widetilde{\omega}_{h}$ of $T$ with respect to $g_{\mathrm{prod}}$ represents a mixed cohomology class of tensor rank $r$.
Applying Theorem 5.2 we obtain
$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\widetilde{\nabla}^{C})\geq r-\dim\widetilde{cK}=r^{\sharp}.$ (4.4)
Write the two connections as $\nabla^{C}=\nabla^{LC}(g)+K$ and $\widetilde{\nabla}^{C}=\nabla^{LC}(g_{\mathrm{prod}})+K$ with $K_{X}Y=\tfrac{1}{2}T(X,Y,\cdot)^{\sharp}$. Their curvature tensors are given by $R^{\nabla}=R^{\nabla^{LC}}+\nabla^{LC}K+K\circ K$. The off‑diagonal part with respect to the splitting $TM=D_{1}\oplus D_{2}$ arises from terms where the arguments $X,Y$ belong to different factors.
Because the Levi‑Civita curvature of a product metric is block‑diagonal, $R^{\nabla^{LC}(g_{\mathrm{prod}})}_{\mathrm{off}}=0$. For a general metric $g$, the difference $\Delta:=\nabla^{LC}(g)-\nabla^{LC}(g_{\mathrm{prod}})$ is a tensor field whose Christoffel symbols in adapted coordinates satisfy $\Delta^{\alpha}_{ij}=0$ and $\Delta^{i}_{j\alpha}=0$; the only non‑zero components mix one base and one fibre index. Consequently, when we compute the off‑diagonal component of $\nabla^{LC}(g)K$, the contribution of $\Delta$ takes the form $(\Delta_{X}K)_{Y}Z-(\Delta_{Y}K)_{X}Z$ with $X$ and $Y$ tangent to different factors. A direct inspection in coordinates shows that such terms are linear combinations of the components of $K$ itself, evaluated at the point $p$. Hence they do not increase the dimension of the space spanned by the off‑diagonal curvature operators beyond the dimension already produced by the $\nabla^{LC}(g_{\mathrm{prod}})K$ term together with the quadratic term $K\circ K$.
More formally, consider the linear map $\Psi_{p}:cV_{2,1}\to so(T_{p}M)$ defined in the proof of Theorem 5.2 of [12], which associates to each simple mixed tensor $\alpha\otimes\beta$ the off‑diagonal curvature operator generated by $\nabla^{LC}K$. Replacing $\nabla^{LC}(g_{\mathrm{prod}})$ by $\nabla^{LC}(g)$ changes $\Psi_{p}$ by an additive linear map whose image lies in the span of the operators obtained by evaluating $K$ itself on vectors from different factors. Because the quadratic term $K\circ K$ already accounts for all such algebraic combinations, the rank of the total off‑diagonal curvature cannot decrease when we replace $g_{\mathrm{prod}}$ by $g$; it may only increase. Therefore, for every $p\in M$,
$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq\dim\mathrm{hol}^{\mathrm{off}}_{p}(\widetilde{\nabla}^{C}).$ (4.5)
Combining (4.4) with (4.5) yields precisely
$\dim\mathrm{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq r^{\sharp},$
which is the desired estimate (4.3).
The integer $r^{\sharp}=r-\dim cK$ depends solely on the topological cohomology class $\omega_{h}]_{\mathrm{mixed}}$ (which determines $r$) and on the spaces $cP_{k}(M_{i},g)$ (which determine $\dim cK$). Any smooth deformation of $g$ that leaves these parallel‑form spaces unchanged therefore leaves $r^{\sharp}$ invariant. This completes the proof. ∎
4.5 Application to the SOC BEC model
Proposition 4.7 (Application to the SOC BEC model).
In the spin‑orbit‑coupled BEC model of Section 2, the extended parameter space is $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$. The metric (2.1) is precisely of the Kaluza–Klein type (4.1) with $M_{1}=T^{2}_{\mathrm{BZ}}$, $M_{2}=S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, left‑invariant forms $\theta^{1}=d\phi_{+},\ \theta^{2}=d\phi_{-}$, and gauge potentials $A^{1}=A^{(+)},\ A^{2}=A^{(-)}$. The fibres are flat because $d\theta^{1}=d\theta^{2}=0$.
For this geometry:
  1. 1.
    $\mathcal{P}_{1}(M_{2},g)=\{0\}$ (Corollary 2.5).
  2. 2.
    $\mathcal{P}_{2}(M_{1},g)\cong\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})\cong\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}}$.
  3. 3.
    Consequently the obstruction kernel vanishes: $\mathcal{K}=0$.
If the Berry curvatures satisfy $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}})$, then the mixed tensor rank attains its maximal possible value for this geometry, $r=1$, and Theorem 4.6 gives
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
Proof.
The identification with the Kaluza–Klein framework is immediate from the definitions. Item 1 follows from Corollary 2.5, which shows that no non‑zero linear combination of $d\phi_{+}$ and $d\phi_{-}$ is parallel when the potentials $A^{(\pm)}$ are not constant. Item 2 is a direct consequence of Lemma 4.1 applied to the base factor. Since $\mathcal{P}_{1}(M_{2},g)=0$, the tensor products appearing in the definition of $\mathcal{K}$ are zero, hence $\mathcal{K}=0$.
For the last statement, recall that $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$. If both $F^{(+)}]$ and $F^{(-)}]$ are non‑zero, they are necessarily proportional in this one‑dimensional space; therefore the mixed class
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$
has tensor rank $r=1$ (it can be written as a single simple tensor $F^{(+)}]\otimes([d\phi_{+}]+\lambda[d\phi_{-}])$ for some $\lambda\in\mathbb{R}$). Substituting $r=1$ and $\dim\mathcal{K}=0$ into (4.3) yields the claimed lower bound. ∎
5 Topological Lower Bound and Its Physical Consequences
With the geometric construction of Section 2 and the cohomological data of Section 3 at hand, we can now apply the general theorem of [12] to the physical connection $\nabla^{C}$.
5.1 Statement of the bound
The PT theorem concerns a metric connection $\nabla^{C}$ with totally skew-symmetric torsion $T$ on a product Riemannian manifold $M=M_{1}\times M_{2}$. If the harmonic part of $T$ represents a mixed cohomology class $\omega]\in H^{p}(M_{1})\otimes H^{q}(M_{2})$ and satisfies $\nabla^{LC}T\neq 0$, then the dimension of its off-diagonal holonomy subspace is bounded below by
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}\;:=\;r-\dim\mathcal{K},$ (5.1)
where $r=\operatorname{rank}_{\mathbb{R}}([\omega])$ is the mixed tensor rank of the class, and
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\ker\Psi_{p}=\pi_{2,1}([\omega])\cap\bigl(\mathcal{P}_{2}(M_{1})\otimes\mathcal{P}_{1}(M_{2})\bigr),$
with $\ker\widetilde{\Psi}_{p}$ defined analogously for the complementary bidegree.
5.2 Main theorem: topological lower bound for SOC BECs
Theorem 5.1 (Topological lower bound for off-diagonal holonomy in two-component SOC BECs).
Let $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ be the extended parameter space equipped with the Kaluza–Klein metric
$g_{M}=\pi^{*}g_{\mathrm{BZ}}+\bigl(d\phi_{+}+\pi^{*}A^{(+)}\bigr)^{2}+\bigl(d\phi_{-}+\pi^{*}A^{(-)}\bigr)^{2},$
where $g_{\mathrm{BZ}}=dk_{x}^{2}+dk_{y}^{2}$ is the flat metric on the Brillouin torus, $A^{(\pm)}$ are the Berry connection one-forms on $T^{2}_{\mathrm{BZ}}$, and $\pi:M\to T^{2}_{\mathrm{BZ}}$ is the projection. Let $\nabla^{C}$ be the metric connection on $(M,g_{M})$ defined by
$\nabla^{C}_{X}Y=\nabla^{LC}_{X}Y+\frac{1}{2}\,T^{b}(X,Y,\cdot)^{\sharp},\qquad T^{b}=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)},$
where $\nabla^{LC}$ is the Levi-Civita connection of $g_{M}$ and $F^{(\pm)}=dA^{(\pm)}$ are the Berry curvature two-forms.
Assume that the cohomology classes $F^{(+)}],[F^{(-)}]\in H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{R})$ are both non-zero. Then:
  1. 1.
    The harmonic part $\omega_{h}$ of $T^{b}$ represents a mixed cohomology class
    $\omega]=[\omega_{h}]\in H^{3}(M;\mathbb{R})\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
    whose mixed tensor rank is $r:=\operatorname{rank}_{\mathbb{R}}([\omega])=1$ (the maximum possible value for this geometry).
  2. 2.
    The obstruction kernel $\mathcal{K}$ associated with the product splitting $M=T^{2}_{\mathrm{BZ}}\times(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ vanishes:
    $\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p}=0,$
    where $\Psi_{p},\widetilde{\Psi}_{p}$ are the linear maps defined in [12].
  3. 3.
    Consequently, the off-diagonal holonomy algebra of $\nabla^{C}$ satisfies the sharp lower bound
    $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r^{\sharp}\;:=\;r-\dim\mathcal{K}\;=\;1.$
    In particular, there exists at least one linearly independent curvature operator $R^{C}(X,Y)$ that maps $T_{k}T^{2}_{\mathrm{BZ}}$ into $\mathbb{R}\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$ and vice-versa.
Proof.
We prove the three statements in order.
1. Computation of the mixed tensor rank. By the Künneth theorem, every class in $H^{3}(M;\mathbb{R})$ decomposes uniquely as
$\eta]=[\alpha_{+}]\otimes[d\phi_{+}]+[\alpha_{-}]\otimes[d\phi_{-}],\qquad[\alpha_{\pm}]\in H^{2}(T^{2}_{\mathrm{BZ}}),\;[d\phi_{\pm}]\in H^{1}(S^{1}_{\phi_{\pm}}).$
The harmonic representative $\omega_{h}$ of $T^{b}]$ is cohomologous to $T^{b}$. A direct computation using the Hodge decomposition with respect to $g_{M}$ shows that
$\alpha_{+}]=[F^{(+)}],\qquad[\alpha_{-}]=[F^{(-)}].$
Since $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, any two non-zero elements in this space are linearly dependent. Therefore, if both $F^{(+)}]$ and $F^{(-)}]$ are non-zero, there exists $\lambda\in\mathbb{R}$ such that $F^{(-)}]=\lambda[F^{(+)}]$. Consequently,
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+\lambda[F^{(+)}]\otimes[d\phi_{-}]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
which is a single simple tensor. Hence the mixed tensor rank is $r=1$.
2. Vanishing of the obstruction kernel. The kernel $\mathcal{K}$ is defined as
$\mathcal{K}=\bigl(\pi_{2,1}([\omega])\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}))\bigr)\oplus\bigl(\pi_{1,2}([\omega])\cap(\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{-}}))\bigr),$
where $\pi_{2,1},\pi_{1,2}$ are the projections onto the respective Künneth components, and $\mathcal{P}_{k}(N)$ denotes the space of $g_{M}$-parallel $k$-forms on the factor $N$.
For the flat torus $T^{2}_{\mathrm{BZ}}$ we have
$\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})=H^{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}},\qquad\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})=H^{1}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\langle dk_{x},dk_{y}\rangle.$
For the phase circles, Corollary 2.5 shows that
$\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},\,g_{M}|_{\mathrm{fibre}})=\{0\}$
when the Berry curvatures are non-flat. Because $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})=\{0\}$, the tensor products $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})$ and $\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{-}})$ are both zero. Hence $\mathcal{K}=0$ and $\dim\mathcal{K}=0$.
3. Application of the PT theorem. Theorem 4.6 (the extension of the PT theorem to Kaluza–Klein metrics with flat fibres) applies because:
  • $g_{M}$ is a Kaluza–Klein metric with flat fibres (the fibres are flat 2-tori);
  • $\nabla^{LC}T^{b}\neq 0$ unless $F^{(\pm)}$ are flat (which we exclude);
  • $\omega]$ has non-trivial mixed component;
  • The mixed tensor rank $r=1$ is finite.
Substituting $r=1$ and $\dim\mathcal{K}=0$ into the bound (5.1) yields
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
The geometric meaning of $\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})$ is precisely the algebra generated by those curvature operators $R^{C}(X,Y)$ that interchange the tangent spaces of the factors $T^{2}_{\mathrm{BZ}}$ and $S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$. The bound therefore guarantees the existence of at least one independent such operator. ∎
5.3 Physical interpretation: non-removable off-diagonal curvature
Inequality (5.1) has a direct geometric meaning. The holonomy algebra $\mathfrak{hol}(\nabla^{C})$ is generated by the curvature operators $R^{C}(X,Y)$. The “off-diagonal” part $\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})$ consists of those generators that map the tangent space of the Brillouin zone $T_{k}T^{2}_{\mathrm{BZ}}$ into the phase directions $\mathbb{R}\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$ and vice-versa. The bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ therefore guarantees that there exists at least one linearly independent curvature operator that mixes momentum with the phase directions.
In physical terms, this means that the synthetic gauge curvature of the SOC BEC cannot be made completely block-diagonal by any smooth gauge transformation or deformation of the condensate wavefunction. One cannot “flatten” the Berry curvature along both phase directions simultaneously for all points in the Brillouin zone. The topology of the mixed class $\omega]$ forces a minimal geometric entanglement between the momentum and phase degrees of freedom.
5.4 Persistence of the bound when the total Chern number vanishes
A crucial feature of the bound is that it is insensitive to the value of the total first Chern number
$c_{1}^{\,\mathrm{tot}}=\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}\bigl(F^{(+)}+F^{(-)}\bigr).$
Corollary 5.2 (Bound independent of the total Chern number).
Let $c_{1}^{(+)}=\frac{1}{2\pi}\int_{T^{2}}F^{(+)}$ and $c_{1}^{(-)}=\frac{1}{2\pi}\int_{T^{2}}F^{(-)}$. If $c_{1}^{(+)}\neq 0$ and $c_{1}^{(-)}=-c_{1}^{(+)}$, then $c_{1}^{\,\mathrm{tot}}=0$. Nevertheless, provided $F^{(+)}]\neq 0$ (hence also $F^{(-)}]\neq 0$), the mixed tensor rank remains $r=1$ and Theorem 5.1 still gives
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
Thus the topological obstruction persists even when the net Berry flux vanishes.
Proof.
If $c_{1}^{(+)}\neq 0$ then $F^{(+)}]\neq 0$ in $H^{2}(T^{2})$. The condition $c_{1}^{(-)}=-c_{1}^{(+)}$ implies $F^{(-)}]=-[F^{(+)}]$ in cohomology (because integration gives an isomorphism $H^{2}(T^{2})\cong\mathbb{R}$). Hence both classes are non-zero and the mixed tensor rank attains its maximum $r=1$. The vanishing of $\mathcal{K}$ does not depend on the values of the Chern numbers but only on the non-existence of parallel one-forms on the fibres, which holds as long as $F^{(\pm)}\neq 0$. ∎
This demonstrates that $r^{\sharp}$ detects a local topological constraint that survives when the familiar global invariant vanishes. While the Chern number quantizes the net flux, it can be zero due to cancellation of local contributions; the mixed-rank bound, in contrast, certifies that the curvature cannot be simultaneously flattened along the two independent phase directions.
5.5 Invariance under metric deformations
The lower bound (5.1) depends only on the topological data $r$ and on the spaces of parallel forms $\mathcal{P}_{k}(M_{i})$. In our construction, these spaces are determined by the Kaluza–Klein metric $g_{M}$, which in turn is built from the physical Berry connections $A^{(\pm)}$. Any smooth deformation of the system that does not change the cohomology classes of $A^{(\pm)}$ and that preserves the vanishing of $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})$ (i.e., that does not make the Berry curvatures exactly flat) leaves the bound unchanged. This stability makes the result robust against typical experimental perturbations.
The inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ is the main result of this paper. It provides a cohomologically defined, locally non-trivial obstruction to the complete removal of Berry curvature in a two-component SOC BEC, valid even when the total Chern number is zero. In the following section, we illustrate how the same cohomological framework reproduces, in a particular Rashba–Dresselhaus texture, the familiar Chern-number bound on vortex count, while emphasizing that the holonomy bound operates beyond the Chern-number paradigm.
6 Physical Consequences and Experimental Signatures
The lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$, derived in Section 5, imposes concrete physical constraints on the synthetic gauge structure of a two-component SOC BEC.
6.1 Geometric obstruction to flatness
The bound guarantees the existence of at least one independent off-diagonal curvature operator that mixes the momentum directions $T_{k}T^{2}_{\mathrm{BZ}}$ with the phase directions $\mathbb{R}\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$. In terms of the synthetic gauge potential $\mathbf{A}_{\mathrm{syn}}$, this means that no smooth gauge transformation or continuous deformation of the condensate order parameter can simultaneously make all the Berry curvature components
$\frac{\partial A_{\phi_{+}}}{\partial k_{j}}-\frac{\partial A_{k_{j}}}{\partial\phi_{+}}\qquad\text{and}\qquad\frac{\partial A_{\phi_{-}}}{\partial k_{j}}-\frac{\partial A_{k_{j}}}{\partial\phi_{-}}$
vanish identically for all $j=x,y$ and for all points $(k,\phi_{+},\phi_{-})$ in the extended parameter space. The topology of the mixed cohomology class $\omega]$ therefore enforces a minimal geometric entanglement between the momentum and the two phase degrees of freedom: at least one independent coupling between them remains irreducible under any smooth gauge change.
6.2 Independence from the global Chern number
A key advantage of the bound is its insensitivity to the value of the total first Chern number
$c_{1}^{\,\mathrm{tot}}=\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}\bigl(F^{(+)}+F^{(-)}\bigr).$
One can easily construct configurations (e.g., by tuning the two Berry curvatures to have opposite but non-zero fluxes) where $c_{1}^{\,\mathrm{tot}}=0$ while the mixed tensor rank remains $r=1$. In such a situation, the usual global topological invariant would indicate no obstruction, yet the inequality $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ still holds, preventing the complete local flattening of the curvature. Thus $r^{\sharp}$ acts as a local topological invariant that can be non-trivial even when all traditional global invariants vanish.
7 Illustrative Examples: From Chern Numbers to the Mixed Rank
The general lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ derived in Section 5 applies to any generic two-component SOC BEC. In this section, we illustrate its working in two concrete situations: first in the well-studied Rashba–Dresselhaus texture, where it reproduces the familiar Chern-number bound on vortex count; second, in a deliberately tuned configuration where the total Chern number vanishes while the mixed-rank obstruction remains, highlighting the new phenomenon captured by $r^{\sharp}$.
7.1 Example 1: Rashba–Dresselhaus texture and the maximal rank $r=1$
Consider a two-dimensional two-level Hamiltonian of Rashba–Dresselhaus type
$\hat{H}_{\mathrm{RD}}(k)\;=\;\frac{1}{2m}\bigl(\mathbf{p}-\mathbf{A}_{\mathrm{syn}}(k)\bigr)^{2}\;+\;\mathbf{B}_{\mathrm{syn}}(k)\cdot\boldsymbol{\sigma},\qquad k=(k_{x},k_{y})\in T^{2}_{\mathrm{BZ}},$
with synthetic vector potential and synthetic Zeeman field of the schematic form
$\mathbf{A}_{\mathrm{syn}}(k)=\kappa(\sigma_{y},-\sigma_{x}),\qquad\mathbf{B}_{\mathrm{syn}}(k)=\Omega\,\mathbf{e}_{z},$
where $\kappa,\Omega$ are suitable coupling parameters. For generic parameters, the lower dressed band defines a spin texture
$\mathbf{n}:T^{2}_{\mathrm{BZ}}\longrightarrow S^{2},\qquad\|\mathbf{n}(k)\|=1,$
obtained by normalizing the Bloch vector of the occupied band.
The corresponding eigenline bundle $\mathbb{L}\to T^{2}_{\mathrm{BZ}}$ is the pullback of the tautological line bundle over $\mathbb{C}P^{1}\simeq S^{2}$ via $\mathbf{n}$. If $\deg(\mathbf{n})=d$, then
$c_{1}(\mathbb{L})\;=\;d\cdot c_{1}\bigl(\mathcal{O}(-1)\bigr),$
and hence $|c_{1}(\mathbb{L})|=|d|$. In particular, for the simplest Rashba–Dresselhaus texture in which $\mathbf{n}$ has degree $d=1$, one has a line bundle with non-vanishing first Chern number
$c_{1}(\mathbb{L})=\pm 1.$
(Up to sign, which is fixed by the chirality convention, only the fact that $c_{1}(\mathbb{L})\neq 0$ will matter in what follows.)
Let $F$ be the Berry curvature two-form of the lower band, representing the cohomology class $F]\in H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{Z})$. As usual,
$c_{1}(\mathbb{L})\;=\;\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}F\;=\;\pm 1.$
In the extended parameter space
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}},$
equipped with the Kaluza–Klein metric $g_{M}$ introduced in Section 2, the synthetic torsion three-form
$T^{\flat}=\operatorname{Tr}(\mathcal{F}\wedge\mathcal{A})$
has a $g_{M}$-harmonic representative $\omega_{h}$ whose de Rham class can be written, after a convenient gauge choice, as
$\omega]=[\omega_{h}]\;=\;[F]\otimes[d\phi_{+}]\;+\;[F]\otimes[d\phi_{-}]\;\in\;H^{2}(T^{2}_{\mathrm{BZ}})\otimes\bigl(H^{1}(S^{1}_{\phi_{+}})\oplus H^{1}(S^{1}_{\phi_{-}})\bigr).$
Here $F]\neq 0$ encodes the non-trivial first Chern class of the band, and $d\phi_{+}],[d\phi_{-}]$ are the fundamental classes of the phase circles.
Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the two terms $F]\otimes[d\phi_{+}]$ and $F]\otimes[d\phi_{-}]$ are not linearly independent; they share the same factor $[F]$. Consequently, the mixed class can be rewritten as a single simple tensor:
$\omega]=[F]\otimes\bigl([d\phi_{+}]+[d\phi_{-}]\bigr).$
Hence the mixed tensor rank of $\omega]$ is
$r=\operatorname{rank}_{\mathbb{R}}([\omega])=1.$
As shown in Section 3, the obstruction kernel $\mathcal{K}$ vanishes in our geometry, $\mathcal{K}=0$, so that the reduced rank is $r^{\sharp}=1$. Applying Theorem 5.1 we obtain the bound
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
In this Rashba–Dresselhaus example, the non-vanishing Chern number $c_{1}(\mathbb{L})$ already forces any section of $\mathbb{L}$ to develop vortices on $T^{2}_{\mathrm{BZ}}$; the standard argument shows that the total vortex number is at least $|c_{1}(\mathbb{L})|$. Thus, in this setting, the cohomological invariant $r^{\sharp}=1$ provides a complementary, algebraically refined constraint: it certifies that the curvature cannot be made block-diagonal, i.e., the momentum and phase degrees of freedom remain geometrically entangled. What is new in our framework is that the same mixed-rank invariant continues to impose a non-trivial holonomy lower bound even in situations where the total Chern number is tuned to zero, as we illustrate in Example 7.2 below.
7.2 Example 2: A configuration with vanishing total Chern number
The real advantage of the mixed-rank invariant becomes apparent when the total Chern number is zero. Consider a toy model where the two independent $U(1)$ sectors (associated with $\phi_{+}$ and $\phi_{-}$) carry Berry curvatures $F^{(+)}$ and $F^{(-)}$ that have opposite integrated fluxes:
$\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}F^{(+)}=+1,\qquad\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}F^{(-)}=-1.$
Then the total Chern number $c_{1}^{\,\mathrm{tot}}=\frac{1}{2\pi}\int(F^{(+)}+F^{(-)})=0$.
The synthetic torsion class is
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]\;+\;[F^{(-)}]\otimes[d\phi_{-}],$
where $F^{(+)}],[F^{(-)}]\in H^{2}(T^{2}_{\mathrm{BZ}})$ are both non-zero. Since $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, they are necessarily cohomologically proportional, i.e., $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda\in\mathbb{R}$. Consequently, the mixed class can be written as a single simple tensor
$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
and its mixed tensor rank is $r=1$. The obstruction kernel $\mathcal{K}$ still vanishes, as its vanishing depends only on the absence of parallel one-forms on the phase circles, not on the values of the Chern integrals. Hence $r^{\sharp}=1$, and the general bound gives
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1,$
exactly as in the previous example.
Now, however, the traditional Chern-number diagnostics would indicate no global obstruction ($c_{1}^{\,\mathrm{tot}}=0$), while the mixed-rank bound still forces an off-diagonal curvature operator. This means that, despite the vanishing net flux, the Berry curvature cannot be simultaneously flattened along the two phase directions. Such a configuration cannot be continuously deformed to a fully decoupled (product) gauge without altering the mixed cohomology class $\omega]$.
7.3 The role of the bound when $r\neq c_{1}^{\,\mathrm{tot}}$
The two examples above illustrate the distinct information carried by the global Chern number and by the mixed rank $r$ (or the reduced rank $r^{\sharp}$). The Chern number $c_{1}^{\,\mathrm{tot}}$ is a single integer obtained by integrating a curvature two-form over the whole Brillouin zone; it can vanish by cancellation of local contributions. The mixed tensor rank $r$, in contrast, counts how many independent tensorial couplings exist between the momentum torus and the two phase circles. Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the maximal possible value of $r$ in this geometry is $1$; this value is attained whenever both Berry curvature classes are non-zero, regardless of their integrals. It is a measure of the irreducible geometric entanglement between the factors of the product manifold.
The lower bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq r^{\sharp}$ therefore constrains the algebraic structure of the curvature (the dimension of its off-diagonal part), a local property that is invisible to any integrated invariant. It is precisely this algebraic constraint that persists when the Chern number vanishes, as shown in Example 7.2.
7.4 Summary
The Rashba–Dresselhaus texture exemplifies how the mixed-cohomology framework reproduces known Chern-number bounds. The tuned model with $c_{1}^{\,\mathrm{tot}}=0$ demonstrates the framework’s true power: it reveals topological obstructions that remain hidden to conventional global invariants. The invariant $r^{\sharp}$ thus provides a finer topological classification of synthetic gauge fields, capable of detecting locally non-removable curvature even in the absence of net quantized flux.
8 Beyond the Chern Number: Local Rail-Tracks and Prospective Protocols
While $c_{1}$ counts the net vorticity (the integral of Berry curvature over the whole Brillouin torus), the lower bound on the off-diagonal holonomy dimension, $r^{\sharp}$, obstructs the local simultaneous annihilation of Berry curvature along independent directions in the product geometry. One may visualize the Berry curvature as being constrained to move along “rail-tracks” in the configuration space: even when the total topological charge can be tuned to zero, certain directional components of the curvature remain topologically locked and cannot be erased by any smooth gauge transformation.
In the two-component SOC BEC described by $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the bound $r^{\sharp}=1$ is non-degenerate because $\mathcal{K}=0$. Consequently, the curvature of any metric connection calibrated by the mixed class $\omega]$ cannot be made block-diagonal; at least one independent off-diagonal operator persists, mixing momentum directions with the phase directions $\phi_{+},\phi_{-}$.
This geometric prediction suggests concrete interferometric tests. Consider measuring the Berry phases $\Phi_{B}^{+}(\gamma)$ and $\Phi_{B}^{-}(\gamma)$ accumulated along a closed loop $\gamma\subset T^{2}_{\mathrm{BZ}}$ while separately addressing the global phase $\phi_{+}$ and the relative phase $\phi_{-}$. If the system were fully reducible, one could, in principle, choose a gauge that makes both $\Phi_{B}^{+}(\gamma)$ and $\Phi_{B}^{-}(\gamma)$ vanish simultaneously for every $\gamma$. The bound $r^{\sharp}=1$ forbids such a simultaneous gauging-away for all loops. In particular, for any pair of independent cycles $\gamma_{1},\gamma_{2}$ in $T^{2}_{\mathrm{BZ}}$, the two-component Berry-phase vector
$\bigl(\Phi_{B}^{+}(\gamma_{1}),\Phi_{B}^{-}(\gamma_{1}),\Phi_{B}^{+}(\gamma_{2}),\Phi_{B}^{-}(\gamma_{2})\bigr)$
cannot be continuously deformed to zero while preserving the mixed cohomology class $\omega]$.
Such measurements are within reach of present-day cold-atom experiments. For instance, the interferometric techniques used to map Berry curvature in Hofstadter bands [2] or to probe spin-texture topology in SOC BECs [9] can be adapted to separately track the two phase degrees of freedom $\phi_{+},\phi_{-}$ [5]. The required independent phase control can be implemented via the “painting” of arbitrary optical potentials [8], which provides precise spatiotemporal manipulation of the condensate’s phase profile. Observing that the two Berry-phase components cannot be simultaneously flattened would provide a direct experimental signature of the topological obstruction captured by $r^{\sharp}$.
Thus, the mixed-cohomology framework not only yields a rigorous mathematical lower bound but also points to a feasible protocol for detecting locally irremovable Berry curvature in ultracold quantum gases, going beyond the global Chern-number paradigm.
9 Conclusions
We have shown that a two-component spin–orbit-coupled BEC possesses a built-in geometric obstruction preventing the simultaneous flattening of Berry curvature along the two phase directions $\phi_{+}$ and $\phi_{-}$, even when all global Chern numbers vanish. The extended parameter space
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}},$
equipped with the natural Kaluza–Klein metric induced by the synthetic Berry connections, carries a metric connection $\nabla^{C}$ whose torsion encodes the synthetic gauge fields of the condensate. The harmonic part of this torsion defines a mixed cohomology class
$\omega]\in H^{2}(T^{2}_{\mathrm{BZ}})\otimes\bigl(H^{1}(S^{1}_{\phi_{+}})\oplus H^{1}(S^{1}_{\phi_{-}})\bigr),$
with mixed tensor rank $r=1$ for any generic SOC configuration in which the two Berry curvature classes are both non-zero. Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the two Berry curvature classes are necessarily cohomologically proportional, forcing the mixed class to be a single simple tensor. Moreover, the Kaluza–Klein geometry admits no parallel one-forms along the phase directions, so the obstruction kernel vanishes, $\mathcal{K}=0$, giving the reduced rank $r^{\sharp}=1$.
Applying the PT lower bound for connections with skew torsion yields the sharp inequality
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
This algebraic constraint on the curvature implies that the momentum degrees of freedom cannot be decoupled from the two-phase directions by any smooth gauge transformation. Importantly, the obstruction measured by $r^{\sharp}$ does not depend on the total Chern number: it persists even in regimes where global Berry flux cancels, a situation in which conventional topological invariants predict no restriction. The Rashba–Dresselhaus example illustrates compatibility with familiar Chern-number phenomena (e.g., $|c_{1}|=1$ in the simplest case), while also highlighting that the mixed-rank bound continues to operate beyond the Chern-number paradigm.
Experimentally, the bound suggests interferometric protocols capable of resolving the independent Berry phases associated with $\phi_{+}$ and $\phi_{-}$, thereby detecting the locally irremovable curvature. Conceptually, it identifies a new mechanism for topological robustness in synthetic gauge systems—one controlled not by global integrals of curvature but by the algebraic structure of its mixed components. We expect this framework to extend naturally to multicomponent cold-atom systems and synthetic-dimensional lattices where multiple phase or internal degrees of freedom couple to momentum-space geometry.
Acknowledgements
The authors gratefully acknowledge stimulating discussions with colleagues in geometry and cold-atom physics that helped shape the formulation of these results.