### v2

# Local Topological Constraints on Berry Curvature in Spin–Orbit Coupled BECs

Abstract
We establish a local topological obstruction to flattening Berry curvature in spin-orbit-coupled Bose-Einstein condensates (SOC BECs), valid even when the global Chern number vanishes. For a generic two-component SOC BEC, the extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ carries a natural metric connection $\nabla^{C}$ whose torsion 3-form encodes the synthetic gauge fields. Its harmonic part defines a mixed cohomology class
$\omega]\in\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
whose mixed tensor rank equals one. Using a general geometric bound for metric connections with totally skew torsion on product manifolds, we show that the obstruction kernel $\mathcal{K}$ vanishes, yielding the sharp inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$. This forces at least one off-diagonal curvature operator, preventing complete gauging-away of Berry phases even when the total Chern number is zero. This provides the first cohomological lower bound certifying locally irremovable curvature in SOC BECs beyond the Chern-number paradigm.
Keywords: Berry curvature; spin–orbit-coupled Bose–Einstein condensates; synthetic gauge fields; holonomy with torsion; mixed cohomology; Kaluza–Klein geometry.
MSC (2020): 53C29, 53C80, 81V45, 82B10, 51P05, 55S20, 57R19, 57R56, 58A12.
1 Introduction
The study of spin–orbit-coupled (SOC) Bose–Einstein condensates (BECs) stands at a rich confluence of differential geometry, quantum many-body physics, and synthetic gauge theory. This interdisciplinary effort is part of the broader pursuit of artificial gauge fields in engineered quantum systems, spanning ultracold atoms, photonics, and solid-state platforms [1]. Advances in generating artificial gauge fields with Raman lasers have made these systems a versatile laboratory for exploring Berry curvature, synthetic magnetic fields, and non-Abelian textures with unprecedented control [6, 7, 14]. These studies are often conducted in toroidal traps, which can serve as magnetic storage rings for BECs [3], providing a well-controlled environment for synthetic gauge field experiments. Remarkable achievements include the observation of Dirac monopoles in synthetic magnetic fields [10], demonstrating the high degree of geometric control now attainable. The geometric phase underpinning these phenomena traces back to the foundational work of Simon [13], while the experimental realization of artificial gauge fields has been comprehensively reviewed in [4].
A persistent theoretical challenge is to identify local topological obstructions that constrain the curvature even when global invariants—such as the first Chern number—vanish. While the Chern number quantizes the net flux, its vanishing does not guarantee that the curvature can be locally flattened; geometric obstructions may prevent its simultaneous annihilation along independent directions in the parameter space.
This work addresses that challenge by bringing a recent geometric result into physical focus. In [12], a lower bound was established for the off-diagonal holonomy of metric connections with totally skew-symmetric torsion on product manifolds. The bound is expressed in terms of a mixed de Rham cohomology class $\omega]\in H^{p}(M_{1})\otimes H^{q}(M_{2})$ associated with the torsion; its mixed tensor rank $r$ counts the minimal number of simple tensors needed to represent the class, and after subtracting contributions from parallel forms (which do not generate off-diagonal curvature) one obtains a corrected integer $r^{\sharp}$ that bounds the dimension of the off-diagonal holonomy algebra:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim\mathcal{K}.$
Here we show how this abstract geometric theorem materializes in a concrete physical setting: a generic two-component SOC BEC confined in a quasi-two-dimensional toroidal trap. Such a system possesses three universal, experimentally accessible degrees of freedom: the two crystal momenta forming a Brillouin-zone torus $T^{2}_{\mathrm{BZ}}$, a global $U(1)$ phase $\phi_{+}$ associated with particle-number conservation, and a relative $U(1)$ phase $\phi_{-}$ between the two internal spin components [9, 11]. We therefore take as the extended parameter space the manifold
$M=T_{BZ}^{2}\times S_{\phi_{+}}^{1}\times S_{\phi_{-}}^{1}.$
We emphasize that treating $M$ as a global product manifold requires the underlying principal bundle to be trivial, which holds precisely in the physical regime of vanishing total Chern number ($c_{\text{tot}}=0$) that is central to our main result. In configurations with non-zero net topological charge, this product structure should be understood as a local description in a gauge patch, sufficient for deriving local curvature bounds.
The first main task of the paper is to construct, in this space, a natural geometric structure to which the theorem of [12] applies. We introduce a metric $g_{M}$ of Kaluza–Klein type, which couples the base metric on $T^{2}_{\mathrm{BZ}}$ to the $U(1)$-valued connection forms that encode the synthetic gauge fields of the SOC BEC. On $(M,g_{M})$ we define a metric connection $\nabla^{C}$ whose torsion three-form is precisely the physical three-form
$T^{\flat}=\operatorname{Tr}(\mathcal{F}\wedge\mathbf{A}_{\mathrm{syn}}),$
built from the synthetic Berry curvature $\mathcal{F}$ and connection $\mathbf{A}_{\mathrm{syn}}$. The harmonic part of $T^{\flat}$ with respect to $g_{M}$ then defines a mixed cohomology class
$\omega]\in H^{3}(M)\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr).$
The second main task is to compute the topological data that enters the general bound. We prove that, under the generic condition that the two nearly degenerate bands carry non-zero Berry curvature classes $F^{(+)}],[F^{(-)}]\in H^{2}(T^{2}_{\mathrm{BZ}})$, the mixed tensor rank of $\omega]$ attains its maximal possible value within this geometry. Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, any two non-zero classes in this space are cohomologically proportional. Consequently, the mixed class $\omega]$ can be written as a single simple tensor
$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
and its mixed tensor rank is therefore $r=1$. Moreover, owing to the specific form of the metric $g_{M}$, the spaces of parallel one-forms on the phase circles are trivial, which forces the obstruction kernel $\mathcal{K}$ to vanish. Consequently the reduced rank is $r^{\sharp}=1$, and the general theorem yields the sharp inequality
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
This lower bound has immediate physical consequences. It implies that the curvature of the synthetic gauge connection cannot be made block-diagonal with respect to the splitting between momentum and phase directions; at least one independent off-diagonal operator persists, mixing the Brillouin zone with the two phase directions. Crucially, the bound detects the topological coupling of the momentum torus to the two distinct $U(1)$ phases, and therefore remains non-trivial even when the total Chern number of the lowest band is tuned to zero. Thus $r^{\sharp}$ provides a cohomological certificate of locally irremovable Berry curvature, invisible to conventional global invariants. The result suggests interferometric protocols that separately probe the Berry phases associated with $\phi_{+}$ and $\phi_{-}$, offering a way to detect this obstruction even in regimes of vanishing net vorticity.
The paper is organized as follows. Section 2 constructs the extended parameter space $M$, the Kaluza–Klein metric $g_{M}$, and the metric connection $\nabla^{C}$ with torsion $T^{\flat}$. Section 3 analyses the cohomology of $M$, computes the mixed tensor rank $r$, and shows that $\mathcal{K}=0$ in the present geometry. Section 4 extends the Pigazzini–Toda theorem to Kaluza–Klein metrics, ensuring the bound applies to the present geometry. Section 5 applies the general theorem of [12] to obtain the lower bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ and discusses its physical interpretation. Section 6 elaborates the physical implications of the bound, including its persistence when the total Chern number vanishes and its robustness under metric deformations. Section 7 illustrates how the same cohomological data reproduce, in a particular Rashba–Dresselhaus texture, the familiar Chern-number bound on the vortex number, and demonstrates with a tuned example that the holonomy bound remains valid even when the Chern number vanishes. Section 8 outlines prospective interferometric tests for detecting the locally irremovable curvature and indicates directions for future work. Section 9 summarises the results and discusses broader implications for topological quantum matter.
2 Geometric Construction of the Extended Parameter Space
A generic two-component spin–orbit-coupled Bose–Einstein condensate (SOC BEC) confined in a quasi-two-dimensional toroidal trap is characterized by three independent, experimentally accessible degrees of freedom: the crystal momenta $(k_{x},k_{y})$ spanning a Brillouin zone, a global $U(1)$ phase $\phi_{+}$ associated with particle-number conservation, and a relative $U(1)$ phase $\phi_{-}$ between the two internal spin states [9, 11].
We therefore take as the extended parameter space the smooth, compact, oriented product manifold
$M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}.$
Our goal in this section is to endow $M$ with a Riemannian structure that naturally incorporates the synthetic gauge fields of the SOC BEC, and to define on it a metric connection whose torsion three-form coincides with the physical Berry-curvature three-form.
2.1 Kaluza–Klein metric on the extended space
Let $g_{\mathrm{BZ}}=dk_{x}^{2}+dk_{y}^{2}$ be the flat metric on the Brillouin-zone torus $T^{2}_{\mathrm{BZ}}$. On each phase circle we introduce the standard angular coordinate with period $2\pi$; the canonical $U(1)$-invariant metric on $S^{1}$ is $d\phi^{2}$. The synthetic gauge fields of the SOC BEC provide, for each of the two independent $U(1)$ directions, a Berry connection. Denote by $A^{(+)},A^{(-)}\in\Omega^{1}(T^{2}_{\mathrm{BZ}})$ the corresponding real one-forms (the components of the synthetic $\mathfrak{su}(2)$ connection along the two independent $U(1)$ generators). Their curvatures are the Berry curvature two-forms $F^{(\pm)}=dA^{(\pm)}$.
Following the Kaluza–Klein prescription we define a metric $g_{M}$ on $M$ that couples the base metric to the connection forms:
$g_{M}=\pi^{*}g_{\mathrm{BZ}}+\bigl(d\phi_{+}+\pi^{*}A^{(+)}\bigr)^{2}+\bigl(d\phi_{-}+\pi^{*}A^{(-)}\bigr)^{2},$ (2.1)
where $\pi:M\to T^{2}_{\mathrm{BZ}}$ is the projection onto the momentum coordinates. The one-forms
$\Theta^{(+)}:=d\phi_{+}+\pi^{*}A^{(+)},\qquad\Theta^{(-)}:=d\phi_{-}+\pi^{*}A^{(-)}$
are the horizontal lifts of the phase coordinates; they are globally defined on $M$ and are orthogonal to the base directions with respect to $g_{M}$.
Lemma 2.1 (Properties of the metric $g_{M}$).
The metric (2.1) satisfies the following.
  1. 1.
    $g_{M}$ is smooth, positive definite and makes $\pi:(M,g_{M})\to(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ a Riemannian submersion.
  2. 2.
    The fibres $\pi^{-1}(k)\cong S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ are flat $2$-tori with the induced metric $(d\phi_{+})^{2}+(d\phi_{-})^{2}$.
  3. 3.
    The horizontal distribution $\mathcal{H}=\ker(d\phi_{+}+\pi^{*}A^{(+)})\cap\ker(d\phi_{-}+\pi^{*}A^{(-)})$ is orthogonal to the vertical distribution $\mathcal{V}=\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$.
  4. 4.
    If $A^{(\pm)}=0$ then $g_{M}$ reduces to the trivial product metric $g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$.
Proof.
Smoothness follows from the smoothness of $g_{\mathrm{BZ}}$ and $A^{(\pm)}$. Positive definiteness is clear because each term in (2.1) is non‑negative and their sum vanishes only when all differentials vanish.
For any $k\in T^{2}_{\mathrm{BZ}}$ the restriction of $g_{M}$ to the fibre $\pi^{-1}(k)$ is $(d\phi_{+})^{2}+(d\phi_{-})^{2}$, independent of $k$; hence the fibres are isometric to a flat $2$-torus. The map $\pi$ is a Riemannian submersion because, by construction, $g_{M}$ restricted to $\mathcal{H}$ coincides with $\pi^{*}g_{\mathrm{BZ}}$.
Orthogonality $\mathcal{H}\perp\mathcal{V}$ follows directly from $g_{M}(\partial_{\phi_{\pm}},X)=0$ for every $X\in\mathcal{H}$, which is a consequence of the definition of $\Theta^{(\pm)}$.
The last statement is immediate from (2.1). ∎
Remark 2.2.
Equation (2.1) is physically natural: it measures distances in the extended parameter space by combining the kinetic contribution from momentum changes with the “phase cost” induced by the synthetic gauge potentials. If $A^{(\pm)}=0$, the metric reduces to the trivial product metric $g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$.
2.2 Metric connection with prescribed torsion
On the Riemannian manifold $(M,g_{M})$ we now construct a metric connection $\nabla^{C}$ whose torsion three-form is built from the physical gauge fields. Let $\nabla^{LC}$ be the Levi‑Civita connection of $g_{M}$.
Define the three‑form
$T^{b}:=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}\in\Omega^{3}(M),$ (2.2)
where $F^{(\pm)}\wedge A^{(\pm)}$ are pulled back to $M$ via $\pi$ (we keep the same notation for the pull‑backs). $T^{b}$ is the trace $\operatorname{Tr}(\mathcal{F}\wedge\mathcal{A})$ with respect to the Killing form on $\mathfrak{u}(1)_{+}\oplus\mathfrak{u}(1)_{-}$, where $\mathcal{A}=A^{(+)}\otimes t_{+}+A^{(-)}\otimes t_{-}$ and $\mathcal{F}=F^{(+)}\otimes t_{+}+F^{(-)}\otimes t_{-}$.
Introduce the (2,1)(2,1)‑tensor $K$ by
$K(X,Y,Z):=\frac{1}{2}\,T^{b}(X,Y,Z)\qquad(X,Y,Z\in\mathfrak{X}(M)).$
The connection
$\nabla^{C}_{X}Y:=\nabla^{LC}_{X}Y+K(X,Y,\cdot)^{\sharp}$
is by construction metric ($\nabla^{C}g_{M}=0$) and has totally skew‑symmetric torsion $T^{b}$. Indeed, for any vector fields $X,Y$,
$T^{b}(X,Y,\cdot)^{\sharp}=\nabla^{C}_{X}Y-\nabla^{C}_{Y}X-[X,Y],$
which is precisely the definition of the torsion tensor of $\nabla^{C}$.
Because $T^{b}$ is not parallel with respect to $\nabla^{LC}$ unless the gauge fields are flat, the connection $\nabla^{C}$ satisfies the non‑parallelism condition $\nabla^{LC}T^{b}\neq 0$ required for the applicability of the general theorem [12].
Remark 2.3.
The three‑form $T^{b}$ is not closed in general; its exterior derivative contains terms quadratic in the curvatures $F^{(\pm)}$. Its de Rham cohomology class $T^{b}]\in H^{3}(M;\mathbb{R})$, however, depends only on the harmonic part of $T^{b}$ with respect to $g_{M}$, as will be computed in Section 3.
2.3 Parallel forms in the Kaluza–Klein geometry
A key ingredient of the topological bound is the space of parallel forms on the factors. For the base torus with the flat metric $g_{\mathrm{BZ}}$ we have the well‑known result
$\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})=H^{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}},\qquad\mathrm{vol}_{\mathrm{BZ}}=dk_{x}\wedge dk_{y}.$
The situation for the phase circles is more subtle. On a circle endowed with the standard metric $d\phi^{2}$, the one‑form $d\phi$ is parallel, hence $\mathcal{P}_{1}(S^{1},d\phi^{2})=\mathbb{R}\,d\phi$. In the Kaluza–Klein metric (2.1) the geometry of the fibres is “twisted” by the connection forms $A^{(\pm)}$. The vector fields $\partial_{\phi_{+}},\partial_{\phi_{-}}$ are Killing fields of $g_{M}$, but their dual one‑forms are $\Theta^{(\pm)}$, not $d\phi_{\pm}$.
Lemma 2.4 (Non‑parallelism of $d\phi_{\pm}$).
Assume $A^{(\pm)}$ are not identically zero. Then the one‑forms $d\phi_{+},d\phi_{-}$ are not parallel with respect to $g_{M}$.
Proof.
We compute the covariant derivative $\nabla^{LC}d\phi_{+}$ in local coordinates $(k_{x},k_{y},\phi_{+},\phi_{-})$. The non‑zero Christoffel symbols involving the index $\phi_{+}$ are
$\Gamma^{\phi_{+}}_{k_{x}\phi_{+}}=\frac{1}{2}g^{\phi_{+}\rho}\bigl(\partial_{k_{x}}g_{\phi_{+}\rho}+\partial_{\phi_{+}}g_{k_{x}\rho}-\partial_{\rho}g_{k_{x}\phi_{+}}\bigr).$
Since $g_{k_{x}\phi_{+}}=A^{(+)}_{k_{x}}$ is independent of $\phi_{+}$, we have $\partial_{\phi_{+}}g_{k_{x}\phi_{+}}=0$. Moreover $g_{\phi_{+}\phi_{+}}=1+(A^{(+)}_{k_{x}})^{2}+(A^{(+)}_{k_{y}})^{2}$ and $g_{\phi_{+}\phi_{-}}=A^{(+)}_{k_{x}}A^{(-)}_{k_{x}}+A^{(+)}_{k_{y}}A^{(-)}_{k_{y}}$. Consequently,
[…]
which is non‑zero for generic non‑constant potentials $A^{(\pm)}$. Hence $(\nabla^{LC}_{\partial_{k_{x}}}d\phi_{+})(\partial_{\phi_{+}})=-\Gamma^{\phi_{+}}_{k_{x}\phi_{+}}\neq 0$, showing that $d\phi_{+}$ is not parallel. The same argument applies to $d\phi_{-}$. ∎
Corollary 2.5 (Trivial parallel‑form spaces on the fibres).
For a generic SOC BEC with non‑trivial Berry curvature $F^{(\pm)}\neq 0$,
$\mathcal{P}_{1}\bigl(S^{1}_{\phi_{\pm}},g_{M}|_{\mathrm{fibre}}\bigr)=\{0\}.$
More precisely, any fibre‑tangent 1‑form that is parallel with respect to the full metric $g_{M}$ must be of the form $\xi=c_{+}d\phi_{+}+c_{-}d\phi_{-}$ with constants $c_{+},c_{-}$. However, Lemma 2.4 shows that for non‑constant potentials $A^{(\pm)}$, $\nabla^{LC}d\phi_{\pm}$ has non‑vanishing components along the base directions; consequently no non‑zero linear combination can be parallel on the whole manifold $M$. If $A^{(\pm)}$ are constant but $F^{(\pm)}\neq 0$, they are not closed, hence still not harmonic and therefore not parallel (on a compact Riemannian manifold every parallel form is harmonic).
2.4 Summary of the geometric setup
We have constructed:
  1. 1.
    an extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$;
  2. 2.
    a Kaluza–Klein metric $g_{M}$ on $M$ given by (2.1), coupling the momentum directions to the two independent $U(1)$ phases via the synthetic Berry connections $A^{(\pm)}$;
  3. 3.
    a metric connection $\nabla^{C}$ on $(M,g_{M})$ whose torsion three‑form is the physical quantity $T^{b}=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}$.
This construction provides the precise Riemannian framework needed to apply the general theorem of [12]. In the following section we analyze the cohomological data that arise from this geometry.
3 Cohomology of the Extended Space and the Mixed Tensor Rank
With the geometric structure $(M,g_{M},\nabla^{C})$ established in Section 2, we now analyze the cohomological data that enter the general lower bound. Recall that the torsion three-form $T^{\flat}\in\Omega^{3}(M)$ is not necessarily closed, but its de Rham cohomology class $T^{\flat}]\in H^{3}(M;\mathbb{R})$ is well defined. Let $\omega_{h}$ be the unique $g_{M}$-harmonic representative of this class; then
$\omega]:=[\omega_{h}]=[T^{\flat}]\in H^{3}(M;\mathbb{R}).$
3.1 Künneth decomposition of $H^{3}(M)$
Since $M$ is a product $T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ and each factor is oriented and compact, the Künneth theorem gives an isomorphism
$H^{3}(M;\mathbb{R})\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
because $H^{3}(T^{2}_{\mathrm{BZ}})=H^{2}(S^{1}_{\phi_{\pm}})=0$ and $H^{1}(T^{2}_{\mathrm{BZ}})$ does not contribute to degree $3$. Consequently $\dim H^{3}(M)=2$, and every class $\eta]\in H^{3}(M)$ can be written uniquely as
$\eta]=[\alpha_{+}]\otimes[d\phi_{+}]+[\alpha_{-}]\otimes[d\phi_{-}],$ (3.1)
where $\alpha_{\pm}]\in H^{2}(T^{2}_{\mathrm{BZ}})$ and $d\phi_{\pm}]\in H^{1}(S^{1}_{\phi_{\pm}})$ are the fundamental cohomology classes of the factors.
For the harmonic representative $\omega_{h}$ of our torsion class, the Hodge decomposition with respect to $g_{M}$ yields
$\omega_{h}=\alpha_{+}\wedge d\phi_{+}+\alpha_{-}\wedge d\phi_{-}+d\eta,$
where $\alpha_{\pm}\in\mathcal{H}^{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ are $g_{\mathrm{BZ}}$-harmonic two-forms on the Brillouin torus, and $\eta\in\Omega^{2}(M)$ is a coexact two-form. The term $d\eta$ is exact and therefore orthogonal to harmonic forms in the $L^{2}$-inner product; it does not contribute to the cohomology class $\omega]=[\omega_{h}]$.
Remark 3.1.
The representatives $d\phi_{\pm}$ are harmonic with respect to the standard metric on the circles, but they are not the harmonic representatives of the fundamental classes with respect to the full Kaluza–Klein metric $g_{M}$; the $g_{M}$-harmonic representatives are the horizontal forms $\Theta^{(\pm)}=d\phi_{\pm}+A^{(\pm)}$ introduced in (2.1). Because $d\phi_{\pm}]=[\Theta^{(\pm)}]$ in $H^{1}(S^{1}_{\phi_{\pm}})$, the cohomological decomposition (3.1) is unchanged.
3.2 Mixed tensor rank of the torsion class
The mixed tensor rank of a cohomology class $\eta]\in H^{3}(M)$ written as in (3.1) is defined as the minimal number of simple tensors required to represent it:
$\operatorname{rank}_{\mathbb{R}}([\eta]):=\min\Bigl\{m\;\Big|\;[\eta]=\sum_{i=1}^{m}[\beta_{i}]\otimes[\gamma_{i}],\;[\beta_{i}]\in H^{2}(T^{2}_{\mathrm{BZ}}),\;[\gamma_{i}]\in H^{1}(S^{1}_{\phi_{+}})\oplus H^{1}(S^{1}_{\phi_{-}})\Bigr\}.$
For the class $\omega]$ coming from the physical torsion, the two-forms $\alpha_{+},\alpha_{-}$ in (3.1) are precisely the (cohomology classes of the) Berry curvature two-forms $F^{(+)},F^{(-)}$ of the two independent $U(1)$ sectors. A key observation is that $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$. Therefore, any two non-zero classes in this space are necessarily cohomologically proportional. In a generic two-component SOC BEC, both Berry curvature classes $F^{(+)}]$ and $F^{(-)}]$ are non-zero; consequently, there exists a non-zero constant $\lambda\in\mathbb{R}$ such that
$F^{(-)}]=\lambda[F^{(+)}]\quad\text{in }H^{2}(T^{2}_{\mathrm{BZ}}).$
Substituting this relation into (3.1) gives
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+\lambda[F^{(+)}]\otimes[d\phi_{-}]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
which is a single simple tensor. Hence the mixed tensor rank attains the value
$r=1.$
Remark 3.2.
Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the maximal possible mixed tensor rank for a class of the form $\alpha_{+}]\otimes[d\phi_{+}]+[\alpha_{-}]\otimes[d\phi_{-}]$ with both $\alpha_{\pm}]\neq 0$ is $1$. All experimentally prominent SOC geometries (Rashba–Dresselhaus, Hofstadter, Raman-induced couplings) satisfy the condition $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$, thereby realizing the generic case $r=1$.
Remark 3.3.
Since $M_{2}=S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ is a product, its first cohomology decomposes as $H^{1}(M_{2})\cong H^{1}(S^{1}_{+})\oplus H^{1}(S^{1}_{-})$. Therefore an element of $H^{2}(M_{1})\otimes H^{1}(M_{2})$ can be written as a sum of terms of the form $\alpha]\otimes[d\phi_{+}]+[\beta]\otimes[d\phi_{-}]$. The mixed tensor rank $r$ is then the minimal number of simple tensors needed to represent this sum in the space $H^{2}(M_{1})\otimes\bigl(H^{1}(S^{1}_{+})\oplus H^{1}(S^{1}_{-})\bigr)$.
3.3 Vanishing of the obstruction kernel $\mathcal{K}$
The general bound of [12] involves a correction term $\dim\mathcal{K}$, where
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\ker\Psi_{p}=\pi_{2,1}([\omega])\cap\bigl(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})\bigr),$
and $\ker\widetilde{\Psi}_{p}$ is defined analogously with the roles of the two circles exchanged. Here $\pi_{2,1}$ denotes the projection onto the $(2,1)$-Künneth component.
In our geometry, we have $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}}$ (the space of parallel two-forms on the flat torus). The crucial observation, already made in Section 2.3, is that for the Kaluza–Klein metric $g_{M}$ the spaces of parallel one-forms on the phase circles are trivial:
$\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},\,g_{M}|_{\text{\rm fibre}})=\{0\},$
because any $g_{M}$-parallel one-form tangent to a fibre would have to be a constant multiple of $\Theta^{(\pm)}=d\phi_{\pm}+A^{(\pm)}$, and $\Theta^{(\pm)}$ is not closed (hence not of the form $d\phi_{\pm}$) unless $F^{(\pm)}=0$. Since the Berry curvatures $F^{(\pm)}$ are generically non-zero, we indeed have $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})=\{0\}$ in the physical geometry.
By Lemma 4.5 (with $A^{(\pm)}$ non‑flat), the space of parallel 1‑forms on the phase circles vanishes: $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{M})=\{0\}$. Consequently the tensor products $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})$ and $\mathcal{P}_{1}(S^{1}_{\phi_{+}})\otimes\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})$ (the latter appears in $\ker\widetilde{\Psi}_{p}$) are zero. Thus
$\mathcal{K}=0,\qquad\text{and therefore}\qquad\dim\mathcal{K}=0.$ (3.2)
3.4 Summary of the topological data
Based on (3.2) we obtain the reduced rank
$r^{\sharp}\;:=\;r-\dim\mathcal{K}\;=\;1.$
All ingredients required for the application of the general geometric theorem are now in place: the extended parameter space $(M,g_{M})$ carries a metric connection $\nabla^{C}$ with totally skew torsion whose harmonic part represents a mixed cohomology class $\omega]$ of mixed tensor rank $r=1$, and the obstruction kernel $\mathcal{K}$ vanishes identically. In the next section, we apply the theorem to obtain a concrete lower bound on the off-diagonal holonomy of $\nabla^{C}$.
4 Geometric foundations for the PT theorem in Kaluza–Klein geometry
4.1 Deformation to a product metric and stability
The physical Kaluza–Klein metric (2.1) is a warped structure that does not admit a global Riemannian product decomposition, whereas Theorem 5.2 in [12] is formally stated for genuine product metrics. To bridge this gap, we introduce a smooth deformation to a product metric and prove the stability of the holonomy bound.
Definition 4.1 (Deformation family).
Let $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ be the smooth one-parameter family of metrics on $M$ defined by:
$g_{\varepsilon}=\pi^{*}g_{\mathrm{BZ}}+(d\phi_{+}+\varepsilon\pi^{*}A^{(+)})^{2}+(d\phi_{-}+\varepsilon\pi^{*}A^{(-)})^{2}.$
For $\varepsilon=0$ we obtain the genuine Riemannian product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$, while for $\varepsilon=1$ we recover the physical Kaluza–Klein metric $g_{M}=g_{1}$.
Lemma 4.2 (Stability of the obstruction data).
For the family $g_{\varepsilon}$ in Definition 4.1, the following hold:
  1. 1.
    The spaces of parallel forms satisfy:
    $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\varepsilon})=\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{0})=\mathbb{R}\operatorname{vol}_{\mathrm{BZ}}\quad\text{for all }\varepsilon\in[0,1].$
  2. 2.
    For $\varepsilon>0$, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{\varepsilon})=\{0\}$. For $\varepsilon=0$, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{0})=\mathbb{R}d\phi_{\pm}$.
  3. 3.
    The mixed tensor rank $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ is constant for all $\varepsilon$.
  4. 4.
    The obstruction kernel $\mathcal{K}_{\varepsilon}$ satisfies $\dim\mathcal{K}_{\varepsilon}=0$ for all $\varepsilon\in[0,1]$.
Proof.
(1) follows because the restriction of $g_{\varepsilon}$ to $T^{2}_{\mathrm{BZ}}$ is $g_{\mathrm{BZ}}$, independent of $\varepsilon$. (2) For $\varepsilon>0$, Lemma 2.4 shows $d\phi_{\pm}$ are not parallel; the proof of Lemma 4.5 will show no other parallel 1-forms exist. For $\varepsilon=0$, $d\phi_{\pm}$ are clearly parallel. (3) is Lemma 4.6 (metric invariance).
(4) For $\varepsilon=0$, the harmonic representative of $\omega]$ is
$\omega_{h}=F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}.$
The subspace $\mathcal{P}_{2}(T^{2})\otimes\mathcal{P}_{1}(S^{1}_{+}\times S^{1}_{-})$ consists of those harmonic 3‑forms that can be written as
$\alpha\wedge\beta,$
where $\alpha$ is a parallel 2‑form on $T^{2}$ (hence a constant multiple of $\text{vol}_{\text{BZ}}$) and $\beta$ is a parallel 1‑form on the fibres (hence a constant linear combination of $d\phi_{+}$ and $d\phi_{-}$).
If $\omega_{h}$ belonged to this subspace, we would have
$F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}=\alpha\wedge\beta=(\lambda\text{vol}_{\text{BZ}})\wedge(c_{+}d\phi_{+}+c_{-}d\phi_{-}),$
which would imply $F^{(+)}=\lambda c_{+}\text{vol}_{\text{BZ}}$ and $F^{(-)}=\lambda c_{-}\text{vol}_{\text{BZ}}$ as differential forms, not just in cohomology. In a generic SOC texture, $F^{(+)}$ and $F^{(-)}$ are not parallel (they are not constant multiples of $\text{vol}_{\text{BZ}}$ pointwise), hence $\omega_{h}\notin\mathcal{P}_{2}\otimes\mathcal{P}_{1}$. Therefore $\mathcal{K}_{0}=0$. ∎
Lemma 4.3 (Lower semi-continuity of off-diagonal holonomy dimension).
The function $F(\varepsilon):=\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ is lower semi-continuous on $[0,1]$, where $\nabla^{C_{\varepsilon}}$ is the metric connection with fixed torsion $T$ constructed from $g_{\varepsilon}$.
Proof.
The curvature tensor $R^{C_{\varepsilon}}$ depends smoothly on $\varepsilon$. At any point $p\in M$, consider the evaluation map
$\Phi_{\varepsilon}:\bigwedge^{2}T_{p}M\times T_{p}M\to\mathfrak{so}(T_{p}M),\quad\Phi_{\varepsilon}(X,Y,Z)=R^{C_{\varepsilon}}(X,Y)Z.$
The off-diagonal holonomy algebra at $p$ is generated by the image of $\Phi_{\varepsilon}$ restricted to arguments that yield off-diagonal operators. The dimension of a finitely generated subspace is a lower semi-continuous function of the generating set (it can only jump upward at discontinuity points). Since $\Phi_{\varepsilon}$ varies continuously in $\varepsilon$, $F(\varepsilon)$ is lower semi-continuous. ∎
Proposition 4.4 (Persistence of the PT bound under deformation).
If the Pigazzini–Toda theorem gives $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\geq r^{\sharp}$ for the product metric $g_{0}$, then the same inequality holds for all $\varepsilon\in[0,1]$, and in particular for the physical metric $g_{M}=g_{1}$.
Proof.
By Lemma 4.2, the topological data $r^{\sharp}=r-\dim\mathcal{K}_{\varepsilon}$ is constant (=r=r) for all $\varepsilon$. The inequality holds at $\varepsilon=0$ by assumption. By Lemma 4.3, the set $\{\varepsilon:\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq r^{\sharp}\}$ is closed (by lower semi-continuity). Since it contains $\varepsilon=0$ and $[0,1]$ is connected, it must be all of $[0,1]$. ∎
4.2 Parallel forms in Kaluza–Klein metrics
Lemma 4.5 (Parallel forms in Kaluza–Klein metrics).
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a Kaluza–Klein metric
$g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2},$
where $g_{1}$ is a Riemannian metric on $M_{1}$, $\{\theta^{a}\}$ defines a flat left-invariant metric $h=\sum_{a}(\theta^{a})^{2}$ on $M_{2}$, and $A^{a}\in\Omega^{1}(M_{1})$.
  1. 1.
    If all connection forms $A^{a}$ are closed (i.e., $dA^{a}=0$), then every $h$-parallel form on the fibre is also $g$-parallel.
  2. 2.
    If at least one curvature $F^{a}:=dA^{a}$ is not identically zero, then the space of $g$-parallel 1-forms tangent to the fibres is trivial:
    $\mathcal{P}_{1}(M_{2},g)=\{0\}.$
Proof.
Write the metric locally as $g=g_{ij}(x)dx^{i}dx^{j}+\sum_{a=1}^{m}(\Theta^{a})^{2}$, where $\Theta^{a}=\theta^{a}+A^{a}_{i}(x)dx^{i}$ and $\{\theta^{a}\}$ is the orthonormal coframe on $M_{2}$ induced by the flat metric $h$. In our SOC BEC case, $m=2$, $\theta^{1}=d\phi_{+}$, $\theta^{2}=d\phi_{-}$.
Let $\xi=\sum_{a=1}^{m}f_{a}(x,y)\Theta^{a}$ be a 1‑form tangent to the fibres. The Levi‑Civita connection of $g$ satisfies the following relation for a horizontal vector field $X=X^{i}\partial_{i}$ and the vertical coordinate field $V_{a}$ dual to $\theta^{a}$:
$\nabla^{LC}_{X}V_{a}=\tfrac{1}{2}F^{a}(X)^{\#},$
where $F^{a}(X)^{\#}$ is the vertical vector field obtained by raising the index of the 1‑form $F^{a}(X,\cdot)$ with the fibre metric $h$. Formula (i) follows from the Koszul formula and uses only that the fibres are flat.
Now suppose $\xi$ is parallel, $\nabla^{LC}\xi=0$. Evaluating on the pair $(X,V_{a})$ gives
$0=(\nabla^{LC}_{X}\xi)(V_{a})=X\bigl(\xi(V_{a})\bigr)-\xi(\nabla^{LC}_{X}V_{a})=X(f_{a})-\tfrac{1}{2}\sum_{b=1}^{m}f_{b}\,h\!\bigl(F^{a}(X)^{\#},V_{b}\bigr).$
Because the functions $f_{a}$ are constant along the fibres (they depend only on the base coordinates $x$ due to the left‑invariance of the metric $h$), equation (ii) reduces to
$\tfrac{1}{2}\sum_{b=1}^{m}f_{b}(x)\,F^{b}(X,V_{a})=0\qquad\text{for every horizontal }X\text{ and every }a.$
If all $F^{b}=0$, condition (iii) is vacuous and any choice of constant $f_{b}$ yields a parallel form, proving statement (i). If, however, some $F^{b_{0}}$ is not identically zero, we can choose a horizontal vector field $X$ and an index $a$ such that $F^{b_{0}}(X,V_{a})\neq 0$ at some point; then (iii) forces $f_{b_{0}}=0$. Repeating the argument for every independent curvature forces all coefficients $f_{b}$ to vanish, hence $\xi=0$. This proves statement (ii) for 1‑forms.
The argument for higher‑degree forms is analogous, using that a parallel $k$-form on the fibre must have constant coefficients with respect to the left‑invariant coframe, and the same curvature coupling appears in the covariant derivative. ∎
Lemma 4.6 (Harmonic representative and Künneth component).
Let $g$ be a Kaluza–Klein metric as in Definition 4.1 on $M=M_{1}\times M_{2}$. Let $\omega]\in H^{k}(M;\mathbb{R})$ be a cohomology class and $\omega_{h}$ its unique $g$-harmonic representative. Then $\omega_{h}$ respects the Künneth decomposition in the sense that its projection onto the $(p,q)$-bidegree component corresponds to the harmonic representative of the mixed class $\omega]_{p,q}\in H^{p}(M_{1})\otimes H^{q}(M_{2})$. In particular, $\omega_{h}$ admits a unique decomposition
$\omega_{h}=\sum_{i=1}^{r}\alpha_{i}\wedge\beta_{i}+\gamma,$
where $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$, the forms $\alpha_{i},\beta_{i}$ are the lifted harmonic representatives of the factor classes, and $\gamma$ contains terms of different bidegrees.
Proof.
By Hodge theory on compact Riemannian manifolds, every de Rham cohomology class $\omega]$ possesses a unique harmonic representative $\omega_{h}$ with respect to the metric $g$. The Künneth theorem provides a canonical isomorphism $H^{k}(M)\cong\bigoplus_{p+q=k}H^{p}(M_{1})\otimes H^{q}(M_{2})$.
While the Hodge Laplacian $\Delta_{g}$ of a Kaluza-Klein metric does not generally commute with the factor-wise Laplacians $\Delta_{g_{1}}$ and $\Delta_{h}$ due to the presence of the connection forms $A^{a}$, the existence of the harmonic representative for each mixed component is guaranteed by the ellipticity of $\Delta_{g}$. The tensor rank $r$ is a purely algebraic property of the class $\omega]$ in the tensor product of the cohomology groups $H^{*}(M_{i})$ and is therefore independent of the specific metric $g$.
The decomposition follows by projecting the global harmonic form $\omega_{h}$ onto the subspace of forms of pure bidegree, which remains well-defined under the orthogonal splitting $TM=D_{1}\oplus D_{2}$ induced by the Kaluza-Klein structure. ∎
4.3 Persistence of non‑parallelism under metric change
To ensure that the lower bound on the off-diagonal holonomy is a robust feature of the model rather than a consequence of a specific metric choice, we must examine the stability of the non-parallelism condition $\nabla^{LC}T\neq 0$. The following result establishes that this property is open in the space of Riemannian metrics, guaranteeing that the topological obstruction remains effective under generic deformations of the Kaluza–Klein geometry.
Lemma 4.7 (Openness of the non‑parallelism condition).
Let $M$ be a smooth manifold and $T\in\Omega^{3}(M)$ a fixed 3‑form. Denote by $\mathcal{M}$ the space of Riemannian metrics on $M$ endowed with the $C^{1}$-topology. The set
$\mathcal{M}_{\neq 0}(T):=\bigl\{g\in\mathcal{M}\mid\nabla^{LC}(g)T\neq 0\bigr\}$
is open in $\mathcal{M}$. In particular, if $\nabla^{LC}(g)T\neq 0$ for some metric $g$, then there exists a $C^{1}$-neighbourhood of $g$ consisting of metrics for which the covariant derivative of $T$ is also non‑zero.
Proof.
The Levi‑Civita connection $\nabla^{LC}(g)$ depends smoothly on the components of the metric $g$ and its first derivatives through the Christoffel symbols. Consequently, the map $g\mapsto\nabla^{LC}(g)T$ is continuous with respect to the $C^{1}$-topology on $\mathcal{M}$. Since the condition $\nabla^{LC}(g)T\neq 0$ defines the complement of the zero set of a continuous map, $\mathcal{M}_{\neq 0}(T)$ is an open subset of $\mathcal{M}$. ∎
Remark 4.8.
In the context of spin-orbit coupled BECs, Lemma 4.7 ensures the structural stability of our topological bound. Since the Kaluza–Klein metric $g$ satisfies $\nabla^{LC}(g)T\neq 0$ due to the non-triviality of the synthetic gauge curvature, any sufficiently small perturbation of the metric—arising, for instance, from experimental fluctuations in the trapping potentials or the SOC strength—will preserve the non-parallelism of the torsion. This confirms that the lower bound on the off-diagonal holonomy is not a fine-tuned property but a robust feature of the SOC BEC geometry.
4.4 Harmonic representatives in Kaluza–Klein geometry
Lemma 4.9 (Metric invariance of the mixed tensor rank).
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold. For any cohomology class $\omega]\in H^{k}(M;\mathbb{R})$, let $\omega]_{\mathrm{mixed}}$ be its projection onto the mixed components of the Künneth decomposition
$\bigoplus_{p+q=k,p,q\geq 1}H^{p}(M_{1};\mathbb{R})\otimes H^{q}(M_{2};\mathbb{R}).$
The mixed tensor rank $r:=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ is a topological invariant of the class $\omega]$. In particular, $r$ is independent of the choice of the Riemannian metric $g$ on $M$.
Proof.
The mixed tensor rank is defined algebraically as the minimal number of simple tensors (elementary tensors of the form $\alpha\otimes\beta$) required to represent $\omega]_{\mathrm{mixed}}$ within the finite-dimensional vector space $\bigoplus H^{p}(M_{1})\otimes H^{q}(M_{2})$. Since the de Rham cohomology groups $H^{p}(M_{i};\mathbb{R})$ are topological invariants associated with the factor manifolds $M_{i}$, the algebraic structure of their tensor product is fixed once the manifold topology is given.
Because the projection of a cohomology class onto its Künneth components is canonical and independent of any Riemannian structure, the rank $r$ remains invariant under any smooth deformation of the metric $g$ on $M$. This ensures that the value of $r$ computed for a Kaluza–Klein metric coincides with the value assigned to the underlying topological class. ∎
4.5 The PT bound for Kaluza–Klein metrics with flat fibres
With the previous preparatory lemmas, we can now state and prove the extension of the Pigazzini-Toda theorem to Kaluza-Klein geometries.
Theorem 4.10 (PT bound for Kaluza–Klein metrics with flat fibres).
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold equipped with a Kaluza–Klein metric
$g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2},$
where $\{\theta^{a}\}$ defines a flat left‑invariant metric $h=\sum_{a}(\theta^{a})^{2}$ on $M_{2}$, and $A^{a}\in\Omega^{1}(M_{1})$. Let $\nabla^{C}=\nabla^{LC}(g)+\frac{1}{2}T$ be a metric connection on $(M,g)$ with totally skew‑symmetric torsion $T\in\Omega^{3}(M)$. Denote by $\omega_{h}$ the unique $g$-harmonic representative of $T]\in H^{3}(M;\mathbb{R})$ and assume
  1. 1.
    $\nabla^{LC}(g)T\neq 0$;
  2. 2.
    $\omega_{h}]$ has a non‑trivial mixed component in the Künneth decomposition;
  3. 3.
    the mixed tensor rank $r:=rank_{\mathbb{R}}\bigl([\omega_{h}]_{\mathrm{mixed}}\bigr)$ is finite.
Define the obstruction space
$\mathcal{K}\;:=\;\bigl(\pi_{2,1}([\omega_{h}])\cap(\mathcal{P}_{2}(M_{1},g)\otimes\mathcal{P}_{1}(M_{2},g))\bigr)\oplus\bigl(\pi_{1,2}([\omega_{h}])\cap(\mathcal{P}_{1}(M_{1},g)\otimes\mathcal{P}_{2}(M_{2},g))\bigr),$
and set $r^{\sharp}:=r-\dim\mathcal{K}$.
Then for every point $p\in M$,
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;r^{\sharp}.$ (4.0)
Moreover, $r^{\sharp}$ depends only on the cohomology class $\omega_{h}]_{\mathrm{mixed}}$ and on the spaces $\mathcal{P}_{k}(M_{i},g)$; consequently, it is invariant under smooth deformations of $g$ that preserve those spaces.
Proof.
The result follows from a direct application of the topological bound established in [12] to the Kaluza–Klein geometry $(M,g,\nabla^{C})$. We verify that the structural requirements are met and compute the resulting invariant.
First, we observe that the metric connection $\nabla^{C}$ satisfies the requisite non-degeneracy conditions. The torsion form $T$, constructed from the synthetic gauge fields, is totally skew-symmetric. Moreover, the condition $\nabla^{LC}(g)T\neq 0$ is satisfied whenever the Berry curvature $F=dA\neq 0$, which holds by assumption in the generic SOC regime.
The mixed tensor rank $r=\text{rank}_{\mathbb{R}}([\omega_{h}]_{\mathrm{mixed}})$ is a cohomological invariant of the class $[T]$; it is independent of the Riemannian metric used to define the harmonic representative, ensuring that $r$ is well-defined and finite.
Crucially, the dimension of the obstruction kernel $\mathcal{K}$ is determined by the intersection of the mixed cohomology components with the spaces of parallel forms. By Lemma 4.5, the non-vanishing Berry curvature $F=dA\neq 0$ implies that the space of fibre-parallel 1-forms vanishes:
$\mathcal{P}_{1}(M_{2},g)=\{0\}.$
Consequently, any tensor product involving $\mathcal{P}_{1}(M_{2},g)$ is zero, and the obstruction space $\mathcal{K}$ is identically zero in the generic SOC BEC regime where synthetic gauge fields are present.
Thus, we have $\dim\mathcal{K}=0$, and the general theorem of [12] then states that the dimension of the off-diagonal holonomy is bounded by $r-\dim\mathcal{K}$, yielding
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;r.$
The invariance of this bound follows from the stability of the topological rank $r$ and the robustness of the vanishing condition $\mathcal{P}_{1}(M_{2},g)=\{0\}$ under metric deformations that preserve the non-integrability of the horizontal distribution. ∎
4.6 Application to the SOC BEC model
Proposition 4.11 (Application to the SOC BEC model).
In the spin‑orbit‑coupled BEC model described in Section 2, the extended parameter space is given by the product manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$. The metric (2.1) is of the Kaluza–Klein type $g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2}$ with the base $M_{1}=T^{2}_{\mathrm{BZ}}$ and the fibre $M_{2}=S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$. The left‑invariant forms are $\theta^{1}=d\phi_{+}$ and $\theta^{2}=d\phi_{-}$, with corresponding gauge potentials $A^{1}=A^{(+)}$ and $A^{2}=A^{(-)}$. The fibres are flat since $d\theta^{a}=0$.
For this specific geometry, the following properties hold:
  1. 1.
    $\mathcal{P}_{1}(M_{2},g)=\{0\}$, due to the non-triviality of the synthetic gauge potentials $A^{(\pm)}$.
  2. 2.
    $\mathcal{P}_{2}(M_{1},g)\cong\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})\cong\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}}$.
  3. 3.
    The obstruction space vanishes identically: $\mathcal{K}=0$.
Furthermore, if the Berry curvatures satisfy $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}})$, then the mixed tensor rank is $r=1$, and Theorem 4.10 implies:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
Proof.
The identification of the BEC metric with the Kaluza–Klein framework follows directly from the construction in Section 2. The first item is a direct consequence of Lemma 4.5; as established in the physical model, the synthetic gauge potentials $A^{(\pm)}$ possess non-vanishing Berry curvature, which acts as an obstruction to the existence of parallel 1-forms along the fibres.
On the base manifold $M_{1}=T^{2}_{\mathrm{BZ}}$, the restriction map onto $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ is an isomorphism, identifying the space of parallel 2-forms with the span of the Riemannian volume form. Since $\mathcal{P}_{1}(M_{2},g)=\{0\}$, the tensor products in the definition of $\mathcal{K}$ vanish, ensuring that the obstruction space is trivial.
Furthermore, since $H^{2}(T^{2}_{\mathrm{BZ}})$ is 1-dimensional, any two non-zero classes are linearly dependent. Thus, we may write $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda\in\mathbb{R}$, and the mixed component of the torsion class becomes:
$\omega]_{\mathrm{mixed}}=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
which represents a single elementary tensor, implying $r=1$. Applying Theorem 4.10 with $r=1$ and $\dim\mathcal{K}=0$ yields the lower bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$. ∎
5 Topological Lower Bound and Its Physical Consequences
With the geometric construction of Section 2 and the cohomological data of Section 3 at hand, we can now apply the general theorem of [12] to the physical connection $\nabla^{C}$. This section establishes the fundamental lower bound on the holonomy algebra for the SOC BEC model.
5.1 Statement of the bound
The general PT theorem establishes that for a metric connection $\nabla^{C}$ with totally skew-symmetric torsion $T$ on a product Riemannian manifold $M=M_{1}\times M_{2}$, the dimension of the off-diagonal holonomy is constrained by the topology of the torsion class and the geometry of the metric. Specifically, if the condition $\nabla^{LC}T\neq 0$ holds, we have:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r-\dim\mathcal{K},$ (5.1)
where $r$ is the mixed tensor rank of the class $[T]$ and $\mathcal{K}$ is the obstruction space defined as the intersection of the mixed cohomology components with the spaces of parallel forms on the factors.
5.2 Main theorem: topological lower bound for SOC BECs
Theorem 5.1 (Topological lower bound for SOC BECs).
Let $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ be the parameter space equipped with the Kaluza–Klein metric $g_{M}$ and the metric connection $\nabla^{C}$ with torsion $T^{b}=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}$. Assuming the Berry curvatures satisfy $F^{(\pm)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{R})$, the following hold:
  1. 1.
    The mixed tensor rank of the torsion class $T^{b}]$ is $r=1$, which is the maximal value for this configuration.
  2. 2.
    The obstruction kernel $\mathcal{K}$ associated with the Kaluza–Klein splitting vanishes identically: $\mathcal{K}=\{0\}$.
  3. 3.
    The off-diagonal holonomy algebra of $\nabla^{C}$ satisfies the sharp lower bound:
    $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$ (5.2)
Proof.
The proof proceeds by synthesizing the topological and geometric invariants derived in the preceding sections.
1. Mixed tensor rank. As established in Lemma 4.9, the rank $r$ is a purely topological invariant of the class $\omega]=[T^{b}]$. Under the Künneth decomposition for $H^{3}(M;\mathbb{R})$, the torsion class is expressed as $\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$. Since the cohomology group $H^{2}(T^{2}_{\mathrm{BZ}})$ is one-dimensional, any two non-zero classes are linearly dependent. Thus, we can write $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda\in\mathbb{R}$, which allows the class to be factorized as the simple tensor:
$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr).$
This represents an element of rank $r=1$ in the tensor product of the cohomology groups.
2. Vanishing of the obstruction kernel. The space $\mathcal{K}$ is defined by the intersection of the mixed components of $\omega]$ with the subspaces of forms $\mathcal{P}_{k}(M_{i},g_{M})$ that are parallel with respect to the Levi-Civita connection of $g_{M}$. Crucially, Lemma 4.5 proves that for a Kaluza-Klein metric, the existence of non-vanishing Berry curvature $F\neq 0$ acts as a differential-geometric obstruction to the existence of parallel 1‑forms along the fibres. Consequently, $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}},g_{M})=\{0\}$. Since the mixed components of $\omega]$ involve 1-forms on the fibre factors, their intersection with the parallel-form strata is trivial, leading to $\dim\mathcal{K}=0$.
3. Application of the PT bound. The Kaluza–Klein metric $g_{M}$ satisfies the conditions for the extended PT theorem as stated in Theorem 4.10. Since the non-vanishing Berry curvature acts as a topological obstruction that prevents the existence of global parallel sections, the dimension of the obstruction space $\mathcal{K}$ is zero; consequently, the bound holds directly for the physical Kaluza–Klein metric $g_{M}$ without recourse to auxiliary product metrics. Substituting the invariants $r=1$ and $\dim\mathcal{K}=0$ into the inequality (5.1), we obtain $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$.
The geometric implication is that there exists at least one independent curvature operator $R^{C}(X,Y)$ that interchanges the tangent spaces of the Brillouin zone and the phase fibres. This lower bound ensures that the synthetic gauge field curvature ”locks” the different sectors of the parameter space, preventing a trivial decoupling of the internal and external degrees of freedom in the SOC BEC. ∎
5.3 Physical interpretation: non-removable off-diagonal curvature
Inequality (5.1) has a direct geometric meaning. The holonomy algebra $\mathfrak{hol}(\nabla^{C})$ is generated by the curvature operators $R^{C}(X,Y)$. The “off-diagonal” part $\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})$ consists of those generators that map the tangent space of the Brillouin zone $T_{k}T^{2}_{\mathrm{BZ}}$ into the phase directions $\mathbb{R}\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$ and vice-versa. The bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ therefore guarantees that there exists at least one linearly independent curvature operator that mixes momentum with the phase directions.
In physical terms, this means that the synthetic gauge curvature of the SOC BEC cannot be made completely block-diagonal by any smooth gauge transformation or deformation of the condensate wavefunction. One cannot “flatten” the Berry curvature along both phase directions simultaneously for all points in the Brillouin zone. The topology of the mixed class $\omega]$ forces a minimal geometric entanglement between the momentum and phase degrees of freedom.
5.4 Persistence of the bound when the total Chern number vanishes
A crucial feature of the bound is that it is insensitive to the value of the total first Chern number
$c_{1}^{\,\mathrm{tot}}=\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}\bigl(F^{(+)}+F^{(-)}\bigr).$
Corollary 5.2 (Bound independent of the total Chern number).
Let $c_{1}^{(+)}=\frac{1}{2\pi}\int_{T^{2}}F^{(+)}$ and $c_{1}^{(-)}=\frac{1}{2\pi}\int_{T^{2}}F^{(-)}$. If $c_{1}^{(+)}\neq 0$ and $c_{1}^{(-)}=-c_{1}^{(+)}$, then $c_{1}^{\,\mathrm{tot}}=0$. Nevertheless, provided $F^{(+)}]\neq 0$ (hence also $F^{(-)}]\neq 0$), the mixed tensor rank remains $r=1$ and Theorem 5.1 still gives
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
Thus the topological obstruction persists even when the net Berry flux vanishes.
Proof.
If $c_{1}^{(+)}\neq 0$ then $F^{(+)}]\neq 0$ in $H^{2}(T^{2})$. The condition $c_{1}^{(-)}=-c_{1}^{(+)}$ implies $F^{(-)}]=-[F^{(+)}]$ in cohomology (because integration gives an isomorphism $H^{2}(T^{2})\cong\mathbb{R}$). Hence both classes are non-zero and the mixed tensor rank attains its maximum $r=1$. The vanishing of $\mathcal{K}$ does not depend on the values of the Chern numbers but only on the non-existence of parallel one-forms on the fibres, which holds as long as $F^{(\pm)}\neq 0$. ∎
This demonstrates that $r^{\sharp}$ detects a local topological constraint that survives when the familiar global invariant vanishes. While the Chern number quantizes the net flux, it can be zero due to cancellation of local contributions; the mixed-rank bound, in contrast, certifies that the curvature cannot be simultaneously flattened along the two independent phase directions.
5.5 Invariance under metric deformations
The lower bound (5.1) depends only on the topological data $r$ and on the spaces of parallel forms $\mathcal{P}_{k}(M_{i})$. In our construction, these spaces are determined by the Kaluza–Klein metric $g_{M}$, which in turn is built from the physical Berry connections $A^{(\pm)}$. Any smooth deformation of the system that does not change the cohomology classes of $A^{(\pm)}$ and that preserves the vanishing of $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})$ (i.e., that does not make the Berry curvatures exactly flat) leaves the bound unchanged. This stability makes the result robust against typical experimental perturbations.
The inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ is the main result of this paper. It provides a cohomologically defined, locally non-trivial obstruction to the complete removal of Berry curvature in a two-component SOC BEC, valid even when the total Chern number is zero. In the following section, we illustrate how the same cohomological framework reproduces, in a particular Rashba–Dresselhaus texture, the familiar Chern-number bound on vortex count, while emphasizing that the holonomy bound operates beyond the Chern-number paradigm.
6 Physical Consequences and Experimental Signatures
The lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ imposes fundamental structural constraints on the synthetic gauge architecture of SOC BECs. These manifest as irreducible couplings between the external (momentum) and internal (phase) degrees of freedom.
6.1 Geometric obstruction to gauge flattening
The bound ensures the existence of at least one independent off-diagonal curvature operator that non-trivially mixes the momentum tangent space $T_{k}T^{2}_{\mathrm{BZ}}$ with the phase fibre directions $\mathbb{R}\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$. In terms of the synthetic gauge field components, this implies that no smooth gauge transformation or continuous deformation of the condensate order parameter can simultaneously “flatten” the curvature components:
$\Omega_{j,\pm}:=\frac{\partial A_{\phi_{\pm}}}{\partial k_{j}}-\frac{\partial A_{k_{j}}}{\partial\phi_{\pm}}.$
The topology of the mixed cohomology class $\omega]$ enforces a minimal geometric coupling between the Brillouin zone dynamics and the dual-phase evolution. Unlike trivial gauge fields, where the connection can be locally reduced to a longitudinal form, the non-vanishing rank $r^{\sharp}$ certifies that at least one independent transverse coupling is irreducible. This “topological locking” prevents the independent manipulation of the momentum and the relative phases $\phi_{\pm}$ in any experimental protocol that preserves the SOC texture.
6.2 Locality and robustness beyond the Chern-number paradigm
The primary physical advantage of the holonomy bound lies in its sensitivity to the local structure of the Berry curvature, whereas the total Chern number $c_{1}^{\mathrm{tot}}$ is a global aggregate that may vanish due to cancellation.
In configurations where the individual fluxes satisfy $F^{(+)}]\approx-[F^{(-)}]$, the global invariant $c_{1}^{\mathrm{tot}}$ vanishes, potentially suggesting a topologically trivial regime. However, the inequality $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ persists as long as the local field strength remains non-zero. This identifies $r^{\sharp}$ as a robust local invariant: it detects the persistence of topological effects, such as the non-integrability of the Berry connection, even in compensated systems or “synthetic vacuum” configurations where net topological charges are absent.
From an experimental perspective, this implies that effects typically associated with non-trivial topology—such as non-Abelian transport characteristics or anomalous velocity contributions—can be observed in SOC BECs even when the total Chern number is zero, provided the mixed tensor rank $r$ is preserved.
7 Illustrative Examples: From Chern Numbers to the Mixed Rank
The general lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ derived in Section 5 establishes a universal constraint for generic two-component SOC BECs. In this section, we illustrate the operational significance of this bound in two paradigmatic scenarios: first, the standard Rashba–Dresselhaus texture, where our framework recovers the phenomenology of Chern-number-induced obstructions; second, a configuration with vanishing net Chern flux, where the mixed-rank invariant $r^{\sharp}$ reveals topological features inaccessible to traditional global invariants.
7.1 Example 1: Rashba–Dresselhaus texture and the maximal rank $r=1$
Consider a two-dimensional, two-level Hamiltonian of the Rashba–Dresselhaus (RD) type:
$\hat{H}_{\mathrm{RD}}(k)\;=\;\frac{1}{2m}\bigl(\mathbf{p}-\mathbf{A}_{\mathrm{syn}}(k)\bigr)^{2}\;+\;\mathbf{B}_{\mathrm{syn}}(k)\cdot\boldsymbol{\sigma},\qquad k=(k_{x},k_{y})\in T^{2}_{\mathrm{BZ}},$
where $\mathbf{A}_{\mathrm{syn}}(k)$ and $\mathbf{B}_{\mathrm{syn}}(k)$ represent the synthetic vector potential and Zeeman field, respectively. For generic coupling parameters $\kappa$ and $\Omega$, the lower dressed band induces a spin texture map $\mathbf{n}:T^{2}_{\mathrm{BZ}}\to S^{2}$.
The associated eigenline bundle $\mathbb{L}\to T^{2}_{\mathrm{BZ}}$ is the pullback of the tautological line bundle over $\mathbb{C}P^{1}\cong S^{2}$ via $\mathbf{n}$. The first Chern number $c_{1}(\mathbb{L})$ corresponds to the degree of the map $\mathbf{n}$. For a minimal RD texture with $d=1$, the Berry curvature $F$ represents a non-trivial cohomology class $F]\in H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{Z})$, such that $\frac{1}{2\pi}\int F=\pm 1$.
In the extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the synthetic torsion class $\omega]$ is given by:
$\omega]=[F]\otimes[d\phi_{+}]\;+\;[F]\otimes[d\phi_{-}]\in H^{3}(M;\mathbb{R}).$
Since $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the mixed components are linearly dependent over the base, allowing the class to be expressed as the simple tensor:
$\omega]=[F]\otimes\bigl([d\phi_{+}]+[d\phi_{-}]\bigr).$
Consequently, the mixed tensor rank is $r=1$. Given that the obstruction kernel $\mathcal{K}$ vanishes in this geometry, we obtain $r^{\sharp}=1$. Theorem 5.1 then yields:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
We note that for a global Rashba texture with non-zero Chern number, the bundle is topologically non-trivial, preventing a global product decomposition. However, applying the theorem within a local trivialization demonstrates that the mixed rank $r^{\sharp}$ correctly identifies the ’locking’ of momentum and phase degrees of freedom. In this context, while the non-vanishing Chern number already implies the existence of vortices on the Brillouin zone, the invariant $r^{\sharp}$ provides an algebraically refined constraint, certifying that the momentum and phase degrees of freedom remain fundamentally coupled through the off-diagonal curvature.
7.2 Example 2: Topological persistence under vanishing net Chern flux
The efficacy of the mixed-rank invariant is most evident when global topological charges vanish. Consider a configuration where the two $U(1)$ sectors carry Berry curvatures $F^{(+)}$ and $F^{(-)}$ with equal and opposite integrated fluxes:
$\frac{1}{2\pi}\int_{T_{\mathrm{BZ}}^{2}}F^{(+)}=+1,\qquad\frac{1}{2\pi}\int_{T_{\mathrm{BZ}}^{2}}F^{(-)}=-1.$
In this case, the total Chern number $c_{1}^{\mathrm{tot}}=c_{1}^{(+)}+c_{1}^{(-)}=0$. Traditional Chern-number diagnostics would categorize this regime as topologically trivial. Crucially, the vanishing of the total Chern flux ensures that the synthetic connection $A$ can be defined as a global 1-form, making the product structure $M=T_{\mathrm{BZ}}^{2}\times S^{1}\times S^{1}$ globally well-defined and consistent with our geometric framework. In this globally trivialized setting, the torsion class remains:
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}].$
Since $F^{(-)}]=-[F^{(+)}]$ in $H^{2}(T_{\mathrm{BZ}}^{2})$, the class factorizes into a single simple tensor:
$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]-[d\phi_{-}]\bigr).$
Thus, $r=1$ and $r^{\sharp}=1$, forcing the holonomy bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$. This proves that even in the absence of a net global flux, the Berry curvature cannot be flattened simultaneously along both phase directions. Such a configuration is protected by a local topological obstruction that is invisible to integrated invariants but perfectly captured by the mixed cohomology rank on the product manifold.
To illustrate how the bound materializes geometrically, consider the concrete choice
$A^{(+)}=\frac{1}{4\pi}(k_{x}dk_{y}-k_{y}dk_{x}),\qquad A^{(-)}=-A^{(+)},$
which yields constant Berry curvatures $F^{(\pm)}=\pm\frac{1}{2\pi}dk_{x}\wedge dk_{y}$ and satisfies ($\ast$) on a torus of period $2\pi$ in each direction. The corresponding Kaluza–Klein metric is
$g_{M}=dk_{x}^{2}+dk_{y}^{2}+\Bigl(d\phi_{+}+\tfrac{1}{4\pi}(k_{x}dk_{y}-k_{y}dk_{x})\Bigr)^{2}+\Bigl(d\phi_{-}-\tfrac{1}{4\pi}(k_{x}dk_{y}-k_{y}dk_{x})\Bigr)^{2}.$
A straightforward computation of the Levi-Civita connection gives the non-zero mixed Christoffel symbols
$\Gamma^{k_{x}k_{y}}_{\phi_{+}}=-\Gamma^{k_{y}k_{x}}_{\phi_{+}}=\frac{1}{8\pi},\qquad\Gamma^{k_{x}k_{y}}_{\phi_{-}}=-\Gamma^{k_{y}k_{x}}_{\phi_{-}}=-\frac{1}{8\pi},$
all other mixed components being zero.
The torsion 3-form is
$T=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}=\frac{1}{8\pi^{2}}(dk_{x}\wedge dk_{y})\wedge(k_{x}dk_{y}-k_{y}dk_{x}).$
For the connection $\nabla^{C}=\nabla^{LC}+\tfrac{1}{2}T$, the off-diagonal curvature components evaluate to
$R^{C}(\partial_{k_{x}},\partial_{k_{y}})\,\partial_{\phi_{+}}=\frac{1}{4\pi}\,\partial_{\phi_{-}}\neq 0,\qquad R^{C}(\partial_{k_{x}},\partial_{k_{y}})\,\partial_{\phi_{-}}=-\frac{1}{4\pi}\,\partial_{\phi_{+}}\neq 0,$
where the contributions from the Levi-Civita curvature and the torsion combine additively.
These two operators are linearly independent, confirming that
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 2>1=r^{\sharp}.$
Remark 7.1.
The explicit computation confirms that:
  1. 1.
    The off-diagonal curvature is non‑vanishing even though $c_{1}^{\mathrm{tot}}=0$.
  2. 2.
    The topological bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ is not merely existential; it is realized by concrete curvature operators that can be written down in closed form.
  3. 3.
    The geometric obstruction manifests as a direct coupling between momentum derivatives $(\partial_{k_{x}},\partial_{k_{y}})$ and phase derivatives $(\partial_{\phi_{+}},\partial_{\phi_{-}})$, preventing any gauge transformation from simultaneously diagonalizing the connection with respect to the product splitting $T_{\mathrm{BZ}}^{2}\times S^{1}\times S^{1}$.
7.3 Comparison: Global Invariants vs. Mixed Rank
The distinction between $c_{1}^{\mathrm{tot}}$ and $r^{\sharp}$ reflects the difference between integrated and structural topology. The Chern number $c_{1}^{\mathrm{tot}}$ is a global aggregate that can vanish via local cancellations. In contrast, the mixed rank $r$ counts the number of irreducible tensorial couplings between the factor manifolds.
In our geometry, $r=1$ is attained whenever at least one individual Berry curvature class is non-zero. It serves as a measure of the irreducible geometric coupling between the Brillouin zone torus and the phase fibres. The bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq r^{\sharp}$ therefore constrains the algebraic complexity of the curvature, ensuring that the synthetic gauge field remains ”locked” in a non-trivial configuration even when the net topological charge is zero.
7.4 Summary
The Rashba–Dresselhaus case shows that $r^{\sharp}$ is consistent with established Chern-number results. However, the $c_{1}^{\mathrm{tot}}=0$ case demonstrates that the mixed-rank framework provides a finer classification, detecting locally non-removable curvature that conventional global invariants fail to resolve. This makes $r^{\sharp}$ a superior tool for analyzing the stability of synthetic gauge fields in complex SOC BEC architectures.
8 Beyond the Chern Number: Local Rail-Tracks and Prospective Protocols
While the Chern number $c_{1}$ quantifies the net vorticity through a global integral over the Brillouin zone, the lower bound on the off-diagonal holonomy dimension, $r^{\sharp}$, acts as a local structural constraint. It obstructs the simultaneous annihilation of Berry curvature components along independent directions within the product geometry. This phenomenon can be visualized as the Berry curvature being constrained to move along topological “rail-tracks” in the configuration space: even when the total topological charge is tuned to zero, specific directional components remain topologically locked and cannot be erased by any smooth gauge transformation.
In the two-component SOC BEC manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the bound $r^{\sharp}=1$ is non-degenerate since the non-vanishing Berry curvature breaks the integrability of the horizontal distribution, ensuring $\mathcal{K}=0$. Consequently, the bound holds directly for the physical Kaluza–Klein metric $g_{M}$. The curvature of any metric connection calibrated by the mixed class $\omega]$ is thus irreducible to a block-diagonal form; at least one independent off-diagonal operator persists, fundamentally mixing momentum dynamics with the phase evolution of $\phi_{+}$ and $\phi_{-}$.
This geometric prediction suggests concrete interferometric tests. Consider the simultaneous measurement of the Berry phases $\Phi_{B}^{+}(\gamma)$ and $\Phi_{B}^{-}(\gamma)$ accumulated along a closed loop $\gamma\subset T^{2}_{\mathrm{BZ}}$ by separately addressing the global and relative phase sectors. If the gauge field were fully reducible (i.e., if the holonomy were trivial), one could choose a gauge where both phases vanish for every $\gamma$. The bound $r^{\sharp}=1$ forbids such a simultaneous vanishing across all loops. For any pair of independent cycles $\{\gamma_{1},\gamma_{2}\}$ in $T^{2}_{\mathrm{BZ}}$, the two-component Berry-phase vector:
$\mathbf{\Phi}_{B}=\bigl(\Phi_{B}^{+}(\gamma_{1}),\Phi_{B}^{-}(\gamma_{1}),\Phi_{B}^{+}(\gamma_{2}),\Phi_{B}^{-}(\gamma_{2})\bigr)$
cannot be continuously deformed to zero while preserving the mixed cohomology class $\omega]$.
Such measurements are within the reach of current cold-atom experimental capabilities. Interferometric techniques used to map Berry curvature in Hofstadter bands [2] or to probe spin-texture topology in SOC systems [9] can be adapted to track the two phase degrees of freedom $\phi_{\pm}$ independently [5]. The required control can be implemented via optical “painting” [8], allowing for precise spatiotemporal manipulation of the condensate’s phase profile.
Observing that these Berry-phase components cannot be simultaneously flattened would provide the first direct experimental signature of the topological obstruction captured by $r^{\sharp}$. This demonstrates that the mixed-cohomology framework not only yields a rigorous mathematical lower bound but also provides a feasible protocol for detecting locally irremovable Berry curvature, effectively moving the study of quantum gases beyond the global Chern-number paradigm.
9 Conclusions
We have demonstrated that a two-component spin–orbit-coupled (SOC) BEC possesses an intrinsic geometric obstruction that prevents the simultaneous flattening of Berry curvature along the independent phase directions $\phi_{\pm}$. This obstruction persists even when the net global Chern flux vanishes, identifying a topological regime that eludes traditional integrated invariants.
The extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, equipped with the natural Kaluza–Klein metric $g_{M}$ induced by synthetic gauge fields, carries a metric connection $\nabla^{C}$ with totally skew-symmetric torsion. The harmonic part of this torsion represents a mixed cohomology class $\omega]$ with mixed tensor rank $r=1$ for any generic configuration where the individual Berry curvature classes are non-trivial. Due to the one-dimensionality of $H^{2}(T^{2}_{\mathrm{BZ}})$, these classes are necessarily proportional, forcing $\omega]$ to take the form of a single simple tensor. Crucially, we have shown that the non-vanishing Berry curvature acts as a geometric obstruction to the integrability of the horizontal distribution, ensuring the vanishing of the obstruction kernel $\mathcal{K}=0$. Consequently, the bound $r^{\sharp}=1$ holds directly for the physical metric $g_{M}$.
Applying the PT lower bound yields the sharp inequality:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
This algebraic constraint implies that momentum degrees of freedom are fundamentally “locked” to the phase sectors, an irreducible coupling that cannot be removed by smooth gauge transformations. The persistence of $r^{\sharp}$ in regimes with zero total Chern number—as illustrated by our refined Rashba–Dresselhaus analysis—highlights that the mixed-rank bound operates beyond the Chern-number paradigm, detecting structural features of the gauge field that integrated invariants fail to resolve.
Experimentally, this framework suggests interferometric protocols to resolve independent Berry phases, providing a direct signature of locally irremovable curvature. Conceptually, it establishes a new mechanism for topological robustness in synthetic gauge systems, controlled by the algebraic structure of mixed cohomology components. We expect this framework to be readily applicable to multicomponent cold-atom systems and synthetic-dimensional lattices, where multiple internal degrees of freedom couple to momentum-space geometry.
Acknowledgements
The authors gratefully acknowledge stimulating discussions with colleagues in geometry and cold-atom physics that helped shape the formulation of these results.