### v3

# Local Topological Constraints on Berry Curvature in Spin–Orbit Coupled BECs

Abstract
We establish a local topological obstruction to flattening Berry curvature in spin-orbit-coupled Bose-Einstein condensates (SOC BECs), valid even when the global Chern number vanishes. For a generic two-component SOC BEC, the extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ carries a natural metric connection $\nabla^{C}$ whose torsion 3-form encodes the synthetic gauge fields. Under the physically relevant assumption of constant Berry curvatures, the harmonic part of this torsion defines a mixed cohomology class
$\omega]\in\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
whose mixed tensor rank equals one. Using a general geometric bound for metric connections with totally skew torsion on product manifolds, we show that the obstruction kernel $\mathcal{K}$ vanishes, yielding the sharp inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$. This forces at least one off-diagonal curvature operator, preventing complete gauging-away of Berry phases even when the total Chern number is zero. This provides the first cohomological lower bound certifying locally irremovable curvature in SOC BECs beyond the Chern-number paradigm.
Keywords: Berry curvature; spin–orbit-coupled Bose–Einstein condensates; synthetic gauge fields; holonomy with torsion; mixed cohomology; Kaluza–Klein geometry.
MSC (2020): 53C29, 53C80, 81V45, 82B10, 51P05, 55S20, 57R19, 57R56, 58A12.
1 Introduction
The study of spin–orbit-coupled (SOC) Bose–Einstein condensates (BECs) stands at a rich confluence of differential geometry, quantum many-body physics, and synthetic gauge theory. This interdisciplinary effort is part of the broader pursuit of artificial gauge fields in engineered quantum systems, spanning ultracold atoms, photonics, and solid-state platforms [1]. Advances in generating artificial gauge fields with Raman lasers have made these systems a versatile laboratory for exploring Berry curvature, synthetic magnetic fields, and non-Abelian textures with unprecedented control [6, 7, 14]. These studies are often conducted in toroidal traps, which can serve as magnetic storage rings for BECs [3], providing a well-controlled environment for synthetic gauge field experiments. Remarkable achievements include the observation of Dirac monopoles in synthetic magnetic fields [10], demonstrating the high degree of geometric control now attainable. The geometric phase underpinning these phenomena traces back to the foundational work of Simon [13], while the experimental realization of artificial gauge fields has been comprehensively reviewed in [4].
A persistent theoretical challenge is to identify local topological obstructions that constrain the curvature even when global invariants—such as the first Chern number—vanish. While the Chern number quantizes the net flux, its vanishing does not guarantee that the curvature can be locally flattened; geometric obstructions may prevent its simultaneous annihilation along independent directions in the parameter space.
This work addresses that challenge by bringing a recent geometric result into physical focus. In [12], a lower bound was established for the off-diagonal holonomy of metric connections with totally skew-symmetric torsion on product manifolds. The bound is expressed in terms of a mixed de Rham cohomology class $\omega]\in H^{p}(M_{1})\otimes H^{q}(M_{2})$ associated with the torsion; its mixed tensor rank $r$ counts the minimal number of simple tensors needed to represent the class, and after subtracting contributions from parallel forms (which do not generate off-diagonal curvature) one obtains a corrected integer $r^{\sharp}$ that bounds the dimension of the off-diagonal holonomy algebra:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}=r-\dim\mathcal{K}.$
Clarification of methodological approach.
While Theorem 5.2 of [12] is formally stated for genuine Riemannian product metrics $g=g_{1}\oplus g_{2}$, the physical metric $g_{M}$ in our SOC BEC model is a Kaluza–Klein metric that is not globally a Riemannian product when the Berry curvatures are non-zero. We do not apply the theorem as a black box. Instead, we adapt its proof to our geometry by: (i) using the pointwise orthogonal splitting $\mathcal{H}_{p}\oplus\mathcal{V}_{p}$ to define obstruction maps $\Psi_{p},\widetilde{\Psi}_{p}$ analogously to those in [12]; (ii) showing that the harmonic representative of the torsion class decomposes appropriately; and (iii) proving that the kernel computation leading to $\mathcal{K}=0$ remains valid. This adaptation, supplemented by a deformation argument ($g_{\varepsilon}\to g_{M}$) and lower-semicontinuity, rigorously yields the bound for the physical metric $g_{M}$.
Here we show how this abstract geometric theorem materializes in a concrete physical setting: a generic two-component SOC BEC confined in a quasi-two-dimensional toroidal trap. Such a system possesses three universal, experimentally accessible degrees of freedom: the two crystal momenta forming a Brillouin-zone torus $T^{2}_{\mathrm{BZ}}$, a global $U(1)$ phase $\phi_{+}$ associated with particle-number conservation, and a relative $U(1)$ phase $\phi_{-}$ between the two internal spin components [9, 11]. We therefore take as the extended parameter space the manifold
$M=T_{BZ}^{2}\times S_{\phi_{+}}^{1}\times S_{\phi_{-}}^{1}.$
We emphasize that treating $M$ as a global product manifold requires the underlying principal bundle to be trivial, which holds precisely in the physical regime of vanishing total Chern number ($c_{\text{tot}}=0$) that is central to our main result. In configurations with non-zero net topological charge, this product structure should be understood as a local description in a gauge patch, sufficient for deriving local curvature bounds.
The first main task of the paper is to construct, in this space, a natural geometric structure to which the theorem of [12] applies. We introduce a metric $g_{M}$ of Kaluza–Klein type, which couples the base metric on $T^{2}_{\mathrm{BZ}}$ to the $U(1)$-valued connection forms that encode the synthetic gauge fields of the SOC BEC. For simplicity and to make the geometric arguments fully explicit, we will work in the experimentally relevant setting where the Berry curvatures are constant multiples of the Brillouin–zone volume form. This covers a wide range of SOC BEC realizations and allows us to write the harmonic representative of the torsion class in closed form. On $(M,g_{M})$ we define a metric connection $\nabla^{C}$ whose torsion three-form is precisely the physical three-form
$T^{\flat}=\operatorname{Tr}(\mathcal{F}\wedge\mathbf{A}_{\mathrm{syn}}),$
built from the synthetic Berry curvature $\mathcal{F}$ and connection $\mathbf{A}_{\mathrm{syn}}$. The harmonic part of $T^{\flat}$ with respect to $g_{M}$ then defines a mixed cohomology class
$\omega]\in H^{3}(M)\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr).$
The second main task is to compute the topological data that enters the general bound. We prove that, under the generic condition that the two nearly degenerate bands carry non-zero Berry curvature classes $F^{(+)}],[F^{(-)}]\in H^{2}(T^{2}_{\mathrm{BZ}})$, the mixed tensor rank of $\omega]$ attains its maximal possible value within this geometry. Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, any two non-zero classes in this space are cohomologically proportional. Consequently, the mixed class $\omega]$ can be written as a single simple tensor
$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
and its mixed tensor rank is therefore $r=1$. Moreover, owing to the specific form of the metric $g_{M}$, the spaces of parallel one-forms on the phase circles are trivial, which forces the obstruction kernel $\mathcal{K}$ to vanish. Consequently the reduced rank is $r^{\sharp}=1$, and the general theorem yields the sharp inequality
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
This lower bound has immediate physical consequences. It implies that the curvature of the synthetic gauge connection cannot be made block-diagonal with respect to the splitting between momentum and phase directions; at least one independent off-diagonal operator persists, mixing the Brillouin zone with the two phase directions. Crucially, the bound detects the topological coupling of the momentum torus to the two distinct $U(1)$ phases, and therefore remains non-trivial even when the total Chern number of the lowest band is tuned to zero. Thus $r^{\sharp}$ provides a cohomological certificate of locally irremovable Berry curvature, invisible to conventional global invariants. The result suggests interferometric protocols that separately probe the Berry phases associated with $\phi_{+}$ and $\phi_{-}$, offering a way to detect this obstruction even in regimes of vanishing net vorticity.
The paper is organized as follows. Section 2 constructs the extended parameter space $M$, the Kaluza–Klein metric $g_{M}$, and the metric connection $\nabla^{C}$ with torsion $T^{\flat}$. Section 3 analyses the cohomology of $M$, computes the mixed tensor rank $r$, and shows that $\mathcal{K}=0$ in the present geometry. Section 4 extends the Pigazzini–Toda theorem to Kaluza–Klein metrics, ensuring the bound applies to the present geometry. Section 5 applies the general theorem of [12] to obtain the lower bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ and discusses its physical interpretation. Section 6 elaborates the physical implications of the bound, including its persistence when the total Chern number vanishes and its robustness under metric deformations. Section 7 illustrates how the same cohomological data reproduce, in a particular Rashba–Dresselhaus texture, the familiar Chern-number bound on the vortex number, and demonstrates with a tuned example that the holonomy bound remains valid even when the Chern number vanishes. Section 8 outlines prospective interferometric tests for detecting the locally irremovable curvature and indicates directions for future work. Section 9 summarises the results and discusses broader implications for topological quantum matter.
2 Geometric Construction of the Extended Parameter Space
A generic two-component spin–orbit-coupled Bose–Einstein condensate (SOC BEC) confined in a quasi-two-dimensional toroidal trap is characterized by three independent, experimentally accessible degrees of freedom: the crystal momenta $(k_{x},k_{y})$ spanning a Brillouin zone, a global $U(1)$ phase $\phi_{+}$ associated with particle-number conservation, and a relative $U(1)$ phase $\phi_{-}$ between the two internal spin states [9, 11].
We therefore take as the extended parameter space the smooth, compact, oriented product manifold
$M=T_{\mathrm{BZ}}^{2}\times S_{\phi_{+}}^{1}\times S_{\phi_{-}}^{1}.$
Global structure of the extended parameter space.
The extended parameter space is the total space $M$ of a principal $U(1)_{+}\times U(1)_{-}$-bundle over the Brillouin torus, equipped with the synthetic connections $A^{(+)},A^{(-)}$. As a smooth manifold, $M$ is diffeomorphic to the four-torus $T^{4}$. This diffeomorphism allows us to identify $M$ with a product $T^{2}_{\rm BZ}\times T^{2}_{\rm fibre}$ for the purpose of applying topological results, where $T^{2}_{\rm fibre}\cong S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$.
The bundle is trivial if and only if each individual Chern class vanishes: $c^{(+)}_{1}=0$ and $c^{(-)}_{1}=0$. The physically interesting regime of compensated topology, where $c^{(+)}_{1}=-c^{(-)}_{1}\neq 0$, corresponds to a non-trivial bundle. However, the central results of this paper—constructed from the globally defined metric $g_{M}$ and connection forms $\Theta^{(\pm)}$—do not require bundle triviality, only the global geometric structures on $M\cong T^{4}$. The condition $c^{\rm tot}_{1}=0$ ensures the existence of a global section for the associated $S^{1}$-bundle related to the diagonal subgroup, which is sufficient for the physical interpretation of separable phase variables in a global gauge.
This is precisely the regime where conventional global invariants (like the total Chern number) fail to detect local geometric obstructions, and where the cohomological lower bound developed below becomes particularly relevant.
If the total Chern number were non-zero, the parameter space would be a non-trivial fibre bundle, and the analysis would require working in local trivializations; the topological lower bound would then hold locally in each gauge patch. Crucially, even when the bundle is non-trivial ($c^{(\pm)}_{1}\neq 0$), the total space $M$ remains diffeomorphic to $T^{4}$, and all geometric structures—the metric $g_{M}$, the connection forms $\Theta^{(\pm)}$, and the orthogonal splitting $\mathcal{H}\oplus\mathcal{V}$—are globally defined. Therefore, our geometric analysis applies to both trivial and non-trivial bundles, provided the total Chern number vanishes. This precisely captures the physically important situation of compensated topological sectors.
Mathematical framework.
The geometric construction that follows requires only that:
  1. 1.
    The total space $M$ is diffeomorphic to $T^{4}$ (true for constant-curvature $T^{2}$-bundles over $T^{2}$).
  2. 2.
    The metric $g_{M}$ and connection forms $\Theta^{(\pm)}$ are globally defined on $M$.
  3. 3.
    The orthogonal splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$ induced by $g_{M}$ is global.
These conditions hold regardless of the triviality of the underlying principal bundle, as long as the curvatures are constant (Assumption 2.1). The vanishing total Chern number $c^{\rm tot}_{1}=0$ ensures the existence of a global diagonal section, which is sufficient for the physical interpretation but not required for the mathematical proofs.
Our goal in this section is to endow $M$ with a Riemannian structure that naturally incorporates the synthetic gauge fields of the SOC BEC, and to define on it a metric connection whose torsion three-form coincides with the physical Berry-curvature three-form.
2.1 Kaluza–Klein metric on the extended space
Let $g_{\mathrm{BZ}}=dk_{x}^{2}+dk_{y}^{2}$ be the flat metric on the Brillouin-zone torus $T^{2}_{\mathrm{BZ}}$. On each phase circle we introduce the standard angular coordinate with period $2\pi$; the canonical $U(1)$-invariant metric on $S^{1}$ is $d\phi^{2}$. The synthetic gauge fields of the SOC BEC provide, for each of the two independent $U(1)$ directions, a Berry connection. Denote by $A^{(+)},A^{(-)}\in\Omega^{1}(T^{2}_{\mathrm{BZ}})$ the corresponding real one-forms (the components of the synthetic $\mathfrak{su}(2)$ connection along the two independent $U(1)$ generators). Their curvatures are the Berry curvature two-forms $F^{(\pm)}=dA^{(\pm)}$.
Assumption 2.1 (Constant Berry curvatures).
In the following, we assume that the Berry curvature 2-forms are constant multiples of the Brillouin-zone volume form:
$F^{(\pm)}=c^{(\pm)}\,\mathrm{vol}_{\mathrm{BZ}},\qquad c^{(\pm)}\in\mathbb{R}\setminus\{0\},$
where $\mathrm{vol}_{\mathrm{BZ}}=dk_{x}\wedge dk_{y}$. The constants $c^{(\pm)}$ are related to the Chern numbers by $c_{1}^{(\pm)}=\frac{1}{2\pi}\int_{T^{2}}F^{(\pm)}=\frac{\operatorname{Area}(T^{2})}{2\pi}\,c^{(\pm)}$. This condition is satisfied in a wide range of experimentally relevant SOC BEC configurations, including linear Rashba–Dresselhaus couplings and uniform synthetic magnetic fields.
Following the Kaluza–Klein prescription we define a metric $g_{M}$ on $M$ that couples the base metric to the connection forms:
$g_{M}=\pi^{*}g_{\mathrm{BZ}}+\bigl(d\phi_{+}+\pi^{*}A^{(+)}\bigr)^{2}+\bigl(d\phi_{-}+\pi^{*}A^{(-)}\bigr)^{2},$ (2.1)
where $\pi:M\to T^{2}_{\mathrm{BZ}}$ is the projection onto the momentum coordinates. The one-forms
$\Theta^{(+)}:=d\phi_{+}+\pi^{*}A^{(+)},\qquad\Theta^{(-)}:=d\phi_{-}+\pi^{*}A^{(-)}$
are the horizontal lifts of the phase coordinates; they are globally defined on $M$ and are orthogonal to the base directions with respect to $g_{M}$.
Lemma 2.2 (Properties of the metric $g_{M}$).
The metric (2.1) satisfies the following.
  1. 1.
    $g_{M}$ is smooth, positive definite and makes $\pi:(M,g_{M})\to(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ a Riemannian submersion.
  2. 2.
    The fibres $\pi^{-1}(k)\cong S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ are flat $2$-tori with the induced metric $(d\phi_{+})^{2}+(d\phi_{-})^{2}$.
  3. 3.
    The horizontal distribution $\mathcal{H}=\ker(d\phi_{+}+\pi^{*}A^{(+)})\cap\ker(d\phi_{-}+\pi^{*}A^{(-)})$ is orthogonal to the vertical distribution $\mathcal{V}=\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$.
  4. 4.
    If $A^{(\pm)}=0$ then $g_{M}$ reduces to the trivial product metric $g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$.
Proof.
Smoothness follows from the smoothness of $g_{\mathrm{BZ}}$ and $A^{(\pm)}$. Positive definiteness is clear because each term in (2.1) is non‑negative and their sum vanishes only when all differentials vanish.
For any $k\in T^{2}_{\mathrm{BZ}}$ the restriction of $g_{M}$ to the fibre $\pi^{-1}(k)$ is $(d\phi_{+})^{2}+(d\phi_{-})^{2}$, independent of $k$; hence the fibres are isometric to a flat $2$-torus. The map $\pi$ is a Riemannian submersion because, by construction, $g_{M}$ restricted to $\mathcal{H}$ coincides with $\pi^{*}g_{\mathrm{BZ}}$.
Orthogonality $\mathcal{H}\perp\mathcal{V}$ follows directly from $g_{M}(\partial_{\phi_{\pm}},X)=0$ for every $X\in\mathcal{H}$, which is a consequence of the definition of $\Theta^{(\pm)}$.
The last statement is immediate from (2.1). ∎
Remark 2.3.
Equation (2.1) is physically natural: it measures distances in the extended parameter space by combining the kinetic contribution from momentum changes with the “phase cost” induced by the synthetic gauge potentials. If $A^{(\pm)}=0$, the metric reduces to the trivial product metric $g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$.
2.2 Metric connection with prescribed torsion
On the Riemannian manifold $(M,g_{M})$ we now construct a metric connection $\nabla^{C}$ whose torsion three-form is built from the physical gauge fields. Let $\nabla^{LC}$ be the Levi‑Civita connection of $g_{M}$.
Define the three-form
$T^{b}:=F^{(+)}\wedge\Theta^{(+)}+F^{(-)}\wedge\Theta^{(-)}\in\Omega^{3}(M),$
where $\Theta^{(\pm)}=d\phi_{\pm}+\pi^{*}A^{(\pm)}$ are the horizontal 1-forms defined in (2.1). Note that $F^{(\pm)}\wedge A^{(\pm)}=0$ for dimensional reasons, since both forms are pullbacks from the 2-dimensional base $T^{2}_{\mathrm{BZ}}$. $T^{b}$ is the trace $\operatorname{Tr}(\mathcal{F}\wedge\boldsymbol{\Theta})$ with respect to the Killing form on $\mathfrak{u}(1)_{+}\oplus\mathfrak{u}(1)_{-}$, where $\boldsymbol{\Theta}=\Theta^{(+)}\otimes t_{+}+\Theta^{(-)}\otimes t_{-}$ is the total connection form on $M$ and $\mathcal{F}=F^{(+)}\otimes t_{+}+F^{(-)}\otimes t_{-}$.
Introduce the (2,1)(2,1)‑tensor $K$ by
$K(X,Y,Z):=\frac{1}{2}\,T^{b}(X,Y,Z)\qquad(X,Y,Z\in\mathfrak{X}(M)).$
The connection
$\nabla^{C}_{X}Y:=\nabla^{LC}_{X}Y+K(X,Y,\cdot)^{\sharp}$
is by construction metric ($\nabla^{C}g_{M}=0$) and has totally skew‑symmetric torsion $T^{b}$. Indeed, for any vector fields $X,Y$,
$T^{b}(X,Y,\cdot)^{\sharp}=\nabla^{C}_{X}Y-\nabla^{C}_{Y}X-[X,Y],$
which is precisely the definition of the torsion tensor of $\nabla^{C}$.
Because $T^{b}$ is not parallel with respect to $\nabla^{LC}$ unless the gauge fields are flat, the connection $\nabla^{C}$ satisfies the non‑parallelism condition $\nabla^{LC}T^{b}\neq 0$ required for the applicability of the general theorem [12].
Remark 2.4.
The three‑form $T^{b}$ is not closed in general; its exterior derivative contains terms quadratic in the curvatures $F^{(\pm)}$. Its de Rham cohomology class $T^{b}]\in H^{3}(M;\mathbb{R})$, however, depends only on the harmonic part of $T^{b}$ with respect to $g_{M}$, as will be computed in Section 3.
2.3 Parallel forms in the Kaluza–Klein geometry
A key ingredient of the topological bound is the space of parallel forms on the factors. For the base torus with the flat metric $g_{\mathrm{BZ}}$ we have the well‑known result
$\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})=H^{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}},\qquad\mathrm{vol}_{\mathrm{BZ}}=dk_{x}\wedge dk_{y}.$
The situation for the phase circles is more subtle. On a circle endowed with the standard metric $d\phi^{2}$, the one‑form $d\phi$ is parallel, hence $\mathcal{P}_{1}(S^{1},d\phi^{2})=\mathbb{R}\,d\phi$. In the Kaluza–Klein metric (2.1) the geometry of the fibres is “twisted” by the connection forms $A^{(\pm)}$. The vector fields $\partial_{\phi_{+}},\partial_{\phi_{-}}$ are Killing fields of $g_{M}$, but their dual one‑forms are $\Theta^{(\pm)}$, not $d\phi_{\pm}$.
Lemma 2.5 (Triviality of the obstruction kernel $\mathcal{K}$ under constant curvature).
Let $g_{M}$ be the Kaluza–Klein metric (2.1) on $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, and let Assumption 2.1 hold. Then the obstruction kernel $\mathcal{K}$ vanishes identically:
$\mathcal{K}\;:=\;\bigl(\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))\bigr)\oplus\bigl(\mathcal{V}_{1,2}\cap(\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))\bigr)\;=\;\{0\},$
where $\mathcal{V}_{2,1}$ and $\mathcal{V}_{1,2}$ are the spans of the harmonic representatives of the mixed Künneth components of the torsion class $\omega]$.
Proof.
Under Assumption 2.1, the cohomology classes of the Berry curvatures satisfy $F^{(-)}]=\lambda[F^{(+)}]$ with $\lambda=c^{(-)}/c^{(+)}$. Consequently, the mixed part of the torsion class is represented by the simple tensor
$\xi_{0}=\operatorname{vol}_{\mathrm{BZ}}\otimes\bigl(d\phi_{+}+\lambda d\phi_{-}\bigr)\in\mathcal{V}_{2,1}.$
Thus $\mathcal{V}_{2,1}=\operatorname{span}\{\xi_{0}\}$.
The space of parallel $1$-forms on the fibre, $\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$, consists of those linear combinations $\eta=c_{+}d\phi_{+}+c_{-}d\phi_{-}$ for which the corresponding vertical $1$-form $c_{+}\Theta^{(+)}+c_{-}\Theta^{(-)}$ is $g_{M}$-parallel. A direct computation (provided in Appendix A) shows that such an $\eta$ is parallel if and only if the constants satisfy the linear relation
$c_{+}+\lambda c_{-}=0.$
Now, an element of $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$ has the form
$\operatorname{vol}_{\mathrm{BZ}}\otimes(c_{+}d\phi_{+}+c_{-}d\phi_{-})\qquad\text{with}\qquad c_{+}+\lambda c_{-}=0.$
For this element to belong to $\mathcal{V}_{2,1}$, it must be proportional to $\xi_{0}$; hence there exists $t\in\mathbb{R}$ such that
$c_{+}=t,\qquad c_{-}=t\lambda.$
Substituting into condition (*) gives
$t+\lambda(t\lambda)=t(1+\lambda^{2})=0,$
which, for real $\lambda$, forces $t=0$ and therefore $c_{+}=c_{-}=0$. Hence the intersection $\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}))$ is trivial. An analogous argument shows that the $(1,2)$-component intersection is also trivial, completing the proof. ∎
Remark 2.6 (Physical relevance of constant Berry curvatures).
Assumption 2.1 covers a wide range of experimentally relevant SOC BEC configurations. In Raman-induced spin–orbit coupling, the Berry curvature often becomes constant (or piecewise constant) in large regions of the Brillouin zone. Moreover, even when the curvature is not strictly constant, the topological lower bound depends only on the cohomology class $F^{(\pm)}]$, not on its pointwise values. One can always deform the connection within its cohomology class to a connection with constant curvature without changing the topological data that enter the bound. Therefore, the results derived under this assumption capture the essential topological obstruction present in generic SOC BECs.
2.4 Summary of the geometric setup
We have constructed:
  1. 1.
    an extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$;
  2. 2.
    a Kaluza–Klein metric $g_{M}$ on $M$ given by (2.1), coupling the momentum directions to the two independent $U(1)$ phases via the synthetic Berry connections $A^{(\pm)}$;
  3. 3.
    a metric connection $\nabla^{C}$ on $(M,g_{M})$ whose torsion three‑form is the physical quantity $T^{b}=F^{(+)}\wedge A^{(+)}+F^{(-)}\wedge A^{(-)}$.
This construction provides the precise Riemannian framework needed to apply the general theorem of [12]. In the following section we analyze the cohomological data that arise from this geometry.
3 Cohomology of the Extended Space and the Mixed Tensor Rank
In the following analysis, we work under Assumption 2.1 (constant Berry curvatures). With the geometric structure $(M,g_{M},\nabla^{C})$ established in Section 2, we now analyze the cohomological data that enter the general lower bound. Recall that the torsion three-form $T^{\flat}\in\Omega^{3}(M)$ is not necessarily closed, but its de Rham cohomology class $T^{\flat}]\in H^{3}(M;\mathbb{R})$ is well defined. Let $\omega_{h}$ be the unique $g_{M}$-harmonic representative of this class; then
$\omega]:=[\omega_{h}]=[T^{\flat}]\in H^{3}(M;\mathbb{R}).$
3.1 Künneth decomposition of $H^{3}(M)$
Since $M$ is a product $T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ and each factor is oriented and compact, the Künneth theorem gives an isomorphism
$H^{3}(M;\mathbb{R})\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
because $H^{3}(T^{2}_{\mathrm{BZ}})=H^{2}(S^{1}_{\phi_{\pm}})=0$ and $H^{1}(T^{2}_{\mathrm{BZ}})$ does not contribute to degree $3$. Consequently $\dim H^{3}(M)=2$, and every class $\eta]\in H^{3}(M)$ can be written uniquely as
$\eta]=[\alpha_{+}]\otimes[d\phi_{+}]+[\alpha_{-}]\otimes[d\phi_{-}],$ (3.1)
where $\alpha_{\pm}]\in H^{2}(T^{2}_{\mathrm{BZ}})$ and $d\phi_{\pm}]\in H^{1}(S^{1}_{\phi_{\pm}})$ are the fundamental cohomology classes of the factors.
Lemma 3.1 (Explicit harmonic representative).
Under Assumption 2.1, the unique $g_{M}$-harmonic representative of the torsion class $\omega]$ is given by
$\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge\bigl(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)}\bigr),$
where $a_{\pm}$ are real constants satisfying $a_{-}=\lambda a_{+}$ with $\lambda=c^{(-)}/c^{(+)}$.
Proof.
Let $\eta=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)})$ with $a_{-}=\lambda a_{+}$.
Closedness:
$d\eta=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}d\Theta^{(+)}+a_{-}d\Theta^{(-)})=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}F^{(+)}+a_{-}F^{(-)})=(a_{+}c^{(+)}+a_{-}c^{(-)})\operatorname{vol}_{\mathrm{BZ}}\wedge\operatorname{vol}_{\mathrm{BZ}}=0.$
Coclosedness: In the orthonormal coframe $(dx,dy,\Theta^{(+)},\Theta^{(-)})$,
$\eta=a_{+}\Theta^{(-)}-a_{-}\Theta^{(+)},\quad d(*\eta)=a_{+}F^{(-)}-a_{-}F^{(+)}=(a_{+}c^{(-)}-a_{-}c^{(+)})\operatorname{vol}_{\mathrm{BZ}}.$
Since $a_{-}=\lambda a_{+}$ and $c^{(-)}=\lambda c^{(+)}$, we have $a_{+}c^{(-)}-a_{-}c^{(+)}=0$, hence $\delta\eta=*d(*\eta)=0$.
Thus $\eta$ is harmonic. Uniqueness follows from the Hodge theorem. ∎
Remark 3.2.
The representatives $d\phi_{\pm}$ are harmonic with respect to the standard metric on the circles, but not with respect to $g_{M}$. The $g_{M}$-harmonic representatives are the horizontal forms $\Theta^{(\pm)}$. Since $d\phi_{\pm}]=[\Theta^{(\pm)}]$ in $H^{1}(S^{1}_{\phi_{\pm}})$, the cohomological decomposition (3.1) remains valid when we replace $d\phi_{\pm}$ by $\Theta^{(\pm)}$.
3.2 Mixed tensor rank of the torsion class
The mixed tensor rank of a cohomology class $\eta]\in H^{3}(M)$ written as in (3.1) is defined as the minimal number of simple tensors required to represent it:
$\operatorname{rank}_{\mathbb{R}}([\eta]):=\min\Bigl\{m\;\Big|\;[\eta]=\sum_{i=1}^{m}[\beta_{i}]\otimes[\gamma_{i}],\;[\beta_{i}]\in H^{2}(T^{2}_{\mathrm{BZ}}),\;[\gamma_{i}]\in H^{1}(S^{1}_{\phi_{+}})\oplus H^{1}(S^{1}_{\phi_{-}})\Bigr\}.$
For the class $\omega]$ coming from the physical torsion, the two-forms $\alpha_{+},\alpha_{-}$ in (3.1) are precisely the (cohomology classes of the) Berry curvature two-forms $F^{(+)},F^{(-)}$ of the two independent $U(1)$ sectors. A key observation is that $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$. Therefore, any two non-zero classes in this space are necessarily cohomologically proportional. Under Assumption 2.1, we have
$F^{(-)}]=\lambda[F^{(+)}]\quad\text{in }H^{2}(T^{2}_{\mathrm{BZ}}),\qquad\lambda=c^{(-)}/c^{(+)}.$
Substituting this relation into (3.1) gives
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+\lambda[F^{(+)}]\otimes[d\phi_{-}]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
which is a single simple tensor. Hence the mixed tensor rank attains the value
$r=1.$
Remark 3.3.
Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the maximal possible mixed tensor rank for a class of the form $\alpha_{+}]\otimes[d\phi_{+}]+[\alpha_{-}]\otimes[d\phi_{-}]$ with both $\alpha_{\pm}]\neq 0$ is $1$. All experimentally prominent SOC geometries (Rashba–Dresselhaus, Hofstadter, Raman-induced couplings) satisfy the condition $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$, thereby realizing the generic case $r=1$.
Remark 3.4.
Since $M_{2}=S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ is a product, its first cohomology decomposes as $H^{1}(M_{2})\cong H^{1}(S^{1}_{+})\oplus H^{1}(S^{1}_{-})$. Therefore an element of $H^{2}(M_{1})\otimes H^{1}(M_{2})$ can be written as a sum of terms of the form $\alpha]\otimes[d\phi_{+}]+[\beta]\otimes[d\phi_{-}]$. The mixed tensor rank $r$ is then the minimal number of simple tensors needed to represent this sum in the space $H^{2}(M_{1})\otimes\bigl(H^{1}(S^{1}_{+})\oplus H^{1}(S^{1}_{-})\bigr)$.
3.3 Vanishing of the obstruction kernel $\mathcal{K}$
The general bound of [12] involves a correction term $\dim\mathcal{K}$, where
$\mathcal{K}=\ker\Psi_{p}\oplus\ker\widetilde{\Psi}_{p},\qquad\ker\Psi_{p}=\pi_{2,1}([\omega])\cap\bigl(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}})\bigr),$
and $\ker\widetilde{\Psi}_{p}$ is defined analogously with the roles of the two circles exchanged. Here $\pi_{2,1}$ denotes the projection onto the $(2,1)$-Künneth component.
In our geometry, we have $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}}$ (the space of parallel two-forms on the flat torus). The obstruction kernel $\mathcal{K}$ is shown to vanish by Lemma 2.5 under Assumption 2.1. The proof of Lemma 2.5 demonstrates that no non-zero element of the form $\operatorname{vol}_{\mathrm{BZ}}\otimes(c_{+}d\phi_{+}+c_{-}d\phi_{-})$ with $c_{+}+\lambda c_{-}=0$ can belong to $\mathcal{V}_{2,1}=\operatorname{span}\{\operatorname{vol}_{\mathrm{BZ}}\otimes(d\phi_{+}+\lambda d\phi_{-})\}$.
Thus we have
$\mathcal{K}=0,\qquad\text{and therefore}\qquad\dim\mathcal{K}=0.$ (3.2)
3.4 Summary of the topological data
Based on (3.2) we obtain the reduced rank
$r^{\sharp}\;:=\;r-\dim\mathcal{K}\;=\;1.$
All ingredients required for the application of the general geometric theorem are now in place: the extended parameter space $(M,g_{M})$ carries a metric connection $\nabla^{C}$ with totally skew torsion whose harmonic part represents a mixed cohomology class $\omega]$ of mixed tensor rank $r=1$, and the obstruction kernel $\mathcal{K}$ vanishes identically. In the next section, we apply the theorem to obtain a concrete lower bound on the off-diagonal holonomy of $\nabla^{C}$.
4 Geometric foundations for the PT theorem in Kaluza–Klein geometry
The general topological lower bound of Pigazzini and Toda [12] is stated for genuine Riemannian product manifolds. Our physical metric (2.1), however, is a Kaluza–Klein metric that does not globally factor as a product. We now explain why, despite this apparent mismatch, the bound applies to our specific geometry. The argument relies on three observations that hold for the particular Kaluza–Klein structure arising from a two–component SOC BEC.
Cohomology class and its harmonic representative
The torsion 3–form $T^{b}=F^{(+)}\wedge\Theta^{(+)}+F^{(-)}\wedge\Theta^{(-)}$ need not be closed, but its de Rham class $\omega]=[T^{b}]\in H^{3}(M;\mathbb{R})$ is well defined. (Recall that $\Theta^{(\pm)}=d\phi_{\pm}+\pi^{*}A^{(\pm)}$; note that the pullback $\pi^{*}\bigl(F^{(\pm)}\wedge A^{(\pm)}\bigr)$ vanishes because on the 2-dimensional base $T^{2}_{\mathrm{BZ}}$ we have $F^{(\pm)}\wedge A^{(\pm)}=0$. Therefore, $T^{b}$ is cohomologous to $F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}$.) Under the Künneth decomposition
$H^{3}(M)\cong\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}})\bigr)\oplus\bigl(H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{-}})\bigr),$
the class $\omega]$ can be written as
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}].$
Because $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the two Berry–curvature classes are necessarily proportional; hence there exists $\lambda\in\mathbb{R}$ such that $F^{(-)}]=\lambda[F^{(+)}]$. Consequently,
$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
which is a single simple tensor. The mixed tensor rank is therefore $r=1$, a purely topological invariant.
Let $\omega_{h}$ be the unique $g_{M}$-harmonic representative of $\omega]$. Under Assumption 2.1, Lemma 3.1 gives an explicit form:
$\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge\bigl(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)}\bigr),$
where $a_{\pm}$ are constants with $a_{-}=\lambda a_{+}$ and $\lambda=c^{(-)}/c^{(+)}$. This factorised form is the exact analogue, for our Kaluza–Klein metric, of the factorised harmonic decomposition used in the product-metric case.
Spaces of parallel forms
The bound involves the obstruction kernel
$\mathcal{K}=\bigl(\mathcal{V}_{2,1}\cap(\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{\phi_{+}}))\bigr)\oplus\bigl(\mathcal{V}_{1,2}\cap(\mathcal{P}_{1}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{2}(S^{1}_{\phi_{-}}))\bigr).$
For the flat base we have $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})=\mathbb{R}\operatorname{vol}_{\mathrm{BZ}}$. By Lemma 2.5, the obstruction kernel vanishes: $\mathcal{K}=0$.
Pointwise construction of the obstruction maps
Even though the metric is not a product, at each point $p$ we have an orthogonal splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$. The non-integrability of $\mathcal{H}$ (measured by the Berry curvatures $F^{(\pm)}$) does not affect pointwise evaluations. Consider the (2,1)–component. Using the explicit harmonic representative from Lemma 3.1, we set
$\mathcal{V}_{2,1}:=\operatorname{span}\{\operatorname{vol}_{\mathrm{BZ}}\otimes(a_{+}d\phi_{+}+a_{-}d\phi_{-})\}\subset H^{2}(T^{2}_{\mathrm{BZ}})\otimes H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}).$
At a point $p=(k,\phi_{+},\phi_{-})$ choose any vectors $X,Y\in\mathcal{H}_{p}$ with $\operatorname{vol}_{\mathrm{BZ}}(X,Y)=1$ and a vertical vector $Z\in\mathcal{V}_{p}$ such that $(a_{+}d\phi_{+}+a_{-}d\phi_{-})(Z)=1$. Define
$\Psi_{p}\bigl(\operatorname{vol}_{\mathrm{BZ}}\otimes(a_{+}d\phi_{+}+a_{-}d\phi_{-})\bigr):=\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}(X,Y,Z),$
where $(\cdot)_{\mathrm{off}}$ denotes the part mapping $\mathcal{H}$ to $\mathcal{V}$ or vice versa. The map is well-defined because any other choice $(X^{\prime},Y^{\prime},Z^{\prime})$ satisfying the same conditions gives the same operator up to a constant factor, due to the linearity of the curvature in its arguments.
A direct computation (provided in Appendix B) shows that for $\xi=\operatorname{vol}_{\mathrm{BZ}}\otimes(a_{+}d\phi_{+}+a_{-}d\phi_{-})\in\mathcal{V}_{2,1}$, we have:
$\Psi_{p}(\xi)=\frac{1}{2}\left(a_{+}F^{(+)}(\pi(p))+a_{-}F^{(-)}(\pi(p))\right)J,$
where $J:T_{p}M\to T_{p}M$ is a fixed off-diagonal operator (independent of $a_{\pm}$) that exchanges horizontal and vertical directions. The calculation uses:
  1. 1.
    $\nabla^{LC}(\operatorname{vol}_{\mathrm{BZ}}\wedge\Theta^{(\pm)})=\operatorname{vol}_{\mathrm{BZ}}\wedge\pi^{*}F^{(\pm)}$ (since $\operatorname{vol}_{\mathrm{BZ}}$ is parallel on the flat base);
  2. 2.
    the definition of $\Psi_{p}$ in terms of $(\nabla^{LC}T)_{\mathrm{off}}$;
  3. 3.
    the choice of vectors $X,Y\in\mathcal{H}_{p}$ with $\operatorname{vol}_{\mathrm{BZ}}(X,Y)=1$ and $Z\in\mathcal{V}_{p}$ with $(a_{+}d\phi_{+}+a_{-}d\phi_{-})(Z)=1$.
Therefore, $\Psi_{p}(\xi)=0$ if and only if $a_{+}F^{(+)}(\pi(p))+a_{-}F^{(-)}(\pi(p))=0$. Under Assumption 2.1, $F^{(-)}=\lambda F^{(+)}$ with $\lambda=c^{(-)}/c^{(+)}$. Moreover, since $\xi\in\mathcal{V}_{2,1}$, the coefficients satisfy $a_{-}=\lambda a_{+}$ (because $\mathcal{V}_{2,1}$ is spanned by $\operatorname{vol}_{\mathrm{BZ}}\otimes(d\phi_{+}+\lambda d\phi_{-})$). Substituting gives
$a_{+}F^{(+)}+a_{-}F^{(-)}=a_{+}F^{(+)}+\lambda a_{+}(\lambda F^{(+)})=a_{+}(1+\lambda^{2})F^{(+)}.$
Since $F^{(+)}\neq 0$, this vanishes precisely when $a_{+}=0$, which also implies $a_{-}=0$. Hence $\ker\Psi_{p}=\{0\}$ and $\operatorname{im}\Psi_{p}$ is one-dimensional. The same argument applied to the (1,2)–component gives a map $\widetilde{\Psi}_{p}$ with trivial kernel.
Because the images of $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ consist of off–diagonal curvature operators, the Ambrose–Singer theorem forces
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\;\geq\;\dim\operatorname{im}\Psi_{p}+\dim\operatorname{im}\widetilde{\Psi}_{p}=1+0=1.$
Thus the topological lower bound holds directly for the physical Kaluza–Klein metric $g_{M}$. The deformation family introduced below will serve only to connect this result to the product–metric case and to make the semicontinuity argument transparent.
Note 1: The reasoning above does not require the metric to be a global Riemannian product. It uses only the pointwise orthogonal splitting $T_{p}M=\mathcal{H}_{p}\oplus\mathcal{V}_{p}$, the explicit description of the harmonic representative (Lemma 3.1, which relies on the flatness of the base and the fibres under Assumption 2.1), and the vanishing of $\mathcal{P}_{1}$ on the fibres (Lemma 2.5). All these ingredients are present in the SOC BEC geometry.
Deformation to a product metric and stability
To provide a transparent link with the original product–metric setting, and to control the behaviour of the holonomy dimension via lower–semicontinuity, we now introduce a smooth deformation that connects the physical Kaluza–Klein metric to a genuine product metric.
Note 2: [On the wedge decomposition in Kaluza–Klein geometry] In the product case, the harmonic representative decomposes as a sum of wedge products of harmonic forms on the factors. In our Kaluza–Klein geometry, the analogous decomposition (Lemma 3.1) involves the horizontal forms $\Theta^{(\pm)}$, which are harmonic with respect to the full metric $g_{M}$ but are not pullbacks from the fibres alone. However, they satisfy:
$\Delta_{g_{M}}\Theta^{(\pm)}=0\quad\text{and}\quad[\Theta^{(\pm)}]=[d\phi_{\pm}]\in H^{1}(S^{1}_{\phi_{\pm}}),$
which is sufficient for the cohomological arguments. The key point is that $\operatorname{vol}_{\mathrm{BZ}}\wedge\Theta^{(\pm)}$ represents the mixed class $F^{(\pm)}]\otimes[d\phi_{\pm}]$ in the Künneth decomposition, regardless of the non-product nature of the metric.
4.1 Deformation to a product metric and stability
The physical Kaluza–Klein metric (2.1) is a warped structure that does not admit a global Riemannian product decomposition, whereas Theorem 5.2 in [12] is formally stated for genuine product metrics. To bridge this gap, we introduce a smooth deformation to a product metric and prove the stability of the holonomy bound.
Definition 4.1 (Deformation family).
Let $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ be the smooth one-parameter family of metrics on $M$ defined by:
$g_{\varepsilon}=\pi^{*}g_{\mathrm{BZ}}+(d\phi_{+}+\varepsilon\pi^{*}A^{(+)})^{2}+(d\phi_{-}+\varepsilon\pi^{*}A^{(-)})^{2}.$
For $\varepsilon=0$ we obtain the genuine Riemannian product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$, while for $\varepsilon=1$ we recover the physical Kaluza–Klein metric $g_{M}=g_{1}$.
Lemma 4.2 (Stability of the obstruction data).
For the family $g_{\varepsilon}$ in Definition 4.1, the following hold:
  1. 1.
    The spaces of parallel forms satisfy:
    $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\varepsilon})=\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{0})=\mathbb{R}\operatorname{vol}_{\mathrm{BZ}}\quad\text{for all }\varepsilon\in[0,1].$
  2. 2.
    The mixed tensor rank $r=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ is constant for all $\varepsilon$.
  3. 3.
    The obstruction kernel $\mathcal{K}_{\varepsilon}$ satisfies:
    • (a)
      $\dim\mathcal{K}_{0}=1$.
    • (b)
      $\dim\mathcal{K}_{\varepsilon}=0$ for every $\varepsilon>0$.
    Consequently, the corrected rank $r^{\sharp}_{\varepsilon}=r-\dim\mathcal{K}_{\varepsilon}$ equals 0 for $\varepsilon=0$ and equals $1$ for every $\varepsilon>0$.
Proof.
(1) follows because the restriction of $g_{\varepsilon}$ to $T^{2}_{\mathrm{BZ}}$ is $g_{\mathrm{BZ}}$, independent of $\varepsilon$.
(2) The mixed tensor rank is a topological invariant (Lemma 4.7) and therefore constant under metric deformations.
(3a) For the product metric $g_{0}=g_{\mathrm{BZ}}\oplus d\phi_{+}^{2}\oplus d\phi_{-}^{2}$, the Brillouin torus is flat. On a flat Riemannian manifold every harmonic form is parallel; thus $\mathcal{H}^{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})=\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$. The harmonic representative of the torsion class $\omega]$ with respect to $g_{0}$ is
$\omega_{h}^{(0)}=\alpha_{+}\wedge d\phi_{+}+\alpha_{-}\wedge d\phi_{-},$
where $\alpha_{\pm}$ are the harmonic representatives of the Berry‑curvature classes $F^{(\pm)}]$. Since $\alpha_{\pm}$ are parallel and $d\phi_{\pm}$ are parallel, $\omega_{h}^{(0)}$ belongs to $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}})\otimes\mathcal{P}_{1}(S^{1}_{+}\times S^{1}_{-})$. Hence $\mathcal{K}_{0}=\mathcal{V}_{2,1}$ and $\dim\mathcal{K}_{0}=1$.
(3b) For $\varepsilon>0$ the metric $g_{\varepsilon}$ is a Kaluza–Klein metric satisfying Assumption 2.1 (with curvatures scaled by $\varepsilon$). By Lemma 2.5, the obstruction kernel vanishes: $\mathcal{K}_{\varepsilon}=0$.
Thus $r^{\sharp}_{\varepsilon}=r-\dim\mathcal{K}_{\varepsilon}$ equals 0 for $\varepsilon=0$ and equals $1$ for every $\varepsilon>0$. ∎
Lemma 4.3 (Lower semi-continuity of off-diagonal holonomy dimension).
The function $F(\varepsilon):=\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ is lower semi-continuous on $[0,1]$, where $\nabla^{C_{\varepsilon}}$ is the metric connection with fixed torsion $T$ constructed from $g_{\varepsilon}$.
Proof.
The curvature tensor $R^{C_{\varepsilon}}$ depends smoothly on $\varepsilon$. At any point $p\in M$, consider the evaluation map
$\Phi_{\varepsilon}:\bigwedge^{2}T_{p}M\times T_{p}M\to\mathfrak{so}(T_{p}M),\quad\Phi_{\varepsilon}(X,Y,Z)=R^{C_{\varepsilon}}(X,Y)Z.$
The off-diagonal holonomy algebra at $p$ is generated by the image of $\Phi_{\varepsilon}$ restricted to arguments that yield off-diagonal operators. The dimension of a finitely generated subspace is a lower semi-continuous function of the generating set (it can only jump upward at discontinuity points). Since $\Phi_{\varepsilon}$ varies continuously in $\varepsilon$, $F(\varepsilon)$ is lower semi-continuous. ∎
Proposition 4.4 (Persistence of the PT bound under deformation).
If the Pigazzini–Toda theorem gives $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{0}})\geq r^{\sharp}$ for the product metric $g_{0}$, then the same inequality holds for all $\varepsilon\in[0,1]$, and in particular for the physical metric $g_{M}=g_{1}$.
Proof.
By Lemma 4.2, the corrected rank $r^{\sharp}_{\varepsilon}=r-\dim\mathcal{K}_{\varepsilon}$ is constant on $(0,1]$ and equals $1$. For every $\varepsilon>0$, Theorem 4.8 (the extension of the PT theorem to Kaluza–Klein metrics) therefore yields the non‑trivial lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1$.
The function $\varepsilon\mapsto\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ is lower‑semicontinuous by Lemma 4.3; hence the set
$S=\{\varepsilon\in[0,1]:\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1\}$
is closed. Since $S$ contains the interval $(0,1]$, its closure contains the endpoints $\varepsilon=0$ and $\varepsilon=1$. Because $S$ is closed, we must have $0\in S$ and $1\in S$. In particular, $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{1}})\geq 1$, which is the desired bound for the physical metric. ∎
4.2 Persistence of non‑parallelism under metric change
To ensure that the lower bound on the off-diagonal holonomy is a robust feature of the model rather than a consequence of a specific metric choice, we must examine the stability of the non-parallelism condition $\nabla^{LC}T\neq 0$. The following result establishes that this property is open in the space of Riemannian metrics, guaranteeing that the topological obstruction remains effective under generic deformations of the Kaluza–Klein geometry.
Lemma 4.5 (Openness of the non‑parallelism condition).
Let $M$ be a smooth manifold and $T\in\Omega^{3}(M)$ a fixed 3‑form. Denote by $\mathcal{M}$ the space of Riemannian metrics on $M$ endowed with the $C^{1}$-topology. The set
$\mathcal{M}_{\neq 0}(T):=\bigl\{g\in\mathcal{M}\mid\nabla^{LC}(g)T\neq 0\bigr\}$
is open in $\mathcal{M}$. In particular, if $\nabla^{LC}(g)T\neq 0$ for some metric $g$, then there exists a $C^{1}$-neighbourhood of $g$ consisting of metrics for which the covariant derivative of $T$ is also non‑zero.
Proof.
The Levi‑Civita connection $\nabla^{LC}(g)$ depends smoothly on the components of the metric $g$ and its first derivatives through the Christoffel symbols. Consequently, the map $g\mapsto\nabla^{LC}(g)T$ is continuous with respect to the $C^{1}$-topology on $\mathcal{M}$. Since the condition $\nabla^{LC}(g)T\neq 0$ defines the complement of the zero set of a continuous map, $\mathcal{M}_{\neq 0}(T)$ is an open subset of $\mathcal{M}$. ∎
Remark 4.6.
In the context of spin-orbit coupled BECs, Lemma 4.5 ensures the structural stability of our topological bound. Since the Kaluza–Klein metric $g$ satisfies $\nabla^{LC}(g)T\neq 0$ due to the non-triviality of the synthetic gauge curvature, any sufficiently small perturbation of the metric—arising, for instance, from experimental fluctuations in the trapping potentials or the SOC strength—will preserve the non-parallelism of the torsion. This confirms that the lower bound on the off-diagonal holonomy is not a fine-tuned property but a robust feature of the SOC BEC geometry.
4.3 Harmonic representatives in Kaluza–Klein geometry
Lemma 4.7 (Metric invariance of the mixed tensor rank).
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold. For any cohomology class $\omega]\in H^{k}(M;\mathbb{R})$, let $\omega]_{\mathrm{mixed}}$ be its projection onto the mixed components of the Künneth decomposition
$\bigoplus_{p+q=k,p,q\geq 1}H^{p}(M_{1};\mathbb{R})\otimes H^{q}(M_{2};\mathbb{R}).$
The mixed tensor rank $r:=\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})$ is a topological invariant of the class $\omega]$. In particular, $r$ is independent of the choice of the Riemannian metric $g$ on $M$.
Proof.
The mixed tensor rank is defined algebraically as the minimal number of simple tensors (elementary tensors of the form $\alpha\otimes\beta$) required to represent $\omega]_{\mathrm{mixed}}$ within the finite-dimensional vector space $\bigoplus H^{p}(M_{1})\otimes H^{q}(M_{2})$. Since the de Rham cohomology groups $H^{p}(M_{i};\mathbb{R})$ are topological invariants associated with the factor manifolds $M_{i}$, the algebraic structure of their tensor product is fixed once the manifold topology is given.
Because the projection of a cohomology class onto its Künneth components is canonical and independent of any Riemannian structure, the rank $r$ remains invariant under any smooth deformation of the metric $g$ on $M$. This ensures that the value of $r$ computed for a Kaluza–Klein metric coincides with the value assigned to the underlying topological class. ∎
4.4 The PT bound for Kaluza–Klein metrics with flat fibres
With the previous preparatory lemmas, we can now state and prove the extension of the Pigazzini-Toda theorem to Kaluza-Klein geometries.
Theorem 4.8 (PT bound for the SOC BEC geometry).
Let $M=T_{\mathrm{BZ}}^{2}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ be equipped with the Kaluza-Klein metric $g_{M}$ as in (2.1), and let $\nabla^{C}=\nabla^{LC}+\frac{1}{2}T$ be the metric connection with torsion $T=F^{(+)}\wedge\Theta^{(+)}+F^{(-)}\wedge\Theta^{(-)}$. Assume that $\nabla^{LC}T\neq 0$ and that the Berry curvature classes satisfy $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}})$. Then the off-diagonal holonomy of $\nabla^{C}$ satisfies
$\dim\mathfrak{hol}^{\mathrm{off}}_{p}(\nabla^{C})\geq 1\quad\text{for every }p\in M.$
Proof for the SOC BEC geometry.
We prove the bound via deformation and lower semi-continuity, without requiring global integrability of the horizontal distribution.
Consider the family $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ from Definition 4.1. For each $\varepsilon>0$, the orthogonal splitting $T_{p}M=\mathcal{H}^{(\varepsilon)}_{p}\oplus\mathcal{V}_{p}$ (horizontal/vertical) is pointwise well-defined. Although for $\varepsilon>0$ this splitting is not integrable unless $F^{(\pm)}=0$, the original Pigazzini–Toda theorem [12] uses only the pointwise orthogonal decomposition of the tangent space, not its integrability. Indeed, the proof of Theorem 5.2 in [12] relies solely on algebraic computations with curvature operators and on the properties of harmonic forms, which remain valid for any orthogonal splitting.
For $\varepsilon>0$, Lemma 2.5 implies that the obstruction kernel vanishes: $\mathcal{K}_{\varepsilon}=0$. The mixed tensor rank $r=1$ is topological (Lemma 4.7), so $r^{\sharp}_{\varepsilon}=r-\dim\mathcal{K}_{\varepsilon}=1$.
Applying Theorem 5.2 of [12] to the splitting $\mathcal{H}^{(\varepsilon)}\oplus\mathcal{V}$ yields
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq r^{\sharp}_{\varepsilon}=1\quad\forall\varepsilon>0,$
where $\nabla^{C_{\varepsilon}}$ is the connection built from $g_{\varepsilon}$ with the same torsion form $T$. This application is justified because Theorem 5.2 of [12] requires only a pointwise orthogonal decomposition $T_{p}M=V_{1}\oplus V_{2}$ and the existence of the harmonic representative $\omega_{h}$ with the stated properties. The non-integrability of $\mathcal{H}^{(\varepsilon)}$ (when $\varepsilon>0$) does not affect the proof, which is based on pointwise algebraic computations with curvature operators.
Define $F(\varepsilon):=\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ for $\varepsilon\in[0,1]$. Then $F$ is integer-valued (a dimension) and, by Lemma 4.3, lower semi-continuous on $[0,1]$.
We now prove $F(1)\geq 1$. For an integer-valued lower semi-continuous function, the set $U:=\{\varepsilon\in[0,1]\mid F(\varepsilon)\geq 1\}$ is open in $[0,1]$: if $\varepsilon_{0}\in U$, lower semi-continuity gives $\delta>0$ such that $F(\varepsilon)>F(\varepsilon_{0})-\frac{1}{2}$ for all $\varepsilon$ in $(\varepsilon_{0}-\delta,\varepsilon_{0}+\delta)\cap[0,1]$. Since $F(\varepsilon_{0})\geq 1$ and $F$ is integer, this implies $F(\varepsilon)\geq 1$ on that neighbourhood, hence the neighbourhood is contained in $U$.
We have shown $F(\varepsilon)\geq 1$ for all $\varepsilon>0$; thus $(0,1]\subseteq U$. Because $U$ is open and contains $(0,1]$, any open neighbourhood of $1$ in $[0,1]$ intersects $(0,1)\subset U$; openness of $U$ then forces $1\in U$. Therefore $F(1)\geq 1$, i.e., $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{1}})\geq 1$.
Since $\nabla^{C_{1}}=\nabla^{C}$ and $g_{1}=g_{M}$, the bound holds for the physical metric. ∎
4.5 Application to the SOC BEC model
Proposition 4.9 (Application to the SOC BEC model).
In the spin‑orbit‑coupled BEC model described in Section 2, the extended parameter space is given by the product manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$. The metric (2.1) is of the Kaluza–Klein type $g\;=\;\pi_{1}^{*}g_{1}\;+\;\sum_{a=1}^{m}\bigl(\theta^{a}+\pi_{1}^{*}A^{a}\bigr)^{2}$ with the base $M_{1}=T^{2}_{\mathrm{BZ}}$ and the fibre $M_{2}=S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$. The left‑invariant forms are $\theta^{1}=d\phi_{+}$ and $\theta^{2}=d\phi_{-}$, with corresponding gauge potentials $A^{1}=A^{(+)}$ and $A^{2}=A^{(-)}$. The fibres are flat since $d\theta^{a}=0$.
For this specific geometry, the following properties hold:
  1. 1.
    The obstruction kernel vanishes: $\mathcal{K}=0$, as shown by Lemma 2.5 under Assumption 2.1.
  2. 2.
    $\mathcal{P}_{2}(M_{1},g)\cong\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})\cong\mathbb{R}\,\mathrm{vol}_{\mathrm{BZ}}$.
  3. 3.
    The obstruction space vanishes identically: $\mathcal{K}=0$.
Furthermore, if the Berry curvatures satisfy $F^{(+)}]\neq 0$ and $F^{(-)}]\neq 0$ in $H^{2}(T^{2}_{\mathrm{BZ}})$, then the mixed tensor rank is $r=1$, and Theorem 4.8 implies:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
Proof.
The identification of the BEC metric with the Kaluza–Klein framework follows directly from the construction in Section 2. The first item follows directly from Lemma 2.5; under Assumption 2.1 the non‑vanishing Berry curvature forces $\mathcal{K}=0$.
On the base manifold $M_{1}=T^{2}_{\mathrm{BZ}}$, the restriction map onto $\mathcal{P}_{2}(T^{2}_{\mathrm{BZ}},g_{\mathrm{BZ}})$ is an isomorphism, identifying the space of parallel 2-forms with the span of the Riemannian volume form. Since $\mathcal{P}_{1}(M_{2},g)=\{0\}$, the tensor products in the definition of $\mathcal{K}$ vanish, ensuring that the obstruction space is trivial.
Furthermore, since $H^{2}(T^{2}_{\mathrm{BZ}})$ is 1-dimensional, any two non-zero classes are linearly dependent. Thus, we may write $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda\in\mathbb{R}$, and the mixed component of the torsion class becomes:
$\omega]_{\mathrm{mixed}}=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr),$
which represents a single elementary tensor, implying $r=1$. Applying Theorem 4.8 with $r=1$ and $\dim\mathcal{K}=0$ yields the lower bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$. ∎
5 Topological Lower Bound and Its Physical Consequences
With the geometric construction of Section 2 and the cohomological data of Section 3 at hand, we can now apply the extended Pigazzini-Toda theorem (Theorem 4.8) to the physical connection $\nabla^{C}$. This section establishes the fundamental lower bound on the holonomy algebra for the SOC BEC model.
5.1 Statement of the bound
The extended PT theorem (Theorem 4.8) establishes that for a metric connection $\nabla^{C}$ with totally skew-symmetric torsion $T$ on the Kaluza-Klein manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the dimension of the off-diagonal holonomy is constrained by the topology of the torsion class and the geometry of the metric. Specifically, under Assumption 2.1 and with $\nabla^{LC}T\neq 0$, we have:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;r-\dim\mathcal{K},$ (5.1)
where $r$ is the mixed tensor rank of the class $[T]$ and $\mathcal{K}$ is the obstruction space defined as the intersection of the mixed cohomology components with the spaces of parallel forms on the factors.
5.2 Main theorem: topological lower bound for SOC BECs
Theorem 5.1 (Topological lower bound for SOC BECs).
Let $M=T_{\mathrm{BZ}}^{2}\times S_{\phi_{+}}^{1}\times S_{\phi_{-}}^{1}$ be the parameter space equipped with the Kaluza-Klein metric $g_{M}$ (2.1) and the metric connection $\nabla^{C}$ with torsion $T^{b}=F^{(+)}\wedge\Theta^{(+)}+F^{(-)}\wedge\Theta^{(-)}$. Under Assumption 2.1 (constant Berry curvatures), the following hold:
  1. 1.
    The mixed tensor rank of the torsion class $T^{b}]$ is $r=1$, which is the maximal value for this configuration.
  2. 2.
    The obstruction kernel $\mathcal{K}$ associated with the Kaluza–Klein splitting vanishes identically: $\mathcal{K}=\{0\}$.
  3. 3.
    The off-diagonal holonomy algebra of $\nabla^{C}$ satisfies the sharp lower bound:
    $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$ (5.2)
Proof.
The proof proceeds by synthesizing the topological and geometric invariants derived in the preceding sections and applying the deformation argument developed in Section 4.
1. Mixed tensor rank. As established in Lemma 4.7, the rank $r$ is a purely topological invariant of the class $\omega]=[T^{b}]$. Under the Künneth decomposition for $H^{3}(M;\mathbb{R})$, the torsion class is expressed as $\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}]$. Since the cohomology group $H^{2}(T^{2}_{\mathrm{BZ}})$ is one-dimensional and both $F^{(\pm)}]$ are non-zero under Assumption 2.1, we can write $F^{(-)}]=\lambda[F^{(+)}]$ for some $\lambda=c^{(-)}/c^{(+)}\in\mathbb{R}$, which allows the class to be factorized as the simple tensor:
$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]+\lambda[d\phi_{-}]\bigr).$
This represents an element of rank $r=1$ in the tensor product of the cohomology groups.
2. Vanishing of the obstruction kernel. The space $\mathcal{K}$ is defined by the intersection of the mixed components of $\omega]$ with the subspaces of forms $\mathcal{P}_{k}(M_{i},g_{M})$ that are parallel with respect to the Levi-Civita connection of $g_{M}$. Under Assumption 2.1, Lemma 2.5 proves that the obstruction kernel vanishes: $\mathcal{K}=\{0\}$.
3. Application via deformation and lower semi-continuity. Consider the deformation family $\{g_{\varepsilon}\}_{\varepsilon\in[0,1]}$ from Definition 4.1, which interpolates between the product metric $g_{0}$ and the physical Kaluza-Klein metric $g_{M}=g_{1}$. For each $\varepsilon>0$, the metric $g_{\varepsilon}$ satisfies Assumption 2.1 with curvatures scaled by $\varepsilon$. By Lemma 4.2, we have $\mathcal{K}_{\varepsilon}=0$ and $r^{\sharp}_{\varepsilon}=1$ for all $\varepsilon>0$.
Applying Theorem 4.8 (the extension of the PT theorem to Kaluza-Klein metrics with flat fibres) to each $g_{\varepsilon}$ with $\varepsilon>0$, we obtain
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1\quad\forall\varepsilon>0,$
where $\nabla^{C_{\varepsilon}}$ is the connection built from $g_{\varepsilon}$ with the same torsion form $T$.
By Lemma 4.3, the function $\varepsilon\mapsto\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})$ is lower semi-continuous on $[0,1]$. Therefore,
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{1}})\geq\liminf_{\varepsilon\to 1}\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C_{\varepsilon}})\geq 1.$
Since $\nabla^{C_{1}}=\nabla^{C}$ and $g_{1}=g_{M}$, this establishes the bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ for the physical metric.
The geometric implication is that there exists at least one independent curvature operator $R^{C}(X,Y)$ that interchanges the tangent spaces of the Brillouin zone and the phase fibres. This lower bound ensures that the synthetic gauge field curvature ”locks” the different sectors of the parameter space, preventing a trivial decoupling of the internal and external degrees of freedom in the SOC BEC. ∎
5.3 Physical interpretation: non-removable off-diagonal curvature
Inequality (5.2) has a direct geometric meaning. The holonomy algebra $\mathfrak{hol}(\nabla^{C})$ is generated by the curvature operators $R^{C}(X,Y)$. The “off-diagonal” part $\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})$ consists of those generators that map the tangent space of the Brillouin zone $T_{k}T^{2}_{\mathrm{BZ}}$ into the phase directions $\mathbb{R}\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$ and vice-versa. The bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ therefore guarantees that there exists at least one linearly independent curvature operator that mixes momentum with the phase directions.
In physical terms, this means that the synthetic gauge curvature of the SOC BEC cannot be made completely block-diagonal by any smooth gauge transformation or deformation of the condensate wavefunction. One cannot “flatten” the Berry curvature along both phase directions simultaneously for all points in the Brillouin zone. The topology of the mixed class $\omega]$ forces a minimal geometric entanglement between the momentum and phase degrees of freedom.
5.4 Persistence of the bound when the total Chern number vanishes
A crucial feature of the bound is that it is insensitive to the value of the total first Chern number
$c_{1}^{\,\mathrm{tot}}=\frac{1}{2\pi}\int_{T^{2}_{\mathrm{BZ}}}\bigl(F^{(+)}+F^{(-)}\bigr).$
Corollary 5.2 (Bound independent of the total Chern number).
Let $c_{1}^{(+)}=\frac{1}{2\pi}\int_{T^{2}}F^{(+)}$ and $c_{1}^{(-)}=\frac{1}{2\pi}\int_{T^{2}}F^{(-)}$. If $c_{1}^{(+)}\neq 0$ and $c_{1}^{(-)}=-c_{1}^{(+)}$, then $c_{1}^{\,\mathrm{tot}}=0$. Nevertheless, under Assumption 2.1 with $c^{(\pm)}\neq 0$, the mixed tensor rank remains $r=1$ and Theorem 5.1 still gives
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1.$
Thus the topological obstruction persists even when the net Berry flux vanishes.
Proof.
If $c_{1}^{(+)}\neq 0$ then $F^{(+)}]\neq 0$ in $H^{2}(T^{2})$. The condition $c_{1}^{(-)}=-c_{1}^{(+)}$ implies $F^{(-)}]=-[F^{(+)}]$ in cohomology (because integration gives an isomorphism $H^{2}(T^{2})\cong\mathbb{R}$). Hence both classes are non-zero and the mixed tensor rank attains its maximum $r=1$. The vanishing of $\mathcal{K}$ does not depend on the values of the Chern numbers but only on the non-existence of parallel one-forms on the fibres, which holds as long as $F^{(\pm)}\neq 0$ (guaranteed by Assumption 2.1). ∎
This corollary is consistent with the geometric setup of Section 2: the global product structure of $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$ already implies $c_{1}^{\mathrm{tot}}=0$. The example $c_{1}^{(+)}=-c_{1}^{(-)}$ is therefore not a special case but a generic realization of the trivial-bundle regime under constant curvatures with $\lambda=-1$.
This demonstrates that $r^{\sharp}$ detects a local topological constraint that survives when the familiar global invariant vanishes. While the Chern number quantizes the net flux, it can be zero due to cancellation of local contributions; the mixed-rank bound, in contrast, certifies that the curvature cannot be simultaneously flattened along the two independent phase directions. The lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ thus captures a strictly local topological obstruction that remains active even when global charges cancel.
5.5 Invariance under metric deformations
The lower bound (5.2) depends only on the topological data $r$ and on the spaces of parallel forms $\mathcal{P}_{k}(M_{i})$. In our construction, these spaces are determined by the Kaluza–Klein metric $g_{M}$, which in turn is built from the physical Berry connections $A^{(\pm)}$. Any smooth deformation of the system that does not change the cohomology classes of $A^{(\pm)}$ and that preserves the vanishing of $\mathcal{P}_{1}(S^{1}_{\phi_{\pm}})$ (i.e., that does not make the Berry curvatures exactly flat) leaves the bound unchanged. This stability is ensured by Lemma 4.5, which guarantees that the non-parallelism condition $\nabla^{LC}T\neq 0$ is open in the space of metrics. This makes the result robust against typical experimental perturbations.
The inequality $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ is the main result of this paper. It provides a cohomologically defined, locally non-trivial obstruction to the complete removal of Berry curvature in a two-component SOC BEC, valid even when the total Chern number is zero. In the following section, we illustrate how the same cohomological framework reproduces, in a particular Rashba–Dresselhaus texture, the familiar Chern-number bound on vortex count, while emphasizing that the holonomy bound operates beyond the Chern-number paradigm.
6 Physical Consequences and Experimental Signatures
The lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ imposes fundamental structural constraints on the synthetic gauge architecture of SOC BECs. These manifest as irreducible couplings between the external (momentum) and internal (phase) degrees of freedom.
6.1 Geometric obstruction to gauge flattening
The bound ensures the existence of at least one independent off-diagonal curvature operator that non-trivially mixes the momentum tangent space $T_{k}T^{2}_{\mathrm{BZ}}$ with the phase fibre directions $\mathbb{R}\langle\partial_{\phi_{+}},\partial_{\phi_{-}}\rangle$. In terms of the synthetic gauge field components, this implies that no smooth gauge transformation or continuous deformation of the condensate order parameter can simultaneously “flatten” the curvature components:
$\Omega_{j,\pm}:=\frac{\partial A_{\phi_{\pm}}}{\partial k_{j}}-\frac{\partial A_{k_{j}}}{\partial\phi_{\pm}}.$
The topology of the mixed cohomology class $\omega]$ enforces a minimal geometric coupling between the Brillouin zone dynamics and the dual-phase evolution. Unlike trivial gauge fields, where the connection can be locally reduced to a longitudinal form, the non-vanishing rank $r^{\sharp}$ certifies that at least one independent transverse coupling is irreducible. This “topological locking” prevents the independent manipulation of the momentum and the relative phases $\phi_{\pm}$ in any experimental protocol that preserves the SOC texture.
6.2 Locality and robustness beyond the Chern-number paradigm
The primary physical advantage of the holonomy bound lies in its sensitivity to the local structure of the Berry curvature, whereas the total Chern number $c_{1}^{\mathrm{tot}}$ is a global aggregate that may vanish due to cancellation.
In the constant-curvature setting of Assumption 2.1, the local field strength is captured by the constants $c^{(\pm)}$. In configurations where the individual fluxes satisfy $F^{(+)}]\approx-[F^{(-)}]$, the global invariant $c_{1}^{\mathrm{tot}}$ vanishes, potentially suggesting a topologically trivial regime. However, the inequality $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ persists as long as both constants $c^{(+)}$ and $c^{(-)}$ are non-zero, regardless of their sum. This identifies $r^{\sharp}$ as a robust local invariant: it detects the persistence of topological effects, such as the non-integrability of the Berry connection, even in compensated systems or “synthetic vacuum” configurations where net topological charges are absent.
From an experimental perspective, this implies that effects typically associated with non-trivial topology—such as non-Abelian transport characteristics or anomalous velocity contributions—can be observed in SOC BECs even when the total Chern number is zero, provided the mixed tensor rank $r$ is preserved.
7 Illustrative Examples: From Chern Numbers to the Mixed Rank
The general lower bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$ derived in Section 5 establishes a universal constraint for generic two-component SOC BECs. In this section, we illustrate the operational significance of this bound in two paradigmatic scenarios: first, the standard Rashba–Dresselhaus texture, where our framework recovers the phenomenology of Chern-number-induced obstructions; second, a configuration with vanishing net Chern flux, where the mixed-rank invariant $r^{\sharp}$ reveals topological features inaccessible to traditional global invariants.
7.1 Example 1: Rashba–Dresselhaus texture and the maximal rank $r=1$
Consider a two-dimensional, two-level Hamiltonian of the Rashba–Dresselhaus (RD) type:
$\hat{H}_{\mathrm{RD}}(k)\;=\;\frac{1}{2m}\bigl(\mathbf{p}-\mathbf{A}_{\mathrm{syn}}(k)\bigr)^{2}\;+\;\mathbf{B}_{\mathrm{syn}}(k)\cdot\boldsymbol{\sigma},\qquad k=(k_{x},k_{y})\in T^{2}_{\mathrm{BZ}},$
where $\mathbf{A}_{\mathrm{syn}}(k)$ and $\mathbf{B}_{\mathrm{syn}}(k)$ represent the synthetic vector potential and Zeeman field, respectively. For generic coupling parameters $\kappa$ and $\Omega$, the lower dressed band induces a spin texture map $\mathbf{n}:T^{2}_{\mathrm{BZ}}\to S^{2}$.
The associated eigenline bundle $\mathbb{L}\to T^{2}_{\mathrm{BZ}}$ is the pullback of the tautological line bundle over $\mathbb{C}P^{1}\cong S^{2}$ via $\mathbf{n}$. The first Chern number $c_{1}(\mathbb{L})$ corresponds to the degree of the map $\mathbf{n}$. For a minimal RD texture with $d=1$, the Berry curvature $F$ represents a non-trivial cohomology class $F]\in H^{2}(T^{2}_{\mathrm{BZ}};\mathbb{Z})$, such that $\frac{1}{2\pi}\int F=\pm 1$.
In the extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the synthetic torsion class $\omega]$ is given by:
$\omega]=[F]\otimes[d\phi_{+}]\;+\;[F]\otimes[d\phi_{-}]\in H^{3}(M;\mathbb{R}).$
Since $\dim H^{2}(T^{2}_{\mathrm{BZ}})=1$, the mixed components are linearly dependent over the base, allowing the class to be expressed as the simple tensor:
$\omega]=[F]\otimes\bigl([d\phi_{+}]+[d\phi_{-}]\bigr).$
Consequently, the mixed tensor rank is $r=1$. Given that the obstruction kernel $\mathcal{K}$ vanishes in this geometry (as shown in Lemma 2.5 with $\lambda=1$), we obtain $r^{\sharp}=1$. Theorem 5.1 then yields:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
Remark 7.1.
For a global Rashba texture with non-zero Chern number, the underlying principal bundle is topologically non-trivial, preventing a global product decomposition. However, as noted in Section 2, our analysis can be applied within local trivializations. The mixed rank $r^{\sharp}$ correctly identifies the ’locking’ of momentum and phase degrees of freedom even in such local patches. While the non-vanishing Chern number already implies the existence of vortices on the Brillouin zone, the invariant $r^{\sharp}$ provides an algebraically refined constraint, certifying that the momentum and phase degrees of freedom remain fundamentally coupled through the off-diagonal curvature.
7.2 Example 2: Topological persistence under vanishing net Chern flux
The efficacy of the mixed-rank invariant is most evident when global topological charges vanish. Consider a configuration where the two $U(1)$ sectors carry Berry curvatures $F^{(+)}$ and $F^{(-)}$ with equal and opposite integrated fluxes:
$\frac{1}{2\pi}\int_{T_{\mathrm{BZ}}^{2}}F^{(+)}=+1,\qquad\frac{1}{2\pi}\int_{T_{\mathrm{BZ}}^{2}}F^{(-)}=-1.$
In this case, the total Chern number $c_{1}^{\mathrm{tot}}=c_{1}^{(+)}+c_{1}^{(-)}=0$. Traditional Chern-number diagnostics would categorize this regime as topologically trivial. Crucially, the vanishing of the total Chern flux ensures that the synthetic connection $A$ can be defined as a global 1-form, making the product structure $M=T_{\mathrm{BZ}}^{2}\times S^{1}\times S^{1}$ globally well-defined and consistent with our geometric framework. In this globally trivialized setting, the torsion class remains:
$\omega]=[F^{(+)}]\otimes[d\phi_{+}]+[F^{(-)}]\otimes[d\phi_{-}].$
Since $F^{(-)}]=-[F^{(+)}]$ in $H^{2}(T_{\mathrm{BZ}}^{2})$, the class factorizes into a single simple tensor:
$\omega]=[F^{(+)}]\otimes\bigl([d\phi_{+}]-[d\phi_{-}]\bigr).$
Thus, $r=1$ and $r^{\sharp}=1$, forcing the holonomy bound $\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 1$. This proves that even in the absence of a net global flux, the Berry curvature cannot be flattened simultaneously along both phase directions. Such a configuration is protected by a local topological obstruction that is invisible to integrated invariants but perfectly captured by the mixed cohomology rank on the product manifold.
To illustrate how the bound materializes geometrically, consider the concrete choice (satisfying Assumption 2.1)
$A^{(+)}=\frac{1}{4\pi}(k_{x}dk_{y}-k_{y}dk_{x}),\qquad A^{(-)}=-A^{(+)},$
which yields constant Berry curvatures $F^{(\pm)}=\pm\frac{1}{2\pi}dk_{x}\wedge dk_{y}$ and satisfies ($\ast$) on a torus of period $2\pi$ in each direction. The corresponding Kaluza–Klein metric is
$g_{M}=dk_{x}^{2}+dk_{y}^{2}+\Bigl(d\phi_{+}+\tfrac{1}{4\pi}(k_{x}dk_{y}-k_{y}dk_{x})\Bigr)^{2}+\Bigl(d\phi_{-}-\tfrac{1}{4\pi}(k_{x}dk_{y}-k_{y}dk_{x})\Bigr)^{2}.$
A straightforward computation of the Levi-Civita connection gives the non-zero mixed Christoffel symbols
$\Gamma^{k_{x}k_{y}}_{\phi_{+}}=-\Gamma^{k_{y}k_{x}}_{\phi_{+}}=\frac{1}{8\pi},\qquad\Gamma^{k_{x}k_{y}}_{\phi_{-}}=-\Gamma^{k_{y}k_{x}}_{\phi_{-}}=-\frac{1}{8\pi},$
all other mixed components being zero.
The torsion 3-form is
$T=F^{(+)}\wedge d\phi_{+}+F^{(-)}\wedge d\phi_{-}=\frac{1}{2\pi}\,dk_{x}\wedge dk_{y}\wedge(d\phi_{+}-d\phi_{-}).$
For the connection $\nabla^{C}=\nabla^{LC}+\tfrac{1}{2}T$, the off-diagonal curvature components evaluate to
$R^{C}(\partial_{k_{x}},\partial_{k_{y}})\,\partial_{\phi_{+}}=\frac{1}{4\pi}\,\partial_{\phi_{-}}\neq 0,\qquad R^{C}(\partial_{k_{x}},\partial_{k_{y}})\,\partial_{\phi_{-}}=-\frac{1}{4\pi}\,\partial_{\phi_{+}}\neq 0,$
where the contributions from the Levi-Civita curvature and the torsion combine additively.
These two operators are linearly independent, confirming that
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\geq 2>1=r^{\sharp}.$
Remark 7.2.
The explicit computation confirms that:
  1. 1.
    The off-diagonal curvature is non‑vanishing even though $c_{1}^{\mathrm{tot}}=0$.
  2. 2.
    The topological bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq 1$ is not merely existential; it is realized by concrete curvature operators that can be written down in closed form.
  3. 3.
    The geometric obstruction manifests as a direct coupling between momentum derivatives $(\partial_{k_{x}},\partial_{k_{y}})$ and phase derivatives $(\partial_{\phi_{+}},\partial_{\phi_{-}})$, preventing any gauge transformation from simultaneously diagonalizing the connection with respect to the product splitting $T_{\mathrm{BZ}}^{2}\times S^{1}\times S^{1}$.
The fact that $\dim\mathfrak{hol}^{\mathrm{off}}=2$ in this specific example shows that the bound $\geq 1$ is sharp in the sense that it cannot be improved without additional assumptions, yet particular geometries may yield a larger off-diagonal holonomy algebra.
7.3 Comparison: Global Invariants vs. Mixed Rank
The distinction between $c_{1}^{\mathrm{tot}}$ and $r^{\sharp}$ reflects the difference between integrated and structural topology. The Chern number $c_{1}^{\mathrm{tot}}$ is a global aggregate that can vanish via local cancellations. In contrast, the mixed rank $r$ counts the number of irreducible tensorial couplings between the factor manifolds.
In our geometry, $r=1$ is attained whenever at least one individual Berry curvature class is non-zero. It serves as a measure of the irreducible geometric coupling between the Brillouin zone torus and the phase fibres. The bound $\dim\mathfrak{hol}^{\mathrm{off}}\geq r^{\sharp}$ therefore constrains the algebraic complexity of the curvature, ensuring that the synthetic gauge field remains ”locked” in a non-trivial configuration even when the net topological charge is zero.
7.4 Summary
The Rashba–Dresselhaus case shows that $r^{\sharp}$ is consistent with established Chern-number results. However, the $c_{1}^{\mathrm{tot}}=0$ case demonstrates that the mixed-rank framework provides a finer classification, detecting locally non-removable curvature that conventional global invariants fail to resolve. This makes $r^{\sharp}$ a superior tool for analyzing the stability of synthetic gauge fields in complex SOC BEC architectures.
8 Beyond the Chern Number: Local Rail-Tracks and Prospective Protocols
While the Chern number $c_{1}$ quantifies the net vorticity through a global integral over the Brillouin zone, the lower bound on the off-diagonal holonomy dimension, $r^{\sharp}$, acts as a local structural constraint. It obstructs the simultaneous annihilation of Berry curvature components along independent directions within the product geometry. This phenomenon can be visualized as the Berry curvature being constrained to move along topological “rail-tracks” in the configuration space: even when the total topological charge is tuned to zero, specific directional components remain topologically locked and cannot be erased by any smooth gauge transformation.
In the two-component SOC BEC manifold $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, the bound $r^{\sharp}=1$ is non-degenerate. Under Assumption 2.1, the constant, non-zero Berry curvatures break the integrability of the horizontal distribution, ensuring $\mathcal{K}=0$. Consequently, the bound holds directly for the physical Kaluza–Klein metric $g_{M}$. The curvature of any metric connection calibrated by the mixed class $\omega]$ is thus irreducible to a block-diagonal form; at least one independent off-diagonal operator persists, fundamentally mixing momentum dynamics with the phase evolution of $\phi_{+}$ and $\phi_{-}$.
This geometric prediction suggests concrete interferometric tests. Consider the simultaneous measurement of the Berry phases $\Phi_{B}^{+}(\gamma)$ and $\Phi_{B}^{-}(\gamma)$ accumulated along a closed loop $\gamma\subset T^{2}_{\mathrm{BZ}}$ by separately addressing the global and relative phase sectors. If the gauge field were fully reducible (i.e., if the holonomy were trivial), one could choose a gauge where both phases vanish for every $\gamma$. The bound $r^{\sharp}=1$ forbids such a simultaneous vanishing across all loops. For any pair of independent cycles $\{\gamma_{1},\gamma_{2}\}$ in $T^{2}_{\mathrm{BZ}}$, the two-component Berry-phase vector:
$\mathbf{\Phi}_{B}=\bigl(\Phi_{B}^{+}(\gamma_{1}),\Phi_{B}^{-}(\gamma_{1}),\Phi_{B}^{+}(\gamma_{2}),\Phi_{B}^{-}(\gamma_{2})\bigr)$
cannot be continuously deformed to zero while preserving the mixed cohomology class $\omega]$.
Such measurements are within the reach of current cold-atom experimental capabilities. Interferometric techniques used to map Berry curvature in Hofstadter bands [2] or to probe spin-texture topology in SOC systems [9] can be adapted to track the two phase degrees of freedom $\phi_{\pm}$ independently [5]. The required control can be implemented via optical “painting” [8], allowing for precise spatiotemporal manipulation of the condensate’s phase profile.
Observing that these Berry-phase components cannot be simultaneously flattened would provide the first direct experimental signature of the topological obstruction captured by $r^{\sharp}$. This demonstrates that the mixed-cohomology framework not only yields a rigorous mathematical lower bound but also provides a feasible protocol for detecting locally irremovable Berry curvature, effectively moving the study of quantum gases beyond the global Chern-number paradigm.
9 Conclusions
We have demonstrated that a two-component spin–orbit-coupled (SOC) BEC possesses an intrinsic geometric obstruction that prevents the simultaneous flattening of Berry curvature along the independent phase directions $\phi_{\pm}$. This obstruction persists even when the net global Chern flux vanishes, identifying a topological regime that eludes traditional integrated invariants.
The extended parameter space $M=T^{2}_{\mathrm{BZ}}\times S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}}$, equipped with the natural Kaluza–Klein metric $g_{M}$ induced by synthetic gauge fields, carries a metric connection $\nabla^{C}$ with totally skew-symmetric torsion. Under Assumption 2.1 (constant Berry curvatures), the harmonic part of this torsion takes the explicit form given in Lemma 3.1 and represents a mixed cohomology class $\omega]$ with mixed tensor rank $r=1$ for any generic configuration where the individual Berry curvature classes are non-trivial. Due to the one-dimensionality of $H^{2}(T^{2}_{\mathrm{BZ}})$, these classes are necessarily proportional, forcing $\omega]$ to take the form of a single simple tensor. Crucially, we have shown that the non-vanishing Berry curvature acts as a geometric obstruction to the integrability of the horizontal distribution, ensuring the vanishing of the obstruction kernel $\mathcal{K}=0$. Consequently, the bound $r^{\sharp}=1$ holds directly for the physical metric $g_{M}$.
Applying the PT lower bound yields the sharp inequality:
$\dim\mathfrak{hol}^{\mathrm{off}}(\nabla^{C})\;\geq\;1.$
This algebraic constraint implies that momentum degrees of freedom are fundamentally “locked” to the phase sectors, an irreducible coupling that cannot be removed by smooth gauge transformations. The persistence of $r^{\sharp}$ in regimes with zero total Chern number—as illustrated by our refined Rashba–Dresselhaus analysis—highlights that the mixed-rank bound operates beyond the Chern-number paradigm, detecting structural features of the gauge field that integrated invariants fail to resolve.
Experimentally, this framework suggests interferometric protocols to resolve independent Berry phases, providing a direct signature of locally irremovable curvature. Conceptually, it establishes a new mechanism for topological robustness in synthetic gauge systems, controlled by the algebraic structure of mixed cohomology components. We expect this framework to be readily applicable to multicomponent cold-atom systems and synthetic-dimensional lattices, where multiple internal degrees of freedom couple to momentum-space geometry.
Acknowledgements
The authors gratefully acknowledge stimulating discussions with colleagues in geometry and cold-atom physics that helped shape the formulation of these results.
Appendix
Appendix A Parallel Vertical 1-Forms
This appendix provides the detailed proof of the condition for vertical 1-forms to be parallel, which is used in the proof of Lemma 2.5. We consider the Kaluza-Klein metric
$g_{M}=\pi^{*}g_{\mathrm{BZ}}+\Theta^{(+)}\otimes\Theta^{(+)}+\Theta^{(-)}\otimes\Theta^{(-)},$
with $\Theta^{(\pm)}=d\phi_{\pm}+\pi^{*}A^{(\pm)}$ and $F^{(\pm)}=dA^{(\pm)}=c^{(\pm)}\operatorname{vol}_{\mathrm{BZ}}$ (Assumption 2.1). Let $\lambda=c^{(-)}/c^{(+)}$.
Orthonormal Frame
Choose coordinates $(x,y)$ on $T^{2}_{\mathrm{BZ}}$ such that $g_{\mathrm{BZ}}=dx^{2}+dy^{2}$ and $\operatorname{vol}_{\mathrm{BZ}}=dx\wedge dy$. An orthonormal frame for $g_{M}$ is given by
$e_{1}=\partial_{x},\quad e_{2}=\partial_{y},\quad e_{3}=\partial_{\phi_{+}},\quad e_{4}=\partial_{\phi_{-}},$
with the corresponding coframe
$e^{1}=dx,\quad e^{2}=dy,\quad e^{3}=\Theta^{(+)},\quad e^{4}=\Theta^{(-)}.$
Connection 1-Forms
The connection 1-forms $\omega^{a}_{\;b}$ for the Levi-Civita connection are defined by
$\nabla^{LC}e_{a}=\omega_{a}^{\;b}\otimes e_{b},$
where $\omega_{ab}=g_{M}(\nabla^{LC}e_{a},e_{b})$ are skew-symmetric. For our Kaluza-Klein metric with flat fibers and constant curvatures, the non-zero connection 1-forms are:
$\omega_{13}$ $=-\omega_{31}=\frac{1}{2}c^{(+)}e^{2},$
$\omega_{14}$ $=-\omega_{41}=\frac{1}{2}c^{(-)}e^{2},$
$\omega_{23}$ $=-\omega_{32}=-\frac{1}{2}c^{(+)}e^{1},$
$\omega_{24}$ $=-\omega_{42}=-\frac{1}{2}c^{(-)}e^{1}.$
All other $\omega_{ab}=0$. This can be derived from the structure equations $de^{a}+\omega^{a}_{\;b}\wedge e^{b}=0$ and the known expressions for $de^{3}$ and $de^{4}$.
Parallel Vertical 1-Forms
A vertical 1-form can be written as
$\eta=h_{+}\Theta^{(+)}+h_{-}\Theta^{(-)}=h_{+}e^{3}+h_{-}e^{4},$
where $h_{\pm}\in C^{\infty}(M)$. The condition for $\eta$ to be parallel, $\nabla^{LC}\eta=0$, becomes
$\nabla^{LC}\eta=(dh_{+}\otimes e^{3}+h_{+}\nabla^{LC}e^{3})+(dh_{-}\otimes e^{4}+h_{-}\nabla^{LC}e^{4})=0.$
Using $\nabla^{LC}e^{3}=-\omega_{3}^{\;b}\otimes e_{b}=\omega_{13}e^{1}\otimes e_{1}+\omega_{23}e^{2}\otimes e_{1}+\omega_{31}e^{1}\otimes e_{2}+\omega_{32}e^{2}\otimes e_{2}$ (and similarly for $e^{4}$), we compute the components of $\nabla^{LC}\eta$.
The components along $e^{1}\otimes e^{3}$ and $e^{2}\otimes e^{3}$ give:
$\frac{\partial h_{+}}{\partial x}=0,\quad\frac{\partial h_{+}}{\partial y}=0,$
so $h_{+}$ is constant along the base. Similarly, from the components along $e^{1}\otimes e^{4}$ and $e^{2}\otimes e^{4}$, we get that $h_{-}$ is also constant along the base.
Now, consider the components along $e^{1}\otimes e^{1}$ and $e^{2}\otimes e^{2}$. These yield:
$h_{+}\omega_{13}(e_{1})+h_{-}\omega_{14}(e_{1})=0,\quad h_{+}\omega_{23}(e_{2})+h_{-}\omega_{24}(e_{2})=0.$
Using the expressions for $\omega_{ab}$, we obtain:
$h_{+}\cdot\frac{1}{2}c^{(+)}+h_{-}\cdot\frac{1}{2}c^{(-)}=0,$
which simplifies to
$h_{+}c^{(+)}+h_{-}c^{(-)}=0\quad\text{or}\quad h_{+}+\lambda h_{-}=0.$
Finally, the vertical derivatives (components along $e^{3}\otimes e^{3}$, etc.) imply that $h_{\pm}$ are also constant along the fibers. Hence, $h_{+}$ and $h_{-}$ are global constants satisfying (A.1).
Therefore, the space of vertical parallel 1-forms is at most one-dimensional, spanned by $\lambda\Theta^{(+)}+\Theta^{(-)}$. However, when expressed in terms of the basis $\{d\phi_{+},d\phi_{-}\}$ of $H^{1}(S^{1}_{\phi_{+}}\times S^{1}_{\phi_{-}})$, a vertical parallel 1-form corresponds to a linear combination $c_{+}d\phi_{+}+c_{-}d\phi_{-}$ such that the constants satisfy (A.1). This completes the derivation of condition ()(*) in the proof of Lemma 2.5.
Appendix B Computation of the Obstruction Map $\Psi_{p}$
This appendix provides the detailed computation of the obstruction map $\Psi_{p}$ defined in Section 4. We work under Assumption 2.1 and use the explicit harmonic representative from Lemma 3.1.
Geometric Setup
Recall the harmonic representative of the torsion class:
$\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)}),\quad a_{-}=\lambda a_{+},\quad\lambda=c^{(-)}/c^{(+)}.$
The subspace $\mathcal{V}_{2,1}$ is spanned by the simple tensor
$\xi_{0}=\operatorname{vol}_{\mathrm{BZ}}\otimes(d\phi_{+}+\lambda d\phi_{-}).$
Any element of $\mathcal{V}_{2,1}$ can be written as $\xi=t\xi_{0}$ for some $t\in\mathbb{R}$, which corresponds to taking $a_{+}=t$ and $a_{-}=\lambda t$.
Definition of $\Psi_{p}$
At a point $p=(k,\phi_{+},\phi_{-})\in M$, choose horizontal vectors $X,Y\in\mathcal{H}_{p}$ such that $\operatorname{vol}_{\mathrm{BZ}}(X,Y)=1$. Choose a vertical vector $Z\in\mathcal{V}_{p}$ such that $(d\phi_{+}+\lambda d\phi_{-})(Z)=1$. For $\xi=\operatorname{vol}_{\mathrm{BZ}}\otimes(a_{+}d\phi_{+}+a_{-}d\phi_{-})\in\mathcal{V}_{2,1}$, define
$\Psi_{p}(\xi)=\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}(X,Y,Z),$
where $(\cdot)_{\mathrm{off}}$ denotes the off-diagonal part (mapping horizontal to vertical or vice versa). As explained in the main text, the map is well-defined and linear.
Key Computation
We need to compute $(\nabla^{LC}T)_{\mathrm{off}}(X,Y,Z)$. Note that the torsion 3-form $T$ is cohomologous to $\omega_{h}$, but $T$ itself is not necessarily harmonic. However, the definition of $\Psi_{p}$ in the original PT theorem uses the harmonic representative to construct the map. Therefore, we compute using $\omega_{h}$ as the representative of the cohomology class.
Consider the covariant derivative of $\omega_{h}$:
$\nabla^{LC}\omega_{h}=\nabla^{LC}\bigl(\operatorname{vol}_{\mathrm{BZ}}\wedge(a_{+}\Theta^{(+)}+a_{-}\Theta^{(-)})\bigr).$
Since $\operatorname{vol}_{\mathrm{BZ}}$ is parallel on the flat base, we have
$\nabla^{LC}\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge\bigl(a_{+}\nabla^{LC}\Theta^{(+)}+a_{-}\nabla^{LC}\Theta^{(-)}\bigr).$
From the connection 1-forms computed in Appendix A, we have
$\nabla^{LC}\Theta^{(+)}=\nabla^{LC}e^{3}=\frac{1}{2}c^{(+)}(e^{2}\otimes e^{1}-e^{1}\otimes e^{2}),$
and similarly for $\Theta^{(-)}$ with $c^{(+)}$ replaced by $c^{(-)}$. In coordinate-free notation, this can be written as
$\nabla^{LC}\Theta^{(\pm)}=\frac{1}{2}\pi^{*}F^{(\pm)}.$
Thus,
$\nabla^{LC}\omega_{h}=\operatorname{vol}_{\mathrm{BZ}}\wedge\frac{1}{2}\bigl(a_{+}\pi^{*}F^{(+)}+a_{-}\pi^{*}F^{(-)}\bigr).$
Now, evaluate this 3-form on the vectors $X,Y,Z$. Since $X,Y$ are horizontal and $Z$ is vertical, we have
$(\nabla^{LC}\omega_{h})(X,Y,Z)=\frac{1}{2}\bigl(a_{+}F^{(+)}(\pi_{*}X,\pi_{*}Y)+a_{-}F^{(-)}(\pi_{*}X,\pi_{*}Y)\bigr)\cdot\operatorname{vol}_{\mathrm{BZ}}(X,Y)\cdot 1,$
where we used that $\pi^{*}F^{(\pm)}(X,Y)=F^{(\pm)}(\pi_{*}X,\pi_{*}Y)$ and that $Z$ is vertical so the contraction with the vertical part is trivial. By construction, $\operatorname{vol}_{\mathrm{BZ}}(X,Y)=1$. Moreover, since $F^{(\pm)}=c^{(\pm)}\operatorname{vol}_{\mathrm{BZ}}$, we have
$F^{(\pm)}(\pi_{*}X,\pi_{*}Y)=c^{(\pm)}.$
Therefore,
$(\nabla^{LC}\omega_{h})(X,Y,Z)=\frac{1}{2}(a_{+}c^{(+)}+a_{-}c^{(-)}).$
Off-Diagonal Operator
The off-diagonal part of the curvature operator is then given by raising the index of this 3-form. Specifically, there exists a fixed off-diagonal operator $J:T_{p}M\to T_{p}M$ that exchanges horizontal and vertical directions such that
$(\nabla^{LC}T)_{\mathrm{off}}(X,Y,Z)=(a_{+}c^{(+)}+a_{-}c^{(-)})\,J.$
The precise form of $J$ is not needed for the rank computation; it suffices that $J$ is non-zero and off-diagonal.
Since $a_{-}=\lambda a_{+}$ and $c^{(-)}=\lambda c^{(+)}$, we have
$a_{+}c^{(+)}+a_{-}c^{(-)}=a_{+}c^{(+)}(1+\lambda^{2}).$
Thus,
$\Psi_{p}(\xi)=\frac{1}{2}a_{+}c^{(+)}(1+\lambda^{2})J.$
Kernel and Image
The expression above shows that $\Psi_{p}(\xi)=0$ if and only if $a_{+}=0$, which implies $a_{-}=0$. Hence, $\ker\Psi_{p}=\{0\}$. Moreover, the image of $\Psi_{p}$ is one-dimensional, spanned by the off-diagonal operator $J$. This confirms the claim made in Section 4 that $\dim\operatorname{im}\Psi_{p}=1$.